### v1

# Elastic, Quasielastic, and Superelastic Electron Scattering from Thermal Lattice Distortions in Perfect Crystals

Abstract
In conventional treatments of electron transport, momentum relaxation within a perfect, defect-free crystal is commonly assumed to require phonon creation or annihilation. Here we treat the crystal as finite and isolated, retaining the lattice center-of-mass (recoil) degree of freedom and enforcing conservation of total mechanical momentum alongside discrete crystal pseudomomentum. Starting from the density-density form of the electron-lattice interaction, we find that an electron in the interior of a perfect crystal admits strong, and in some regimes dominant, elastic momentum-relaxing scattering channels, in which momentum is conserved by recoil of the lattice background without phonon excitation. In addition, we identify mixed quasielastic and superelastic channels in which phonon occupations change but do not account fully for the electron’s momentum transfer. These results provide a microscopic basis for momentum relaxation that does not rely on local energy dissipation. They naturally reconcile a wide range of experimental observations—including weak localization, quantum oscillations, ultrasonic attenuation, and the separation of momentum and energy relaxation times—with predominantly elastic scattering in clean crystals. The framework clarifies how diffusive transport, including Planckian-scale diffusion, can emerge from elastic dynamics in a time-dependent lattice background.
preprint: APS/123-QED
I Introduction
Electron-lattice interactions are among the most important processes in condensed matter physics. They govern momentum relaxation, energy dissipation, and the loss of quantum coherence that underlies electrical resistance. In conventional transport theory, electron momentum relaxation in a perfect, defect-free crystal is commonly identified with inelastic scattering processes in which phonons are created or annihilated. Within this view, deflection of an electron by thermal lattice motion is inseparable from energy transfer to internal vibrational degrees of freedom.
This perspective is codified in the Bloch-Grüneisen and Boltzmann approaches to resistivity, which treat electronic transport as an incoherent sequence of scattering events and assume no electronic phase coherence beyond that retained in lowest-order scattering matrix elements. It is further reinforced by the Fröhlich interaction [10], which couples electrons explicitly to phonon creation and annihilation operators and therefore enforces inelasticity at first order in perturbation theory. Within this canonical framework, thermal deflection of electrons in a perfect crystal is taken to require phonon excitation, while elastic scattering is associated primarily with defects or impurities [39, 4].
Here we present evidence that this identification is overly restrictive. We show that thermally induced electron scattering in a pure crystal is often largely elastic when viewed on an event-by-event basis. Elastic and quasi-elastic momentum-relaxing channels exist even in a perfectly periodic lattice, without the necessity of phonon creation or annihilation. Previously, elastic scattering was almost exclusively associated with explicit disorder; we demonstrate that this association is not required by symmetry, conservation laws, or kinematics.
Early discussions of electronic transport already contained hints of this distinction. Peierls emphasized that electrical resistance requires degradation of electronic (crystal) momentum rather than dissipation of electronic energy, and noted explicitly that the lattice, as a massive collective object, can absorb momentum without requiring significant energy transfer [27]. Pippard similarly argued that defects and impurities are important but not indispensable, and that momentum relaxation could, in principle, proceed through elastic or quasi-elastic mechanisms even in very pure crystals [29]. These insights acknowledged the logical separation between momentum relaxation and energy dissipation, although this separation was not consistently maintained in subsequent formulations of transport theory.
In contrast, the canonical Fröhlich Hamiltonian and Bloch-Grüneisen framework [7] implicitly exclude rigid translation of the crystal as a whole from the dynamical description. The lattice center-of-mass (recoil) degree of freedom is effectively clamped, and momentum transfer is assigned exclusively to internal phonon excitations or to Umklapp processes. Once these assumptions are imposed, momentum loss is forced to coincide with energy dissipation, and elastic deflection by a perfect lattice is excluded by construction.
We revisit this assumption by treating the crystal as finite, isolated, and freely floating, and by explicitly retaining its total mechanical momentum. Enforcing conservation of total momentum more completely cannot be formally incorrect; at worst it may prove quantitatively unimportant. As we show below, however, restoring the lattice center-of-mass degree of freedom qualitatively alters the available scattering channels for an electron deep inside a pure conducting crystal. In particular, it reveals an elastic recoil channel that is absent by design in the conventional formulation.
The physical content of this channel is elementary. If a particle embedded in a composite object receives a sudden impulse, the momentum of the whole must change immediately, even if the internal degrees of freedom remain largely unexcited. A familiar example is provided by Mössbauer scattering, where momentum is absorbed predominantly by recoil of the entire lattice while phonon excitation is strongly suppressed. As we shall show, an analogous mechanism operates for electrons interacting with a thermally fluctuating but otherwise perfect crystal.
The ambiguity of elastic scattering
Elastic scattering is most familiar in the context of impurity or defect scattering. In that setting, momentum conservation is usually justified in one of two ways. First, the presence of an impurity breaks discrete translational symmetry, so crystal pseudomomentum need not be conserved. Second, within a particle picture, one argues that if an electron enters with one momentum and exits with another, the atomic lattice must recoil so as to conserve total mechanical momentum. Both statements are essentially correct. However, neither addresses the internal bookkeeping problem that arises for an electron scattering deep inside a perfectly periodic crystal: how does a change in the electron’s pseudomomentum communicate with the mechanical momentum of the lattice as a whole?
This unresolved point lies at the heart of the transport problem in a translationally invariant system. As we show below, it is resolved once three elements are made explicit: (i) the lattice is treated as a fully dynamical object rather than a fixed reference, (ii) pseudomomentum is assigned not only to the electron but also to the crystal background in the absence of the electron density (see the next section), and (iii) the electron-lattice interaction is formulated in density-density form [24, 10], which makes the exchange of pseudomomentum between electron and lattice explicit even in a perfectly periodic crystal.
In any untethered collection of atoms, an impulse imparted to even a single constituent instantaneously changes the total momentum of the assembly. While the subsequent redistribution of that momentum among internal degrees of freedom proceeds causally and at finite velocity, the change in the collective momentum itself is immediate. This total momentum is not an abstract bookkeeping device but a genuine dynamical variable that appears explicitly in the many-body wavefunction.
In the present formulation, this momentum conservation is enforced transparently. As shown in Eq. 12, a momentum transfer $\mathbf{q}$ to the foreground electron is accompanied by an equal and opposite momentum $\mathbf{q}$ carried by the remaining degrees of freedom, which we identify as the lattice background. This statement holds independently of whether phonons are created or destroyed in the process.
Finally, we clarify our use of the term elastic. In a finite-temperature many-body system, strictly elastic processes do not exist in an absolute sense. Even the dominant “elastic” line in Mössbauer scattering exhibits slight thermal broadening, reflecting the thermal motion of the emitting nucleus and its finite velocity autocorrelation time. The term elastic is nevertheless standard in this context, and we adopt it here with the same understood margin. When we use the term quasielastic, we instead refer to processes involving energy exchange that is appreciably larger than the thermal broadening scale, but still too small to account for the full momentum change of the electron.
The landscape seen by an electron
At temperatures of order $T\sim 100\,$K, thermally excited acoustic distortions in a typical metal generate substantial deformation-potential landscapes on nanometer length scales. Using standard values for the deformation potential ($D\sim 5$$15\,\mathrm{eV}$), sound velocity ($v_{s}\sim 3$$5\times 10^{3}\,\mathrm{m/s}$), and the characteristic thermal phonon wavevector $q_{T}\sim k_{B}T/\hbar v_{s}$, one finds order-of-magnitude estimates for the resulting deformation-potential field gradients of
$|\nabla V_{\rm def}|\sim 10^{4}\text{–}10^{6}\ \mathrm{V/cm}\qquad(T\sim 100\,\mathrm{K}).$ (1)
Over a representative lateral scale of $L\sim 10\,$nm, this corresponds to peak-to-peak variations of the screened scalar potential of order
$\Delta V_{\rm pp}\sim|\nabla V_{\rm def}|\,L\sim 10\text{–}300\ \mathrm{mV}.$ (2)
[…]
II Foreground, Background, and Pseudomomentum
The conceptual distinction between momentum, crystal momentum, and electrical current has been recognized since the early development of band theory, but its implications for microscopic transport have remained subtle. In a particularly prescient analysis, Kohn emphasized that crystal momentum is a conserved quantity associated with translational symmetry, while the electrical current need not be proportional to it, and that momentum may be exchanged with the lattice without the necessity of phonon excitation or energy dissipation [17]. This observation already points to the possibility that momentum conservation can be satisfied through coherent motion of the lattice background rather than through inelastic scattering processes. What has remained largely implicit, however, is how this distinction should be implemented at the level of a local scattering event occurring deep in the interior of a crystal. In particular, the possibility that total crystal may be explicitly partitioned between a tagged electron and the remaining lattice, both with pseudomomenta appearing as dynamical variables in the many-body wavefunction, has not been systematically developed.
The present work makes this partition explicit and operational, thereby extending Kohn’s insight to a microscopic framework in which elastic deflection, coherent lattice recoil, and diffusive transport can coexist naturally in a perfect crystal. In what follows, we formalize a symmetric foreground-background structure, derive the elastic scattering amplitude within first-order perturbation theory, and compute the elastic fraction explicitly. We show that this fraction is close to unity for copper at room temperature. We also clarify how the present formulation extends the Fröhlich framework without contradicting its successful predictions, and discuss the implications for the interpretation of transport phenomena in crystalline solids.
We distinguish explicitly between foreground and background degrees of freedom. The foreground is the object of interest, here a single electron or a collection of them. The background consists of all remaining degrees of freedom, including the lattice atoms and any other electrons not treated explicitly. This distinction is essential for identifying the relevant symmetries and conserved quantities.
In a crystalline environment, the electronic Hamiltonian is invariant under discrete translations of the electron by lattice vectors, with the background held fixed. The associated conserved quantity is crystal momentum (pseudomomentum); for an electron in a periodic potential this is the familiar Bloch momentum $\hbar\mathbf{k}$.
Crucially, the same discrete translational symmetry applies to the background (see Fig. 1). If the electron (foreground A) coordinate is held fixed and the entire background B is translated by a lattice vector, the Hamiltonian is again unchanged. A consistent application of translational symmetry therefore assigns pseudomomentum not only to the electron but also to the background, even though only their sum corresponds to a mechanical momentum.
One might object that the problem can always be formulated in the rest frame of the background, with only the electron moving. However, this choice of frame obscures the conservation of total momentum of the combined system and suppresses the dynamical role of background recoil, which is essential for understanding local scattering events in the interior of a perfect crystal.
Figure 1: Sliding Blocks. Floating in space, two blocks attract one another and slide frictionlessly along their common long axis, $x$. At the top, both blocks carry conserved mechanical momenta, and clearly $P_{\rm tot}=p_{A}+p_{B}$. At the bottom, the same relation holds, but $p_{A}$ and $p_{B}$ are now pseudomomenta; their sum nevertheless equals the conserved mechanical momentum. The two blocks serve as surrogates for a foreground electron (A) in a perfect background crystal (B), illustrating that both the foreground and the background acquire pseudomomentum in each other’s presence.
Let $\hat{U}(\mathbf{h})$ translate all coordinates, foreground and background, by an arbitrary vector $\mathbf{h}\in\mathbb{R}^{3}$. For an isolated electron+background system,
[…]
so the corresponding Noether generator is the conserved total mechanical momentum, denoted by $\hat{\mathbf{Q}}$.
In the foreground/background formulation $\hat{\mathbf{Q}}$ is obtained as the sum of two subsystem pseudomomenta. Because $\hat{U}(\mathbf{h})$ factorizes into a translation of the foreground coordinate and an equal translation of the background configuration,
$\hat{U}(\mathbf{h})=\hat{U}_{\mathrm{f}}(\mathbf{h})\,\hat{U}_{\mathrm{b}}(\mathbf{h}),$ (4)
we have
$\hat{\mathbf{Q}}=\hat{\mathbf{P}}+\hat{\mathbf{p}}$ (5)
Here $\hat{\mathbf{p}}$ is the pseudomomentum carried by the foreground electron (the generator of its translation), and $\hat{\mathbf{P}}$ is the pseudomomentum carried by the background. (the generator of translating the background as a whole). Their sum $\hat{\mathbf{Q}}=\hat{\mathbf{P}}+\hat{\mathbf{p}}$ is the conserved mechanical momentum of the isolated system. For convenience, we can set it to 0; it never changes and we can now ignore it.
Earlier we remarked that a pseudomomentum cannot be combined with a mechanical momentum. We have now seen, however, that two pseudomomenta can combine to form a mechanical momentum. These rules are analogous to those governing the collision of two objects in free space, except that here the interaction is periodic in their relative coordinate, apart from defects and lattice vibrations.
The total momentum $\hat{\mathbf{Q}}$ is conserved, but the foreground and background momenta, $\hat{\mathbf{p}}$ and $\hat{\mathbf{P}}$, undergo equal and opposite changes, in exact analogy with collisions in free space. This exchange does not require the creation or annihilation of internal lattice excitations and occurs immediately at the collision event (see Fig. 2).
When a crystal containing an internal electron is treated as a closed many-body system, the total background momentum $\mathbf{P}=\sum_{i}\mathbf{p}_{i}$ is defined at each instant as the generator of translations of the background degrees of freedom. Since the foreground is held fixed in this construction, $\mathbf{P}$ is a pseudomomentum. An impulse imparted to a single atom in the background by a foreground object changes the total background momentum immediately by the same amount, resulting in a shift of the center-of-mass velocity $\Delta\mathbf{V}_{\mathrm{CM}}=\Delta\mathbf{p}/M$, independent of any subsequent internal redistribution of momentum.
The finite-speed propagation of stress, phonons, or other internal excitations governs how momentum is shared among individual degrees of freedom, but it does not affect the instantaneous value of the conserved quantity itself. This kinematic distinction follows directly from translational invariance and does not imply nonlocal dynamics or superluminal signaling; rather, it reflects the separation between global conservation laws and the causal evolution of local observables.
Figure 2: The total momentum (black) remains constant as the foreground (electron, blue) and background (lattice, red) momenta jump abruptly at the collision time $t^{*}$. There is no delay in the total lattice momentum change, even though only a single atom initially carries the momentum. While this is elementary, we emphasize it because there is a common tendency to associate momentum flow directly with energy flow in the medium. Indeed there is momentum flow among particles, but the total is fixed immediately. The total appears in the wavefunctions.
Finite-window symmetry
The interaction is periodic only over a finite distance $L$, as required for any physically realizable system. This situation is directly analogous to the classic particle in a box. Momentum is not strictly conserved in such a system; nevertheless, traveling wave packets carrying well-defined momentum can be constructed as linear combinations of standing-wave eigenstates, which individually carry no momentum. Over finite times these wave packets propagate as if momentum were conserved, until a collision with a boundary reverses their motion.
This type of approximate symmetry, valid only over finite spatial and temporal windows, is sometimes implicit in discussions of wave-packet dynamics in finite systems. For clarity, we refer to it here as finite-window symmetry. If one insists on working with finite crystals, this concept must be adopted: translational symmetry holds only within a spatial and temporal window set by the system size.
In a large crystal this construction is an excellent approximation. Localized corpuscles can be formed that propagate nearly as Bloch waves with well-defined pseudomomentum and experience long travel times before encountering a boundary. We emphasize that this corpuscle construction is a calculational convenience that permits the use of the notation and results associated with the formally unphysical Born-von Kármán boundary conditions, without assuming their literal validity.
III Lattice degrees of freedom and zero modes
We consider a finite crystal consisting of atoms with coordinates $\hat{\mathbf{R}}_{j}$ and masses $m_{j}$, with total lattice mass $M=\sum_{j}m_{j}$, minus the foreground electron. We define the center-of-mass coordinate of the lattice background as
$\hat{\mathbf{R}}\equiv\frac{1}{M}\sum_{j}m_{j}\,\hat{\mathbf{R}}_{j}.$ (6)
It is convenient to decompose the atomic coordinates as
$\hat{\mathbf{R}}_{j}$ $=\hat{\mathbf{R}}+\bm{\xi}^{(0)}_{j}+\hat{\mathbf{u}}_{j},$ (7)
$\sum_{j}m_{j}\bm{\xi}^{(0)}_{j}$ $=\mathbf{0},\qquad\sum_{j}m_{j}\hat{\mathbf{u}}_{j}=\mathbf{0},$
where $\hat{\mathbf{R}}$ is the exact lattice center-of-mass coordinate, $\bm{\xi}^{(0)}_{j}$ denote fixed reference positions measured relative to the center of mass, and $\hat{\mathbf{u}}_{j}$ represent internal, dynamical lattice displacements. The mass-weighted constraints ensure that all internal degrees of freedom carry zero net momentum, so that $\hat{\mathbf{R}}$ describes the unique translational zero mode.
The classical and quantum theory of lattice vibrations explicitly separates rigid motion from internal vibrations. Born and Huang showed that the equations of motion of a finite crystal admit three zero-frequency normal modes corresponding to rigid translation of the crystal as a whole, while the internal normal modes satisfy the mass-weighted constraint $\sum_{j}m_{j}\hat{\mathbf{u}}_{j}=\mathbf{0},$ so that internal vibrations carry no net mechanical momentum [8]. All mechanical momentum of the crystal resides in the translational zero mode.
The total center-of-mass degree of freedom of the combined foreground-background system is always conserved and need not be carried explicitly. For an interior electron, it is crucial to track the pseudomomentum of the foreground electron and separately the background,
IV Electron-Lattice interaction
IV.1 Foreground and background Hamiltonians
Given our aim of formulating transport without assuming pre-existing quasiparticles, care is required in defining the Hamiltonian. We write the total Hamiltonian as
$H=H_{\rm fg}+H_{\rm bg}+H_{\rm int},$ (8)
where $H_{\rm fg}$ describes the foreground electron, $H_{\rm bg}$ the background degrees of freedom, and $H_{\rm int}$ their coupling.
For a single foreground electron we take
$H_{\rm fg}=\frac{\hat{\mathbf{p}}^{2}}{2m},$ (9)
where $\hat{\mathbf{p}}$ denotes the generator of translations of the foreground coordinate $\hat{\mathbf{r}}$. The use of a bare kinetic-energy operator reflects the fact that the electron has not been pre-diagonalized into Bloch states. All effects of the periodic crystal potential, including band structure, pseudomomentum exchange, and scattering, are instead generated dynamically through the interaction Hamiltonian $H_{\rm int}$. This is the same starting point adopted by Fröhlich [10].
The remaining electrons of the metal are treated as part of the background. Together with the ions they establish the effective lattice potential, provide screening, and ensure mechanical stability. Their degrees of freedom are therefore implicit in $H_{\rm bg}$.
Density operator and exact factorization
The lattice density operator admits an exact factorization into a center-of-mass translation and an internal part:
$\hat{\rho}_{\mathbf{q}}$ $=\sum_{j}e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}_{j}}=e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{\mathbf{q}},$ (10)
$\hat{\rho}^{\rm int}_{\mathbf{q}}$ $=\sum_{j}e^{-i\mathbf{q}\cdot(\bm{\xi}^{(0)}_{j}+\hat{\mathbf{u}}_{j})}.$
This identity expresses the fact that a uniform translation of the lattice multiplies all density Fourier components by the phase $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$, while all information about internal lattice dynamics is contained in $\hat{\rho}^{\rm int}_{\mathbf{q}}$.
Density–density interaction and pseudomomentum conservation
With this factorization, the electron–lattice interaction may be written in density–density form as
$H_{\rm int}=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}\,\hat{\rho}_{-\mathbf{q}},$ (11)
which follows the standard Born-approximation scattering formalism introduced by Van Hove [34]. Using Eq. (10), this becomes
$H_{\rm int}$ $=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}\,e^{i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{-\mathbf{q}}$ (12)
$=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}\,\hat{\rho}^{\rm int}_{-\mathbf{q}}.$
[…]
Crucially, this exchange of pseudomomentum applies whether or not there is a change a change in phonon occupation numbers. Elastic matrix elements of $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ are generically nonzero. The background zero mode is therefore capable of absorbing finite pseudomomentum with negligible energy cost, providing the kinematic basis for elastic but momentum-relaxing scattering. Any combination of phonon involment and lattice background recoil is allowed. It seems this should matter to measureable properties.
V Elastic internal scattering and the Debye-Waller fraction
In this section we show that, even within a strictly perturbative treatment based on Bloch electrons and phonon number states, robust elastic scattering channels exist already at first order in the electron–lattice coupling. These channels become explicit once the full lattice density operator is retained and the translational zero mode of the lattice is treated as a dynamical degree of freedom. The resulting processes are elastic with respect to internal lattice excitations, despite involving momentum exchange with the lattice as a whole.
Density-density coupling
We consider a single tagged (foreground) electron with coordinate $\hat{\mathbf{r}}$. The interaction between the foreground electron and the background lattice is taken to be of density-density form,
$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,\rho_{\rm fg}(\mathbf{q})\,\rho_{\rm bg}(-\mathbf{q}).$ (13)
Separating the background density into its static equilibrium expectation value and fluctuations, $\rho_{\rm bg}=\langle\rho_{\rm bg}\rangle+\delta\rho_{\rm bg}$, the static component generates a periodic potential acting on the foreground electron,
$U_{\rm Bloch}(\hat{\mathbf{r}})=\sum_{\mathbf{G}}V(\mathbf{G})\,e^{i\mathbf{G}\cdot\hat{\mathbf{r}}}\,\langle\rho_{\rm bg}(-\mathbf{G})\rangle,$ (14)
where $\mathbf{G}$ are reciprocal lattice vectors. The resulting foreground Hamiltonian is therefore
[…]
while the remaining part of $H_{\rm int}$ describes the dynamical coupling to background fluctuations. In this way the familiar Bloch Hamiltonian emerges directly from the same density-density interaction, without being assumed a priori.
The lattice degrees of freedom (the background) are described by a Hamiltonian $H_{\rm bg}$. The coupling between foreground and background is retained in the form
$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,\rho_{\rm fg}(\mathbf{q})\,\rho_{\rm bg}(-\mathbf{q}),$ (16)
where the tagged-electron density operator is
$\rho_{\rm fg}(\mathbf{q})=e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}$, (17)
and the full background density operator is
$\rho_{\rm bg}(-\mathbf{q})=\sum_{j}e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}_{j}}=e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (18)
with $\hat{\mathbf{R}}_{j}=\hat{\mathbf{R}}+\bm{\xi}^{(0)}_{j}+\hat{\mathbf{u}}_{j}$ as defined previously.
Initial and final states
For a perturbative treatment we consider product states of foreground and background,
$\Psi_{i}=\Psi_{n\mathbf{k}}(\mathbf{r})\,\Psi^{\rm bg}_{i}(\{\mathbf{u}_{j}\}),\qquad\Psi_{f}=\Psi_{n^{\prime}\mathbf{k}^{\prime}}(\mathbf{r})\,\Psi^{\rm bg}_{f}(\{\mathbf{u}_{j}\}),$ (19)
where $\Psi^{\rm bg}_{i,f}$ are phonon number states or superpositions thereof. The electronic states are Bloch eigenstates of the static lattice potential.
Using the factorized form of the interaction, any nonvanishing matrix element $\langle\Psi_{f}|H_{\rm int}|\Psi_{i}\rangle$ implies the pseudomomentum transfers
$\Delta\mathbf{p}=+\hbar\mathbf{q},\qquad\Delta\mathbf{P}=-\hbar\mathbf{q},$ (20)
so that the total mechanical momentum of the combined system is conserved. The recoil pseudomomentum is carried by the lattice center-of-mass degree of freedom.
While much stronger coupling regimes can be treated nonperturbatively—for example within wave-on-wave (WoW) simulations—the purpose of the present analysis is to emphasize that even within standard perturbation theory there exist elastic scattering processes that are usually excluded by more restrictive formulations.
Elastic internal channel and the Debye-Waller factor
For a harmonic lattice, the internal displacements admit the normal-mode expansion
$\hat{\mathbf{u}}_{j}=\frac{1}{\sqrt{N}}\sum_{\mathbf{k},\lambda}\sqrt{\frac{\hbar}{2M\omega_{\mathbf{k}\lambda}}}\,\mathbf{e}_{\mathbf{k}\lambda}\left(a_{\mathbf{k}\lambda}e^{i\mathbf{k}\cdot\bm{\xi}^{(0)}_{j}}+a^{\dagger}_{\mathbf{k}\lambda}e^{-i\mathbf{k}\cdot\bm{\xi}^{(0)}_{j}}\right)$. (21)
The operator $e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}$ therefore contains all orders in the phonon operators and automatically generates elastic, quasielastic, and inelastic processes.
In particular, in a phonon number state basis the diagonal matrix elements are generically nonzero,
$\left\langle\{n\}\left|e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\right|\{n\}\right\rangle\neq 0,$ (22)
so that a scattering channel strictly elastic with respect to internal lattice excitations is present already at first order in $H_{\rm int}$. Upon thermal averaging (or in the ground state), this diagonal factor reduces to the Debye-Waller form
$\left\langle e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\right\rangle_{T}=\exp[-W(\mathbf{q},T)],$ (23)
which is the direct analog of the recoil-free (zero-phonon-line) fraction in Mössbauer spectroscopy and in neutron or x-ray scattering.
Relation to the Fröhlich truncation
In the original Fröhlich formulation, the exact factorization Eq. (10) is not employed. The lattice center-of-mass coordinate $\hat{\mathbf{R}}$ is not retained explicitly, which is equivalent to setting the recoil factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$ to unity and treating the lattice as an infinitely massive reference frame. At the same time, the internal displacement exponential is truncated to linear order,
$e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\rightarrow 1-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}.$ (24)
The zeroth-order term reproduces the static periodic potential responsible for Bloch band formation, while the linear term generates single-phonon processes.
Although this procedure is internally consistent and highly successful in the weak-coupling regime, it removes the translational zero mode by construction. As a result, even if the full exponential in $\hat{\mathbf{u}}_{j}$ were retained, the truncated theory would still lack an explicit degree of freedom capable of absorbing the total recoil momentum of the lattice. The standard Fröhlich Hamiltonian therefore describes momentum exchange with internal lattice modes, but not with the lattice as a whole.
Elastic, quasielastic, and inelastic processes
Retaining the full operator $e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}$ in the background density ensures that elastic, quasielastic, and inelastic scattering processes are treated on equal footing already at first order in the density-density interaction. At this level, the interaction transfers a definite total momentum $\hbar\mathbf{q}$ from the foreground electron to the background, independent of how that momentum is subsequently absorbed. It is therefore useful to decompose the momentum transfer as $\mathbf{q}=\mathbf{q}_{\rm el}+\mathbf{q}_{\rm in}$, where $\mathbf{q}_{\rm el}$ corresponds to center-of-mass recoil of the background and $\mathbf{q}_{\rm in}$ is taken up by its internal degrees of freedom. The latter satisfies the exact kinematic relation
$\hbar\mathbf{q}_{\rm in}=-\sum_{\mathbf{k},\lambda}\hbar\mathbf{k}\,\Delta n_{\mathbf{k}\lambda},$ (25)
where $\Delta n_{\mathbf{k}\lambda}$ denotes the change in occupation of the background normal modes. In this formulation, the distinction between elastic and inelastic scattering is determined by the background state rather than by an expansion in ionic displacements.
Thermal averaging and the Debye-Waller fraction
In thermal equilibrium the internal lattice degrees of freedom are described by the density matrix
$\hat{\varrho}=\frac{e^{-\beta\hat{H}_{\mathrm{lat}}}}{Z}=\sum_{n}p_{n}\,|n\rangle\langle n|,\qquad p_{n}=\frac{e^{-\beta E_{n}}}{Z},$ (26)
where $|n\rangle$ denote exact many-body eigenstates of the lattice Hamiltonian $\hat{H}_{\mathrm{lat}}$ with energies $E_{n}$. For elastic internal scattering the relevant quantity is the thermal average of the phonon-diagonal matrix element of the internal density operator,
$\langle\rho^{\rm int}_{\mathbf{q}}\rangle_{T}=\sum_{n}p_{n}\,\langle n|\rho^{\rm int}_{\mathbf{q}}|n\rangle.$ (27)
The corresponding elastic (coherent) scattering intensity is then given by
$I_{\mathrm{el}}(\mathbf{q})=\big|\langle\rho^{\rm int}_{\mathbf{q}}\rangle_{T}\big|^{2}.$ (28)
By contrast, the total scattering probability summed over all final internal states is
$I(\mathbf{q})=\left\langle\rho^{\rm int}_{\mathbf{q}}\rho^{\rm int}_{-\mathbf{q}}\right\rangle_{T}.$ (29)
The elastic internal branching fraction is therefore
$f_{\mathrm{el}}(\mathbf{q},T)=\frac{\sum_{\alpha}p_{n}\,\big|\langle n|\rho^{\rm int}_{\mathbf{q}}|n\rangle\big|^{2}}{\left\langle\rho^{\rm int}_{\mathbf{q}}\rho^{\rm int}_{-\mathbf{q}}\right\rangle_{T}}.$ (30)
For a harmonic lattice one finds[18, 23]
$f_{\mathrm{el}}(\mathbf{q},T)\simeq e^{-2W(\mathbf{q},T)},\qquad W(\mathbf{q},T)=\frac{1}{2}\left\langle(\mathbf{q}\cdot\hat{\mathbf{u}})^{2}\right\rangle_{T},$ (31)
showing that a substantial fraction of momentum-relaxing scattering events can be strictly elastic internally even at finite temperature.
Estimate of the elastic internal fraction at representative momentum transfers
For momentum relaxation in a metal it is natural to consider momentum transfers of order the Fermi momentum, since large-angle scattering dominates the transport relaxation rate through the usual $(1-\cos\theta)$ weighting. For copper, a free-electron estimate yields
$k_{F}\simeq 1.36\ \text{\AA }^{-1},\qquad 2k_{F}\simeq 2.7\ \text{\AA }^{-1}.$ (32)
We therefore evaluate the elastic internal (phonon-diagonal) fraction $f_{\mathrm{el}}(\mathbf{q},T)$ at representative values $q\ll k_{F}$, $q\sim k_{F}$, and $q\sim 2k_{F}$.
Within the isotropic Debye-Waller estimate used above,
$2W(\mathbf{q},T)=\frac{q^{2}\langle u^{2}\rangle_{T}}{3},\qquad f_{\mathrm{el}}(\mathbf{q},T)\simeq e^{-2W(\mathbf{q},T)}.$ (33)
Using the room-temperature mean-square displacement for copper, $\langle u^{2}\rangle_{300\mathrm{K}}\simeq 1.9\times 10^{-2}\ \text{\AA }^{2}$, one finds the values summarized in Table 1.
qq-1) $2W$ $f_{\mathrm{el}}$
$0.3$ $5.7\times 10^{-4}$ $0.9994$
$k_{F}\simeq 1.36$ $1.2\times 10^{-2}$ $0.988$
$2k_{F}\simeq 2.7$ $4.6\times 10^{-2}$ $0.955$
Table 1: Estimated elastic internal fraction $f_{\mathrm{el}}$ for copper at room temperature ($T=300$ K) for representative momentum transfers. Even for large-angle scattering with $q\sim 2k_{F}$, the majority of momentum-relaxing events remain phonon-diagonal, with momentum absorbed predominantly by the lattice background rather than by internal phonon excitation.
These estimates show that internal phonon excitation constitutes only a minor branching channel for momentum relaxation in copper, even at room temperature and for momentum transfers relevant to transport. The dominant mechanism for absorbing electron momentum is thus elastic or quasi-elastic coupling to the lattice background, consistent with the discussion above.
VI Experiments suggesting large elastic contributions
Experimental evidence for predominantly elastic momentum relaxation in metals at low temperature has accumulated for decades, particularly in the mesoscopic-transport and quantum-interference literature. In this section we show that the framework developed here resolves several longstanding disparities between theory and experiment regarding electronic transport. We review a range of experiments that demonstrate a clear separation between momentum relaxation and energy relaxation, consistent with elastic or quasi-elastic scattering dominating momentum randomization.
Hierarchy of momentum and energy relaxation times
Electrical resistivity and optical conductivity probe momentum relaxation, quantum-interference phenomena such as weak localization probe phase coherence, and nonequilibrium measurements including shot noise, Johnson-noise thermometry, and hot-electron relaxation directly access energy exchange. Strikingly, these probes consistently reveal a strong hierarchy of timescales,
$\tau_{\mathrm{p}}\ll\tau_{E},$ (34)
established using independent experimental techniques and across a wide range of materials. Here $\tau_{\mathrm{p}}$ is the timescale on which electronic momentum (or pseudomomentum) is randomized by scattering, while $\tau_{E}$ is the timescale over which electrons exchange energy with the background and relax toward thermal equilibrium.
The consistent appearance of this hierarchy demonstrates that momentum relaxation and energy relaxation are distinct physical processes. While conventional language often associates momentum randomization with inelastic scattering, these experiments show that electronic momentum can be efficiently scrambled by a time-dependent environment without significant energy exchange. The lattice need not act as an energy sink in order to act as a momentum sink. This separation is precisely what is expected when electrons scatter elastically from a dynamically fluctuating lattice background.
Quantum interference and dephasing
Traditionally, the suppression of quantum-interference effects with increasing temperature has been attributed to inelastic scattering processes that transfer energy to the environment [22, 2]. However, it is important to distinguish carefully between dephasing, as operationally defined in transport experiments, and irreversible decoherence in the strict quantum-mechanical sense.
A wide range of experiments demonstrate that genuine energy relaxation remains slow over the same temperature range in which phase coherence is strongly suppressed. In particular, mesoscopic interference experiments show a rapid decrease—and in some cases saturation—of the phase coherence time $\tau_{\phi}$ at low temperatures, even as independent probes indicate that electron-phonon energy exchange rates continue to vanish [26, 16, 35]. This separation of timescales shows that momentum and phase randomization need not be accompanied by significant energy transfer.
As observed in weak-localization experiments on mesoscopic metals, the loss of quantum interference reflects the destruction of coherent backscattering between time-reversed trajectories. This coherence is far more sensitive than the transport relaxation time and can be suppressed by elastic, time-dependent scattering with negligible energy exchange and without irreversible decoherence [22, 26].
From this perspective, dephasing reflects the extreme sensitivity of quantum interference to weak, time-dependent fluctuations of the environment. Even in the absence of thermal excitations, lattice coordinates remain quantum operators with nonvanishing variance due to zero-point motion. As a result, elastic scattering amplitudes are not unity even at $T=0$, as expressed by the Debye-Waller factor. When such quantum fluctuations are slow on electronic timescales—as occurs for long-wavelength modes in finite samples—they act as an effectively time-dependent scattering potential, suppressing interference under ensemble and time averaging. These processes are elastic on electronic energy scales and do not constitute irreversible decoherence.
Hot-electron relaxation
A particularly direct separation of momentum relaxation from energy relaxation is provided by hot-electron experiments in metals [36, 30, 11, 12]. In these measurements, nonequilibrium electron distributions are injected and their subsequent momentum randomization and energy loss are probed independently. Over wide temperature ranges, the momentum-relaxation time inferred from transport or magnetotransport is found to be much shorter than the energy-relaxation time associated with thermalization. Electrons therefore undergo many momentum-deflecting collisions while remaining essentially isoenergetic. Such a pronounced separation of timescales cannot be explained by purely inelastic phonon scattering and instead points to an effectively elastic deflection channel intrinsic to the metallic state.
Johnson-noise thermometry
Johnson-noise thermometry provides a particularly stringent probe of energy relaxation in driven metallic microstructures [36, 30]. In these experiments, a dc bias drives a substantial current while the local electronic temperature is inferred directly from the Johnson-Nyquist noise. Because the noise power is proportional to the electronic energy distribution, this technique provides a quantitative and noninvasive measure of energy flow, independent of momentum transport.
A key observation is that under bias the electrical resistivity and momentum randomization rate increase strongly, while the electronic temperature remains comparatively low and spatially nonuniform. This demonstrates that electrons can undergo many momentum-deflecting scattering events without efficiently transferring energy to the lattice. Such behavior is difficult to reconcile with a picture in which inelastic phonon scattering controls momentum relaxation, but follows naturally if the dominant scattering mechanism is elastic deflection from a dynamically fluctuating lattice background.
VII Discussion
The purpose of this section is twofold. First, we examine the consistency of the present framework with well-established experimental facts, including the Wiedemann-Franz law, weak localization, and quantum oscillations. These topics are not speculative; they address potential objections that might otherwise arise when elastic scattering plays a dominant microscopic role. Second, we outline broader interpretive implications for Planckian transport, linear-in-$T$ resistivity, and the Mott-Ioffe-Regel crossover. These latter points are necessarily more exploratory and are presented as a coherent physical interpretation rather than as definitive conclusions.
VII.1 Implications for the Lorenz ratio
Because elastic scattering randomizes momentum just as effectively as inelastic scattering, no contradiction arises between predominantly elastic microscopic dynamics and the well-established success of conventional transport theory.
In a degenerate metal, the Wiedemann-Franz (WF) law states that the ratio of the electronic thermal conductivity $\kappa$ to the electrical conductivity $\sigma$ times temperature approaches the universal Sommerfeld value
$L\equiv\frac{\kappa}{\sigma T}\;\xrightarrow[T\to 0]{}\;L_{0}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}.$ (35)
While some fraction of inelastic processes must exist to relax the electronic energy current and establish global thermal equilibrium, it has long been understood that these processes need not control the transport coefficients themselves [39, 21, 31].
If the dominant momentum-relaxing channel is elastic and only weakly dependent on energy over the thermal window $\sim k_{B}T$ about the Fermi level, then both electrical and thermal currents are governed by the same relaxation time $\tau_{\mathrm{el}}$,
$\sigma\propto\tau_{\mathrm{el}},\qquad\kappa\propto\tau_{\mathrm{el}},$ (36)
and the Sommerfeld value $L_{0}$ is recovered. Rare inelastic events, with rate $1/\tau_{\mathrm{in}}\ll 1/\tau_{\mathrm{el}}$, serve to transfer energy from the electronic system to internal lattice degrees of freedom, but do not enter the leading expressions for $\sigma$ or $\kappa$. A transport regime dominated by elastic momentum randomization is therefore fully compatible with the WF law, even when true energy relaxation is parametrically slow.
VII.2 Weak localization
Weak localization (WL) provides a particularly clear illustration of the dominance of elastic backscattering in metals at low temperature. The phenomenon arises from constructive interference between pairs of time-reversed electronic trajectories that return to the same point after multiple elastic scattering events [6, 22]. The very existence of a WL correction therefore implies that the dominant backscattering processes are coherent and elastic.
If backscattering were predominantly inelastic, interference between time-reversed paths would be washed out and no WL correction could develop. Instead, experiments show a well-defined suppression of the conductivity at low temperature together with a characteristic negative magnetoresistance when time-reversal symmetry is broken by an applied magnetic field. As temperature is increased, WL is gradually suppressed. Although this suppression is often attributed to “inelastic scattering,” experimentally it reflects the loss of phase coherence rather than rapid energy relaxation.
Independent probes such as hot-electron relaxation and Johnson-noise thermometry show that electron-phonon energy exchange remains slow over the same temperature range. The suppression of WL therefore signals the onset of a time-dependent scattering environment that randomizes phase while leaving electronic energy essentially unchanged. Within the present framework, WL fits naturally into the hierarchy of timescales
$\tau_{p}\ll\tau_{\phi}\ll\tau_{E},$
with elastic scattering producing diffusive transport and slower temporal evolution of the lattice background destroying phase coherence without necessitating phonon creation or significant energy dissipation.
VII.3 Shubnikov-de Haas oscillations
Quantum oscillation phenomena such as the Shubnikov-de Haas and de Haas-van Alphen effects provide an independent and equally stringent demonstration that electronic scattering in metals is predominantly elastic at low temperature [33, 3]. These oscillations arise from the quantization of electronic orbits in a magnetic field and reflect coherent motion around the Fermi surface over many cyclotron periods.
The oscillation amplitude is controlled by the Dingle factor [9],
$\exp\!\left(-\frac{\pi}{\omega_{c}\tau_{q}}\right),$ (37)
where $\tau_{q}$ is the quantum lifetime associated with Landau-level broadening. Importantly, $\tau_{q}$ is primarily sensitive to elastic momentum-randomizing processes, while genuinely inelastic energy exchange rapidly suppresses oscillations. Any scattering event that transfers an energy comparable to $\hbar\omega_{c}$ destroys the coherence required to complete a cyclotron orbit.
The observation of quantum oscillations over broad temperature ranges and down to low magnetic fields therefore implies that electrons undergo many momentum-deflecting collisions without significant energy exchange. As in the case of weak localization, the temperature dependence of the oscillation amplitude reflects phase smearing due to a time-dependent scattering environment rather than rapid inelastic backscattering. Quantum oscillations thus provide direct experimental evidence that elastic scattering dominates electron dynamics in normal metals at low temperature.
VII.4 Sound attenuation
Ultrasonic attenuation in metals provides a complementary probe of electron-lattice dynamics from the lattice side. Experimentally, sound waves propagating through a metal are strongly attenuated even in very clean samples, and this attenuation is often interpreted as evidence for inelastic electron-phonon scattering [28, 25]. However, attenuation of a coherent acoustic wave does not, by itself, imply irreversible energy dissipation.
From a microscopic perspective, attenuation measures the loss of coherence of a propagating phonon mode. This loss can arise through elastic scattering processes that redistribute acoustic momentum among phonon modes or transfer momentum to the electronic subsystem, even when the total energy of the combined electron-lattice system is conserved. In this sense, sound attenuation is directly analogous to the suppression of quantum interference in electronic transport: coherent propagation is degraded without requiring strong energy relaxation [19, 1].
Standard theories of ultrasonic attenuation distinguish regimes based on the relative timescales of the sound frequency and electronic or phononic relaxation rates. In the collisionless (Landau-Rumer) regime, attenuation arises from scattering of sound by thermal excitations and mode conversion rather than from viscous energy dissipation [20, 25]. More generally, the acoustic wave loses forward coherence through elastic momentum exchange with a dynamically fluctuating background, while energy relaxation enters only at longer timescales.
Within the present framework, elastic momentum exchange between electrons and the lattice background provides a natural mechanism for ultrasonic attenuation. Time-dependent lattice distortions scatter sound waves, redirecting acoustic momentum and reducing the amplitude of the forward-propagating mode, while leaving electronic and lattice energies largely unchanged. Energy dissipation occurs subsequently through rare inelastic processes that establish thermal equilibrium.
This interpretation is consistent with experimental observations in normal metals, where substantial ultrasonic attenuation persists even in temperature regimes where independent measurements show that electron-phonon energy transfer is strongly suppressed [28, 25]. As in weak localization and quantum oscillation phenomena, sound attenuation therefore supports a picture in which elastic scattering dominates momentum transfer, while true energy relaxation plays a secondary role.
VII.5 Planckian diffusion and linear-in-$T$ resistivity
Much of the recent literature introduces Planckian behavior as a bound on a microscopic relaxation time $\tau_{\mathrm{Pl}}\sim\hbar/k_{B}T$, often interpreted as the fastest possible rate of inelastic scattering or energy dissipation [37, 14, 13, 32]. Within this narrative, transport coefficients are obtained by inserting $\tau_{\mathrm{Pl}}$ into kinetic or hydrodynamic expressions.
The framework developed in Ref. [38] and here suggests a different interpretation. The primary microscopic result is the emergence of real-space diffusion with a diffusion constant
$D\sim\frac{\hbar}{m^{\ast}},$ (38)
arising from elastic momentum scrambling in a time-dependent lattice background. This form of Planckian diffusion does not rely on quasiparticles, strong inelastic scattering, or rapid energy dissipation; it follows directly from quantum dynamics in a fluctuating potential and has been demonstrated explicitly in wave-on-wave simulations of the full electron-lattice problem [5, 38].
Once diffusion is established, Einstein relations connect $D$ to transport coefficients,
$\sigma=\chi D,\qquad D=\mu\frac{k_{B}T}{e},$ (39)
so that recasting $D$ in terms of an effective relaxation time yields
$\tau_{\mathrm{eff}}\sim\frac{\hbar}{k_{B}T},$ (40)
up to factors of order unity. From this perspective, the Planckian timescale is not a fundamental microscopic bound, but an emergent parametrization of quantum-limited diffusion combined with equilibrium thermodynamics. It constrains how rapidly charge spreads in space, not how rapidly energy must be irreversibly transferred to internal degrees of freedom.
This viewpoint also clarifies the emergence of linear-in-$T$ resistivity across and below the Debye scale. Within a conventional Bloch-Grüneisen picture, the Debye temperature sets the scale at which all phonon modes become available for inelastic scattering. In contrast, within the quantum acoustics framework there is no special anomaly associated with crossing the Debye temperature. Linear-in-$T$ resistivity arises smoothly as coherence is lost in a time-dependent elastic background, while energy relaxation remains comparatively slow. The Debye scale therefore marks neither the onset of momentum relaxation nor a transition to a new inelastic regime, but rather a coherence crossover.
VII.6 Mott-Ioffe-Regel crossover as loss of quasiparticle coherence
The Mott-Ioffe-Regel (MIR) condition $k_{F}\ell\sim 1$ is often interpreted as a boundary beyond which quasiparticles “cease to exist” and transport must become anomalous. From the present perspective, this interpretation is misleading. The MIR condition marks the point at which a semiclassical mean free path loses a literal geometric meaning, not a dynamical singularity in transport.
Within the quantum acoustics framework, transport evolves smoothly through $k_{F}\ell\sim 1$. Diffusion persists, but it is no longer usefully described in terms of well-defined trajectories between discrete scattering events. Crucially, this breakdown of a quasiparticle picture does not coincide with a loss of coherence. Were strongly inelastic processes truly dominant at the MIR scale, interference phenomena such as weak localization and quantum oscillations would be rapidly suppressed, contrary to experimental observation. Instead, experiments show that momentum relaxation can be strong while inelastic energy relaxation remains weak.
Wave-on-wave simulations sharpen this picture by explicitly demonstrating that momentum can be transferred elastically to a thermally distorted lattice background while preserving electronic energy. The MIR crossover therefore reflects a loss of quasiparticle coherence rather than the onset of strong dissipation, fully consistent with elastic-dominated momentum relaxation in a dynamical lattice environment.
VIII Summary
Summary
In the conventional clamped-lattice formulation of electron-phonon scattering, the factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$ associated with rigid translation of the lattice background is removed by assumption. This elimination suppresses an otherwise allowed elastic momentum-transfer channel, forcing momentum relaxation to be identified exclusively with inelastic phonon creation or annihilation. By restoring the lattice translational zero mode, we show that this identification is not required by symmetry, conservation laws, or kinematics. In a finite crystal, elastic pseudomomentum transfer between an electron and the lattice background is generically present and can dominate momentum relaxation at low and intermediate temperatures.
Making this elastic channel explicit requires abandoning several silent but consequential restrictions embedded in the traditional formulation. First, the lattice must be treated as a finite, isolated object so that its center-of-mass degree of freedom remains dynamical. (If the crystal is attached to a laboratory or substrate, this simply enlarges the background without altering the underlying momentum bookkeeping.) Second, discrete translational symmetry must be implemented symmetrically, with foreground and background pseudomomenta treated as equally real and conserved quantities. Third, the electron-lattice interaction must be retained in its full density-density form, rather than truncated in a way that projects out elastic momentum-transfer processes by construction. When these elements are retained, elastic scattering of electrons from thermal lattice distortions is not merely allowed but unavoidable.
In this framework, pseudomomentum is exchanged between the electron and the lattice background, while total mechanical momentum is conserved through recoil of the lattice zero mode, without the necessity of phonon creation. The long-standing identification of lattice motion with inelasticity is thus revealed to be an artifact of an incomplete theoretical description.
Restoring elastic momentum-transfer processes to their proper role provides a more complete and internally consistent foundation for electronic transport in crystals. It offers a natural interpretation of experimental observations that indicate robust elastic scattering in clean materials and clarifies how momentum relaxation can be efficient even when energy relaxation is slow. More broadly, these results raise the possibility that in at least some strange-metal regimes, momentum relaxation is dominated by elastic scattering in a time-dependent background, with genuinely inelastic processes playing a secondary role in establishing energy equilibration rather than setting the transport rate.
IX Acknowledgements
E. Heller thanks B. Halperin for many discussions that provided a thorough understanding of the traditional theory. He also thanks Prof. Martina Hentschel for insightful comments that have improved the manuscript. This work was supported by the U.S. Department of Energy, Office of Science, under Grant No. DE-SC0025489.
Appendix A Elastic and inelastic scattering in a tight-binding, isolated one-dimensional chain
We consider a finite one-dimensional tight-binding chain of $N$ identical atoms, isolated in space, with nearest-neighbor springs and an electron moving in a basis of localized $p$ orbitals. The chain possesses an exact translational zero mode, corresponding to its center-of-mass (CM) motion. This allows the electron to exchange momentum with the lattice as a whole without necessarily exciting internal phonon modes. As a result, at finite temperature there exists a strictly elastic backscattering channel for an electronic wavepacket.
Hamiltonian and separation of the CM mode
Let the atomic coordinates be written as
$x_{n}=R+na+u_{n},\qquad\sum_{n=1}^{N}u_{n}=0,$ (41)
where $R$ is the center-of-mass coordinate, $a$ the equilibrium lattice spacing, and $\{u_{n}\}$ the internal displacements. The lattice Hamiltonian separates exactly into CM and internal parts,
$H_{\rm lat}=\frac{P_{R}^{2}}{2M_{\rm tot}}+\sum_{n=1}^{N}\frac{p_{n}^{2}}{2M}+\frac{K}{2}\sum_{n=1}^{N-1}(u_{n+1}-u_{n})^{2},$ (42)
with total mass $M_{\rm tot}=NM$ and $\sum_{n}p_{n}=0$.
The electron is described by a nearest-neighbor tight-binding Hamiltonian,
$H_{e}=-\sum_{n=1}^{N-1}t_{n}\left(c_{n+1}^{\dagger}c_{n}+c_{n}^{\dagger}c_{n+1}\right),$ (43)
where the hopping amplitudes depend linearly on the bond lengths,
$t_{n}\equiv t(x_{n+1}-x_{n})\simeq t_{0}-\alpha(u_{n+1}-u_{n}).$ (44)
The electron-lattice coupling depends only on internal coordinates; the CM coordinate $R$ drops out exactly.
Strictly elastic backscattering via CM recoil
Because the full electron-lattice system is translationally invariant, there exists an exactly conserved total momentum. The electronic band energy satisfies $\varepsilon(k)=\varepsilon(-k)$, so reversing the electronic momentum does not change the electronic energy.
Consider an incoming electronic wavepacket narrowly peaked at momentum +k+k, and an internal lattice eigenstate $|\nu\rangle$. A scattering process of the form
$|k\rangle\otimes|\nu\rangle\otimes|P_{R}\rangle\;\longrightarrow\;|-k\rangle\otimes|\nu\rangle\otimes|P_{R}+2\hbar k\rangle$ (45)
leaves the internal phonon state unchanged and transfers the momentum $2\hbar k$ entirely to the CM of the chain.
The associated change in CM kinetic energy is
$\Delta E_{\rm CM}=\frac{(P_{R}+2\hbar k)^{2}-P_{R}^{2}}{2M_{\rm tot}}=\frac{4\hbar k\,P_{R}+4\hbar^{2}k^{2}}{2M_{\rm tot}}.$ (46)
Exact energy conservation requires $\Delta E_{\rm CM}=0$, which is satisfied when
$P_{R}=-\hbar k.$ (47)
For this value of the CM momentum, the backscattering process $k\rightarrow-k$ is strictly elastic, involving no excitation or absorption of internal phonon modes. Alternately, if $N$ is very large, the requirement $P_{R}=-\hbar k$ can be relaxed. For a finite, free chain of total mass $M$, electron backscattering $k\rightarrow-k$ changes the electron momentum by $\Delta p_{e}=-2\hbar k$. Momentum conservation then requires the chain to absorb a recoil $\Delta P_{\rm chain}=+2\hbar k$. If the chain initially has CM momentum $P$, energy conservation implies an associated recoil energy
$\Delta E_{\rm recoil}=\frac{(P+2\hbar k)^{2}-P^{2}}{2M}=\frac{2\hbar k\,(P+\hbar k)}{M}.$ (48)
Strictly elastic backscattering without internal excitation is therefore possible only for the special value $P=-\hbar k$; for generic initial CM momentum, an energy mismatch must be accommodated.
For a macroscopic chain, however, $M\propto N$ and the recoil energy scales as $\Delta E_{\rm recoil}\sim 1/N$, becoming parametrically small in the thermodynamic limit. In this regime the energy mismatch can be absorbed without resolving a discrete internal excitation. In particular, an electron wavepacket of spatial extent $L$ has an intrinsic energy spread
$\Delta E_{\rm pkt}\sim\hbar v_{g}\,\Delta k\sim\hbar v_{g}/L,$ (49)
which is independent of $N$ and typically exceeds $\Delta E_{\rm recoil}$ for sufficiently large chains. Equivalently, if the scattering occurs over a finite time $\tau_{\rm scatt}$, the associated time-energy uncertainty $\hbar/\tau_{\rm scatt}$ exceeds the recoil energy, rendering the process effectively elastic from the electron’s perspective.
In a phonon number-state description, the accommodation of this small energy mismatch corresponds to the creation or annihilation of one or more extremely low-frequency acoustic phonons. Importantly, however, such number-state changes do not account for the electron’s momentum reversal $k\rightarrow-k$: a single low-$q$ phonon carries momentum $\hbar q\ll 2\hbar k$. The dominant momentum transfer is instead absorbed by the collective motion of the chain (center-of-mass recoil), while the phonon sector merely provides an energetically soft channel required by exact energy conservation.
Thus, for large $N$, electron backscattering involves an exact momentum transfer of $2\hbar k$ to the background with a vanishing energy cost, and any phonon number-state changes reflect quasi-elastic energy bookkeeping rather than the mechanism of momentum reversal itself.
Role of finite temperature
At finite temperature $T$, the CM momentum is thermally distributed,
$f(P_{R})\propto\exp\!\left(-\frac{P_{R}^{2}}{2M_{\rm tot}k_{B}T}\right),$ (50)
with width $\sigma_{P}=\sqrt{M_{\rm tot}k_{B}T}$. The resonant value $P_{R}=-\hbar k$ therefore occurs with nonzero probability density for any $T>0$. Consequently, strictly elastic electronic backscattering events exist at finite temperature, mediated entirely by recoil of the lattice center of mass.
Bulk and boundary backscattering in a isolated chain
The elastic CM-recoil backscattering channel does not require the electron to encounter the ends of the chain and may occur locally in the bulk. The electron-lattice interaction in Eq. (44) depends only on nearest-neighbor bond distortions and is independent of absolute position.
For a wavepacket localized well away from the ends, with spatial extent $\ell\ll Na$, the matrix element connecting $|k\rangle$ and $|-k\rangle$ takes the form
$\langle-k,P_{R}+2\hbar k|H_{\rm int}|k,P_{R}\rangle\propto\sum_{n}e^{2ikna}\,\langle P_{R}+2\hbar k|P_{R}\rangle,$ (51)
where the sum runs over bulk sites only. The momentum transfer is absorbed entirely by the CM degree of freedom, while the internal lattice state remains unchanged.
Free boundaries of the finite chain can also reflect an electronic wavepacket once it reaches an end. Such boundary reflections are elastic and likewise involve a momentum reversal $k\rightarrow-k$ accompanied by CM recoil. Free boundaries therefore provide a spatial location for reflection but do not act as a momentum sink. Both bulk and boundary backscattering processes are governed by the same momentum conservation principle enforced by the CM zero mode.
Other scattering channels
In addition to the strictly elastic CM recoil channel, a variety of inelastic and hybrid scattering processes are allowed within a isolated chain.
Single-phonon emission and absorption
If momentum is transferred to internal phonon modes, the internal lattice state changes. A generic single-phonon process takes the form
$|k\rangle\otimes|\{n_{q}\}\rangle\longrightarrow|k^{\prime}\rangle\otimes|\{n_{q}^{\prime}\}\rangle,$
with $n_{q}^{\prime}=n_{q}\pm 1$. Energy conservation requires
$\varepsilon(k^{\prime})=\varepsilon(k)\pm\hbar\omega_{q}+\Delta E_{\rm CM}.$
The corresponding Golden-Rule rate is
$W_{k\to k^{\prime};q}=\frac{2\pi}{\hbar}\,|M_{k,k^{\prime};q}|^{2}\,\delta\!\left(\varepsilon(k^{\prime})-\varepsilon(k)\mp\hbar\omega_{q}-\Delta E_{\rm CM}\right),$
with matrix elements scaling as
$M_{k,k^{\prime};q}\sim\alpha\sqrt{\frac{\hbar}{2M\omega_{q}}}\,F_{k,k^{\prime};q}.$
Multi-phonon and hybrid CM-phonon processes
When single-phonon processes are kinematically suppressed, higher-order multi-phonon processes and hybrid processes involving both CM recoil and internal phonon exchange are allowed. These channels relax the strict resonance condition Eq. (47) at the cost of reduced matrix elements.
Kinematic summary and temperature dependence
  • Elastic channels: exact backscattering $k\rightarrow-k$ without phonon excitation occurs via CM recoil and is enabled by translational invariance.
  • Inelastic channels: internal phonon emission and absorption become increasingly important as temperature increases and as phonon spectra permit energy matching.
  • Finite size: the discreteness of the phonon spectrum in a finite chain strongly constrains available inelastic channels and can enhance the relative importance of CM-mediated elastic processes.
Practical consequences
  • In wavepacket simulations, elastic backscattering events are identified by momentum reversal without change in internal phonon occupation, accompanied by a compensating shift in CM momentum.
  • Both bulk and boundary reflections obey the same momentum conservation law and are distinguished only by their spatial location.
  • The CM zero mode plays a central dynamical role and cannot be neglected without qualitatively altering the scattering physics.

### v2

@@
 Abstract
-In conventional treatments of electron transport, momentum relaxation within a perfect, defect-free crystal is commonly assumed to require phonon creation or annihilation.
-Here we treat the crystal as finite and isolated, retaining the lattice center-of-mass (recoil) degree of freedom and enforcing conservation of total mechanical momentum alongside discrete crystal pseudomomentum.
-Starting from the density-density form of the electron-lattice interaction, we find that an electron in the interior of a perfect crystal admits strong, and in some regimes dominant, elastic momentum-relaxing scattering channels, in which momentum is conserved by recoil of the lattice background without phonon excitation.
-In addition, we identify mixed quasielastic and superelastic channels in which phonon occupations change but do not account fully for the electron’s momentum transfer.
-These results provide a microscopic basis for momentum relaxation that does not rely on local energy dissipation.
-They naturally reconcile a wide range of experimental observations—including weak localization, quantum oscillations, ultrasonic attenuation, and the separation of momentum and energy relaxation times—with predominantly elastic scattering in clean crystals.
-The framework clarifies how diffusive transport, including Planckian-scale diffusion, can emerge from elastic dynamics in a time-dependent lattice background.
+In standard treatments of electron transport, momentum relaxation in a perfect, defect-free crystal is linked with phonon creation or annihilation.
+In this work, we reconsider this problem for a finite, isolated crystal, retaining the lattice center-of-mass (recoil) degree of freedom and enforcing conservation of total mechanical momentum together with discrete crystal pseudomomentum.
+Starting from the density–density form of the electron–lattice interaction, we show that an electron in the interior of a perfect crystal admits elastic momentum-transfer channels in which total momentum is conserved by recoil of the lattice background without phonon excitation.
+These elastic channels can provide the leading contribution to momentum relaxation.
+We further identify mixed quasielastic and superelastic processes in which phonon occupations change but do not account entirely for the electron’s momentum transfer.
+The elastic channels arise within the standard microscopic Hamiltonian and do not require additional disorder or defects.
+The resulting framework provides a complementary microscopic perspective on momentum relaxation in clean crystals, and is consistent with experimental phenomena such as weak localization, quantum oscillations, ultrasonic attenuation, and the observed separation of momentum and energy relaxation times.
 preprint: APS/123-QED
 I Introduction
-Electron-lattice interactions are among the most important processes in condensed matter physics.
-They govern momentum relaxation, energy dissipation, and the loss of quantum coherence that underlies electrical resistance.
-In conventional transport theory, electron momentum relaxation in a perfect, defect-free crystal is commonly identified with inelastic scattering processes in which phonons are created or annihilated.
-Within this view, deflection of an electron by thermal lattice motion is inseparable from energy transfer to internal vibrational degrees of freedom.
-This perspective is codified in the Bloch-Grüneisen and Boltzmann approaches to resistivity, which treat electronic transport as an incoherent sequence of scattering events and assume no electronic phase coherence beyond that retained in lowest-order scattering matrix elements.
-It is further reinforced by the Fröhlich interaction [10], which couples electrons explicitly to phonon creation and annihilation operators and therefore enforces inelasticity at first order in perturbation theory.
-Within this canonical framework, thermal deflection of electrons in a perfect crystal is taken to require phonon excitation, while elastic scattering is associated primarily with defects or impurities [39, 4].
-Here we present evidence that this identification is overly restrictive.
-We show that thermally induced electron scattering in a pure crystal is often largely elastic when viewed on an event-by-event basis.
-Elastic and quasi-elastic momentum-relaxing channels exist even in a perfectly periodic lattice, without the necessity of phonon creation or annihilation.
-Previously, elastic scattering was almost exclusively associated with explicit disorder; we demonstrate that this association is not required by symmetry, conservation laws, or kinematics.
-Early discussions of electronic transport already contained hints of this distinction.
-Peierls emphasized that electrical resistance requires degradation of electronic (crystal) momentum rather than dissipation of electronic energy, and noted explicitly that the lattice, as a massive collective object, can absorb momentum without requiring significant energy transfer [27].
-Pippard similarly argued that defects and impurities are important but not indispensable, and that momentum relaxation could, in principle, proceed through elastic or quasi-elastic mechanisms even in very pure crystals [29].
-These insights acknowledged the logical separation between momentum relaxation and energy dissipation, although this separation was not consistently maintained in subsequent formulations of transport theory.
-In contrast, the canonical Fröhlich Hamiltonian and Bloch-Grüneisen framework [7] implicitly exclude rigid translation of the crystal as a whole from the dynamical description.
-The lattice center-of-mass (recoil) degree of freedom is effectively clamped, and momentum transfer is assigned exclusively to internal phonon excitations or to Umklapp processes.
-Once these assumptions are imposed, momentum loss is forced to coincide with energy dissipation, and elastic deflection by a perfect lattice is excluded by construction.
-We revisit this assumption by treating the crystal as finite, isolated, and freely floating, and by explicitly retaining its total mechanical momentum.
-Enforcing conservation of total momentum more completely cannot be formally incorrect; at worst it may prove quantitatively unimportant.
-As we show below, however, restoring the lattice center-of-mass degree of freedom qualitatively alters the available scattering channels for an electron deep inside a pure conducting crystal.
-In particular, it reveals an elastic recoil channel that is absent by design in the conventional formulation.
-The physical content of this channel is elementary.
-If a particle embedded in a composite object receives a sudden impulse, the momentum of the whole must change immediately, even if the internal degrees of freedom remain largely unexcited.
-A familiar example is provided by Mössbauer scattering, where momentum is absorbed predominantly by recoil of the entire lattice while phonon excitation is strongly suppressed.
-As we shall show, an analogous mechanism operates for electrons interacting with a thermally fluctuating but otherwise perfect crystal.
-The ambiguity of elastic scattering
-Elastic scattering is most familiar in the context of impurity or defect scattering.
-In that setting, momentum conservation is usually justified in one of two ways.
-First, the presence of an impurity breaks discrete translational symmetry, so crystal pseudomomentum need not be conserved.
-Second, within a particle picture, one argues that if an electron enters with one momentum and exits with another, the atomic lattice must recoil so as to conserve total mechanical momentum.
-Both statements are essentially correct.
-However, neither addresses the internal bookkeeping problem that arises for an electron scattering deep inside a perfectly periodic crystal: how does a change in the electron’s pseudomomentum communicate with the mechanical momentum of the lattice as a whole?
-This unresolved point lies at the heart of the transport problem in a translationally invariant system.
-As we show below, it is resolved once three elements are made explicit: (i) the lattice is treated as a fully dynamical object rather than a fixed reference, (ii) pseudomomentum is assigned not only to the electron but also to the crystal background in the absence of the electron density (see the next section), and (iii) the electron-lattice interaction is formulated in density-density form [24, 10], which makes the exchange of pseudomomentum between electron and lattice explicit even in a perfectly periodic crystal.
-In any untethered collection of atoms, an impulse imparted to even a single constituent instantaneously changes the total momentum of the assembly.
-While the subsequent redistribution of that momentum among internal degrees of freedom proceeds causally and at finite velocity, the change in the collective momentum itself is immediate.
-This total momentum is not an abstract bookkeeping device but a genuine dynamical variable that appears explicitly in the many-body wavefunction.
-In the present formulation, this momentum conservation is enforced transparently.
-As shown in Eq.
-12, a momentum transfer $\mathbf{q}$ to the foreground electron is accompanied by an equal and opposite momentum $\mathbf{q}$ carried by the remaining degrees of freedom, which we identify as the lattice background.
-This statement holds independently of whether phonons are created or destroyed in the process.
-Finally, we clarify our use of the term elastic.
+Electron–lattice interactions play a central role in condensed matter physics.
+They govern both momentum relaxation and energy exchange in crystalline solids, and therefore underlie electrical resistivity and the loss of electronic coherence.
+Experimentally, however, momentum and energy relaxation often occur on markedly different timescales.
+In clean metals over broad temperature ranges, electrical resistivity and magnetotransport measurements indicate rapid momentum randomization, while hot-electron experiments, Johnson-noise thermometry, and related nonequilibrium probes show that energy relaxation to the lattice can remain comparatively slow.
+The resulting hierarchy
+$\tau_{p}\ll\tau_{E},$ (1)
+where $\tau_{p}$ is the momentum-relaxation time and $\tau_{E}$ the energy-relaxation time, indicates that electronic momentum can be efficiently scrambled without substantial net energy transfer to internal lattice excitations.
+In conventional transport theory, momentum relaxation in a perfect, defect-free crystal is typically described in terms of phonon-mediated scattering processes.
+Within the Bloch–Grüneisen and semiclassical Boltzmann frameworks, transport is treated as a sequence of transitions between electronic states, with rates evaluated using the electron–phonon interaction [35, 3].
+In the standard Fröhlich Hamiltonian [9], the interaction is expressed explicitly in terms of phonon creation and annihilation operators, and momentum transfer is therefore associated, at lowest order, with changes in phonon occupation.
+In this commonly employed description, elastic momentum relaxation is primarily attributed to impurities or other forms of explicit disorder.
+The observed separation between $\tau_{p}$ and $\tau_{E}$ motivates a careful examination of how momentum conservation is implemented microscopically in a perfectly periodic crystal.
+The conceptual distinction between momentum relaxation and energy dissipation has long been recognized.
+Peierls emphasized that electrical resistance requires degradation of electronic (crystal) momentum rather than dissipation of electronic energy, and noted that the lattice, as a massive collective object, can absorb momentum with minimal energy cost [24].
+Pippard similarly stressed that while defects often play an important role, momentum relaxation need not be uniquely tied to strongly inelastic processes [25].
+In this work we revisit the problem starting from the same microscopic electron–lattice Hamiltonian, but retaining explicitly the lattice center-of-mass degree of freedom.
+In conventional treatments the crystal is effectively considered as a fixed reference frame, and the translational zero mode does not appear among the dynamical variables.
+Here we adopt an expanded viewpoint in which the crystal total mechanical momentum is retained.
+This does not modify the underlying Hamiltonian; rather, it makes explicit a degree of freedom that is commonly suppressed.
+When total momentum conservation is enforced in this manner, a transfer of momentum $\mathbf{q}$ to an electron is accompanied by an equal and opposite momentum $\mathbf{q}$ carried by the remaining degrees of freedom, which we identify as the lattice background.
+This exchange can occur without excitation of internal phonon modes, leading to an elastic recoil channel even in a perfectly periodic crystal.
+We also identify mixed quasielastic and superelastic processes in which phonon occupations change but do not fully account for the electron’s momentum transfer.
+These channels arise within the standard density–density form of the electron–lattice interaction [22, 9] and require no additional disorder.
+The physical content of the elastic recoil channel is elementary.
+In any untethered composite object, an impulse imparted to one constituent immediately changes the total momentum of the whole.
+While redistribution of that momentum among internal degrees of freedom proceeds causally and at finite velocity, the collective momentum of the backgroud crystal appears in the wavefunction and is not dependent of the details of redistribution of the momentum among the constitutuents of the whole.
+.
+An instructive analogue is provided by Mössbauer scattering, in which the dominant elastic line is recoil of the entire lattice without internal phonon excitation.
+A conceptual roadblock has been that, in contrast to processes such as Mössbauer and neutron scattering, where elastic, non-Bragg diffuse deflection is widely acknowledged, conduction electrons do not enter and exit; they start and stay in the crystal.
+The roadblock is addressed by assigning pseudomomentum to both a “foreground” electron density and a “background” consisting of everything else.
+We hasten to point out that the question here involves real physics, and not mere representation: Must phonons be created or destroyed when electrons deflect in a pure crystal, or not?
+On the meaning of elastic scattering
+Elastic scattering of interior electrons is most familiar in the context of impurity or defect scattering, where translational symmetry is explicitly broken.
+Here, we restore pseudomomentum to the status of a conserved momentum, and as in free space what then is conserved in such an elastic deflection?
+The answer will not be controversial: the total momentum of the whole system is conserved.
+What we believe is new is that two pseudmomomenta have changed equally and oppositely in such a collision: one belonging to the electron, another to the rest of the crystal.
+The total momentum, the sum of the two, is unchanged.
+In a perfectly periodic crystal, the relationship between changes in electronic pseudomomentum and the mechanical momentum of the lattice background is less often made explicit.
+In the present formulation this bookkeeping becomes transparent once (i) the lattice is treated as fully dynamical, (ii) pseudomomentum is assigned both to the electron and to the background in the absence of the electron density, and (iii) the interaction is expressed in density–density form, making the exchange of pseudomomentum explicit.
+Finally, we clarify our terminology.
 In a finite-temperature many-body system, strictly elastic processes do not exist in an absolute sense.
-Even the dominant “elastic” line in Mössbauer scattering exhibits slight thermal broadening, reflecting the thermal motion of the emitting nucleus and its finite velocity autocorrelation time.
-The term elastic is nevertheless standard in this context, and we adopt it here with the same understood margin.
-When we use the term quasielastic, we instead refer to processes involving energy exchange that is appreciably larger than the thermal broadening scale, but still too small to account for the full momentum change of the electron.
+Even the dominant line in Mössbauer scattering exhibits slight thermal broadening.
+We therefore use the term elastic to denote processes in which phonon occupation does not change, and quasielastic to denote processes involving energy exchange that is small compared with the momentum transfer involved.
 The landscape seen by an electron
@@
 In a particularly prescient analysis, Kohn emphasized that crystal momentum is a conserved quantity associated with translational symmetry, while the electrical current need not be proportional to it, and that momentum may be exchanged with the lattice without the necessity of phonon excitation or energy dissipation [17].
-This observation already points to the possibility that momentum conservation can be satisfied through coherent motion of the lattice background rather than through inelastic scattering processes.
+This observation already points to the possibility that momentum conservation can be satisfied through recoil of the lattice background rather than through inelastic scattering processes.
 What has remained largely implicit, however, is how this distinction should be implemented at the level of a local scattering event occurring deep in the interior of a crystal.
-In particular, the possibility that total crystal may be explicitly partitioned between a tagged electron and the remaining lattice, both with pseudomomenta appearing as dynamical variables in the many-body wavefunction, has not been systematically developed.
-The present work makes this partition explicit and operational, thereby extending Kohn’s insight to a microscopic framework in which elastic deflection, coherent lattice recoil, and diffusive transport can coexist naturally in a perfect crystal.
+Here, we explicitly assign pseudomomentum to both a foreground electron and the background lattice, with both pseudomomenta appearing as dynamical variables in the many-body wavefunction.
+Making this partition explicit is a realization of Kohn’s insight, within a framework in which elastic deflection, coherent lattice recoil, and diffusive transport can coexist naturally in a perfect crystal.
 In what follows, we formalize a symmetric foreground-background structure, derive the elastic scattering amplitude within first-order perturbation theory, and compute the elastic fraction explicitly.
@@
 In a crystalline environment, the electronic Hamiltonian is invariant under discrete translations of the electron by lattice vectors, with the background held fixed.
-The associated conserved quantity is crystal momentum (pseudomomentum); for an electron in a periodic potential this is the familiar Bloch momentum $\hbar\mathbf{k}$.
+The associated conserved quantity is pseudomomentum (crystal momentum); for an electron in a periodic potential this is the familiar Bloch momentum $\hbar\mathbf{k}$.
 Crucially, the same discrete translational symmetry applies to the background (see Fig.
@@
 Figure 1: Sliding Blocks.
-Floating in space, two blocks attract one another and slide frictionlessly along their common long axis, $x$.
+Floating in space, two blocks attract one another and slide frictionlessly along their common long axis.
 At the top, both blocks carry conserved mechanical momenta, and clearly $P_{\rm tot}=p_{A}+p_{B}$.
@@
 For convenience, we can set it to 0; it never changes and we can now ignore it.
-Earlier we remarked that a pseudomomentum cannot be combined with a mechanical momentum.
-We have now seen, however, that two pseudomomenta can combine to form a mechanical momentum.
+Even under the usual rule stating that defects, such as a crystal edge or impurity, scatter electrons elastically, there is a subtlety: the electron changes its pseudomomentum, but the lattice cannot the respond or recoil in kind.
+A pseudomomentum cannot be combined with a mechanical momentum.
+We have now seen, however, that both foreground and background have pseudomomentum, and can exchage it freely.
+Total momentum is conserved for example in the very real deflection of an electron at an edge by equal and opposite pseudomomentum changes.
 These rules are analogous to those governing the collision of two objects in free space, except that here the interaction is periodic in their relative coordinate, apart from defects and lattice vibrations.
@@
 2).
-When a crystal containing an internal electron is treated as a closed many-body system, the total background momentum $\mathbf{P}=\sum_{i}\mathbf{p}_{i}$ is defined at each instant as the generator of translations of the background degrees of freedom.
-Since the foreground is held fixed in this construction, $\mathbf{P}$ is a pseudomomentum.
-An impulse imparted to a single atom in the background by a foreground object changes the total background momentum immediately by the same amount, resulting in a shift of the center-of-mass velocity $\Delta\mathbf{V}_{\mathrm{CM}}=\Delta\mathbf{p}/M$, independent of any subsequent internal redistribution of momentum.
-The finite-speed propagation of stress, phonons, or other internal excitations governs how momentum is shared among individual degrees of freedom, but it does not affect the instantaneous value of the conserved quantity itself.
-This kinematic distinction follows directly from translational invariance and does not imply nonlocal dynamics or superluminal signaling; rather, it reflects the separation between global conservation laws and the causal evolution of local observables.
+The very real recoil of the atom labeled in red in Fig.
+2 may very well prove to have excited no vibrational quanta when measured in a number state basis.
+This is just the elementary fact that a slightly displaced ground state of a harmonic oscillator is still mostly in the ground state, if measured.
+Two measurable subsystem momenta and the total momentum
+Figure 1 shows two variations on a mechanical system. At the top, two smooth blocks interact frictionlessly. At the bottom, both blocks are corrugated, leading to periodic interaction.
 Figure 2: The total momentum (black) remains constant as the foreground (electron, blue) and background (lattice, red) momenta jump abruptly at the collision time $t^{*}$.
@@
 Finite-window symmetry
-The interaction is periodic only over a finite distance $L$, as required for any physically realizable system.
-This situation is directly analogous to the classic particle in a box.
-Momentum is not strictly conserved in such a system; nevertheless, traveling wave packets carrying well-defined momentum can be constructed as linear combinations of standing-wave eigenstates, which individually carry no momentum.
-Over finite times these wave packets propagate as if momentum were conserved, until a collision with a boundary reverses their motion.
-This type of approximate symmetry, valid only over finite spatial and temporal windows, is sometimes implicit in discussions of wave-packet dynamics in finite systems.
-For clarity, we refer to it here as finite-window symmetry.
-If one insists on working with finite crystals, this concept must be adopted: translational symmetry holds only within a spatial and temporal window set by the system size.
-In a large crystal this construction is an excellent approximation.
-Localized corpuscles can be formed that propagate nearly as Bloch waves with well-defined pseudomomentum and experience long travel times before encountering a boundary.
-We emphasize that this corpuscle construction is a calculational convenience that permits the use of the notation and results associated with the formally unphysical Born-von Kármán boundary conditions, without assuming their literal validity.
+In any physically realizable crystal the interaction is periodic only over a finite spatial extent $L$.
+Strict translational invariance is therefore absent at the boundaries, and crystal momentum is not an exact conserved quantity for the full finite system.
+This situation is directly analogous to the textbook particle in a box: the exact eigenstates are standing waves that do not carry definite momentum, yet traveling wave packets with well-defined average momentum can be constructed as superpositions of these eigenstates.
+Over times short compared with the traversal time to the boundary, such wave packets propagate as if momentum were conserved.
+The same reasoning applies to a large but finite crystal.
+Although global translational symmetry is broken by the boundaries, the Hamiltonian is locally periodic in the bulk.
+Wave packets localized far from the boundaries evolve, for long but finite times, according to the translationally invariant bulk Hamiltonian.
+In this regime one may assign a well-defined crystal pseudomomentum to the packet, up to corrections that vanish as $L\to\infty$.
+This type of approximate symmetry is standard in discussions of the thermodynamic limit, bulk transport theory, and wave-packet dynamics in finite systems.
+For clarity, we refer to it here as finite-window symmetry: translational invariance holds within a spatial and temporal window set by the system size.
+No new symmetry is implied; rather, this terminology emphasizes that the use of bulk momentum quantum numbers in a finite crystal is an approximation controlled by the ratio of the observation time and length scales to the system size.
+In a macroscopic crystal this approximation is exceedingly accurate.
+Localized wave packets (“corpuscles”) can be constructed that propagate nearly as Bloch waves with well-defined pseudomomentum and experience long travel times before interacting with boundaries.
+This construction permits the use of the notation and results associated with Born–von Kármán boundary conditions as a calculational device, without assuming literal periodicity of the finite sample.
+In a finite crystal, the exact normal modes are standing waves.
+As exact eigenmodes they have zero net mechanical momentum.
+Any “phonon carrying momentum $\hbar q$” is then not an eigenstate; it’s a wavepacket (a superposition of standing modes) that only behaves like a traveling object transiently until it reflects, mode-mixes, etc.
 III Lattice degrees of freedom and zero modes
@@
 The mass-weighted constraints ensure that all internal degrees of freedom carry zero net momentum, so that $\hat{\mathbf{R}}$ describes the unique translational zero mode.
-The classical and quantum theory of lattice vibrations explicitly separates rigid motion from internal vibrations.
+The classical and quantum theory of lattice vibrations explicitly separates rigid translation from internal vibrations.
 Born and Huang showed that the equations of motion of a finite crystal admit three zero-frequency normal modes corresponding to rigid translation of the crystal as a whole, while the internal normal modes satisfy the mass-weighted constraint $\sum_{j}m_{j}\hat{\mathbf{u}}_{j}=\mathbf{0},$ so that internal vibrations carry no net mechanical momentum [7].
 All mechanical momentum of the crystal resides in the translational zero mode.
-The total center-of-mass degree of freedom of the combined foreground-background system is always conserved and need not be carried explicitly.
-For an interior electron, it is crucial to track the pseudomomentum of the foreground electron and separately the background,
 IV Electron-Lattice interaction
 IV.1 Foreground and background Hamiltonians
-Given our aim of formulating transport without assuming pre-existing quasiparticles, care is required in defining the Hamiltonian. We write the total Hamiltonian as
+. We write the total Hamiltonian as
 $H=H_{\rm fg}+H_{\rm bg}+H_{\rm int},$ (9)
@@
 $\hat{\rho}^{\rm int}_{\mathbf{q}}$ $=\sum_{j}e^{-i\mathbf{q}\cdot(\bm{\xi}^{(0)}_{j}+\hat{\mathbf{u}}_{j})}.$
-This identity expresses the fact that a uniform translation of the lattice multiplies all density Fourier components by the phase $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$, while all information about internal lattice dynamics is contained in $\hat{\rho}^{\rm int}_{\mathbf{q}}$.
+Here $\hat{\mathbf{R}}$ is the lattice center-of-mass operator, $\bm{\xi}^{(0)}_{j}$ are equilibrium positions, and $\hat{\mathbf{u}}_{j}$ are internal displacements.
+Equation (11) is an exact operator identity, independent of any choice of basis (plane waves, Bloch waves, or standing waves).
+This identity expresses a purely kinematic fact: a rigid translation of the lattice multiplies every density Fourier component by the phase $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$, while all internal lattice dynamics are contained entirely in $\hat{\rho}^{\rm int}_{\mathbf{q}}$.
 Density–density interaction and pseudomomentum conservation
-With this factorization, the electron–lattice interaction may be written in density–density form as
+The electron–lattice interaction may be written in density–density form as
 $H_{\rm int}=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}\,\hat{\rho}_{-\mathbf{q}},$ (12)
-which follows the standard Born-approximation scattering formalism introduced by Van Hove [34]. Using Eq. (10), this becomes
-$H_{\rm int}$ $=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}\,e^{i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{-\mathbf{q}}$ (12)
+following the formalism of Van Hove [30]. No assumption about plane-wave eigenstates is required; $\mathbf{q}$ labels Fourier components of the interaction kernel.
+Using Eq. (11), this becomes
+$H_{\rm int}$ $=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i{\mathbf{q}}\cdot\hat{\bm{\rho}}}\,\,\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (13)
 $=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}\,\hat{\rho}^{\rm int}_{-\mathbf{q}}.$
@@
 The background zero mode is therefore capable of absorbing finite pseudomomentum with negligible energy cost, providing the kinematic basis for elastic but momentum-relaxing scattering.
-Any combination of phonon involment and lattice background recoil is allowed.
-It seems this should matter to measureable properties.
+Any combination of phonon involvement and lattice background recoil is allowed.
+It seems this should matter for measurable properties.
 V Elastic internal scattering and the Debye-Waller fraction
@@
 The resulting processes are elastic with respect to internal lattice excitations, despite involving momentum exchange with the lattice as a whole.
-Density-density coupling
-We consider a single tagged (foreground) electron with coordinate $\hat{\mathbf{r}}$. The interaction between the foreground electron and the background lattice is taken to be of density-density form,
-$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,\rho_{\rm fg}(\mathbf{q})\,\rho_{\rm bg}(-\mathbf{q}).$ (13)
-Separating the background density into its static equilibrium expectation value and fluctuations, $\rho_{\rm bg}=\langle\rho_{\rm bg}\rangle+\delta\rho_{\rm bg}$, the static component generates a periodic potential acting on the foreground electron,
+Density–density coupling and emergence of the Bloch potential
+We take the foreground–background interaction in density–density form,
+$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,\rho_{\rm fg}(\mathbf{q})\,\rho_{\rm bg}(-\mathbf{q}),$ (14)
+with the tagged-electron density operator
+$\rho_{\rm fg}(\mathbf{q})=e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}$. (15)
+For the background we use the exact center-of-mass factorization
+$\rho_{\rm bg}(-\mathbf{q})=\sum_{j}e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}_{j}}=e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (16)
+where $\hat{\mathbf{R}}_{j}=\hat{\mathbf{R}}+\bm{\xi}_{j}^{(0)}+\hat{\mathbf{u}}_{j}$.
+Separating the background density operator into its equilibrium expectation value (taken in the background center-of-mass frame) and fluctuations,
+$\rho_{\rm bg}(-\mathbf{q})=\langle\rho_{\rm bg}(-\mathbf{q})\rangle+\delta\rho_{\rm bg}(-\mathbf{q}),$ (17)
+the static component produces an effective periodic potential acting on the foreground electron,
 $U_{\rm Bloch}(\hat{\mathbf{r}})=\sum_{\mathbf{G}}V(\mathbf{G})\,e^{i\mathbf{G}\cdot\hat{\mathbf{r}}}\,\langle\rho_{\rm bg}(-\mathbf{G})\rangle,$ (18)
-where $\mathbf{G}$ are reciprocal lattice vectors. The resulting foreground Hamiltonian is therefore
+where $\mathbf{G}$ are reciprocal lattice vectors.
+The corresponding single-particle foreground Hamiltonian becomes
 […]
-while the remaining part of $H_{\rm int}$ describes the dynamical coupling to background fluctuations.
-In this way the familiar Bloch Hamiltonian emerges directly from the same density-density interaction, without being assumed a priori.
-The lattice degrees of freedom (the background) are described by a Hamiltonian $H_{\rm bg}$. The coupling between foreground and background is retained in the form
-$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,\rho_{\rm fg}(\mathbf{q})\,\rho_{\rm bg}(-\mathbf{q}),$ (16)
-where the tagged-electron density operator is
-$\rho_{\rm fg}(\mathbf{q})=e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}$, (17)
-and the full background density operator is
-$\rho_{\rm bg}(-\mathbf{q})=\sum_{j}e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}_{j}}=e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (18)
-with $\hat{\mathbf{R}}_{j}=\hat{\mathbf{R}}+\bm{\xi}^{(0)}_{j}+\hat{\mathbf{u}}_{j}$ as defined previously.
+which is the usual Bloch Hamiltonian. This periodic potential is not assumed a priori, but emerges as the static mean-field component of the underlying density–density coupling.
+The remaining part of $H_{\rm int}$ couples the foreground electron to background density fluctuations $\delta\rho_{\rm bg}$, but with the crucial difference from the clamped-lattice formulation: the full operator $\rho_{\rm bg}(-\mathbf{q})$ retains the explicit center-of-mass translation factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$, which enforces exact momentum bookkeeping in the finite, untethered crystal.
 Initial and final states
@@
 $\Delta\mathbf{p}=+\hbar\mathbf{q},\qquad\Delta\mathbf{P}=-\hbar\mathbf{q},$ (21)
-so that the total mechanical momentum of the combined system is conserved. The recoil pseudomomentum is carried by the lattice center-of-mass degree of freedom.
-While much stronger coupling regimes can be treated nonperturbatively—for example within wave-on-wave (WoW) simulations—the purpose of the present analysis is to emphasize that even within standard perturbation theory there exist elastic scattering processes that are usually excluded by more restrictive formulations.
+so that the total mechanical momentum of the combined system is conserved. The recoil pseudomomentum is carried by the lattice center-of-mass degree of freedom, regardless of phonon number changes.
 Elastic internal channel and the Debye-Waller factor
@@
 which is the direct analog of the recoil-free (zero-phonon-line) fraction in Mössbauer spectroscopy and in neutron or x-ray scattering.
+While much stronger coupling regimes can be treated nonperturbatively—for example within wave-on-wave (WoW) simulations [4] the present analysis shows that even within standard perturbation theory there exist elastic scattering processes that are absent when the center of mass is hidden or eliminated.
 Relation to the Fröhlich truncation
@@
 (11) is not employed.
-The lattice center-of-mass coordinate $\hat{\mathbf{R}}$ is not retained explicitly, which is equivalent to setting the recoil factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$ to unity and treating the lattice as an infinitely massive reference frame.
+The lattice center-of-mass coordinate $\hat{\mathbf{R}}$ is not retained explicitly, which is equivalent to setting the recoil factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$ to unity and treating the lattice as a clamped reference frame.
 At the same time, the internal displacement exponential is truncated to linear order,
@@
 The zeroth-order term reproduces the static periodic potential responsible for Bloch band formation, while the linear term generates single-phonon processes.
-Although this procedure is internally consistent and highly successful in the weak-coupling regime, it removes the translational zero mode by construction.
+Although this procedure is internally consistent and highly successful in the weak-coupling regime, it removes the background recoil zero mode by construction.
 As a result, even if the full exponential in $\hat{\mathbf{u}}_{j}$ were retained, the truncated theory would still lack an explicit degree of freedom capable of absorbing the total recoil momentum of the lattice.
-The standard Fröhlich Hamiltonian therefore describes momentum exchange with internal lattice modes, but not with the lattice as a whole.
-Elastic, quasielastic, and inelastic processes
-Retaining the full operator $e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}$ in the background density ensures that elastic, quasielastic, and inelastic scattering processes are treated on equal footing already at first order in the density-density interaction.
-At this level, the interaction transfers a definite total momentum $\hbar\mathbf{q}$ from the foreground electron to the background, independent of how that momentum is subsequently absorbed.
-It is therefore useful to decompose the momentum transfer as $\mathbf{q}=\mathbf{q}_{\rm el}+\mathbf{q}_{\rm in}$, where $\mathbf{q}_{\rm el}$ corresponds to center-of-mass recoil of the background and $\mathbf{q}_{\rm in}$ is taken up by its internal degrees of freedom.
-The latter satisfies the exact kinematic relation
-$\hbar\mathbf{q}_{\rm in}=-\sum_{\mathbf{k},\lambda}\hbar\mathbf{k}\,\Delta n_{\mathbf{k}\lambda},$ (25)
-where $\Delta n_{\mathbf{k}\lambda}$ denotes the change in occupation of the background normal modes.
-In this formulation, the distinction between elastic and inelastic scattering is determined by the background state rather than by an expansion in ionic displacements.
+The standard Fröhlich Hamiltonian therefore describes momentum exchange with internal lattice modes, but the 0 mode, the lattice recoil as a whole, is absent.
 Thermal averaging and the Debye-Waller fraction
@@
 where $|n\rangle$ denote exact many-body eigenstates of the lattice Hamiltonian $\hat{H}_{\mathrm{lat}}$ with energies $E_{n}$.
-For elastic internal scattering the relevant quantity is the thermal average of the phonon-diagonal matrix element of the internal density operator,
+For elastic internal scattering the relevant quantity is the thermal expectation value of the internal density operator,
 $\langle\rho^{\rm int}_{\mathbf{q}}\rangle_{T}=\sum_{n}p_{n}\,\langle n|\rho^{\rm int}_{\mathbf{q}}|n\rangle.$ (27)
-The corresponding elastic (coherent) scattering intensity is then given by
+The coherent elastic scattering intensity is
 $I_{\mathrm{el}}(\mathbf{q})=\big|\langle\rho^{\rm int}_{\mathbf{q}}\rangle_{T}\big|^{2}.$ (28)
-By contrast, the total scattering probability summed over all final internal states is
+The total scattering intensity summed over all final internal states is
 $I(\mathbf{q})=\left\langle\rho^{\rm int}_{\mathbf{q}}\rho^{\rm int}_{-\mathbf{q}}\right\rangle_{T}.$ (29)
-The elastic internal branching fraction is therefore
-$f_{\mathrm{el}}(\mathbf{q},T)=\frac{\sum_{\alpha}p_{n}\,\big|\langle n|\rho^{\rm int}_{\mathbf{q}}|n\rangle\big|^{2}}{\left\langle\rho^{\rm int}_{\mathbf{q}}\rho^{\rm int}_{-\mathbf{q}}\right\rangle_{T}}.$ (30)
+The elastic (coherent) branching fraction is therefore
+$f_{\mathrm{el}}(\mathbf{q},T)=\frac{\big|\langle\rho^{\rm int}_{\mathbf{q}}\rangle_{T}\big|^{2}}{\left\langle\rho^{\rm int}_{\mathbf{q}}\rho^{\rm int}_{-\mathbf{q}}\right\rangle_{T}}.$ (30)
 For a harmonic lattice one finds[18, 21]
 $f_{\mathrm{el}}(\mathbf{q},T)\simeq e^{-2W(\mathbf{q},T)},\qquad W(\mathbf{q},T)=\frac{1}{2}\left\langle(\mathbf{q}\cdot\hat{\mathbf{u}})^{2}\right\rangle_{T},$ (31)
-showing that a substantial fraction of momentum-relaxing scattering events can be strictly elastic internally even at finite temperature.
+the familiar Debye–Waller result. Even at finite temperature a substantial fraction of momentum-transfer events leave the lattice in the same internal state.
 Estimate of the elastic internal fraction at representative momentum transfers
@@
 Even for large-angle scattering with $q\sim 2k_{F}$, the majority of momentum-relaxing events remain phonon-diagonal, with momentum absorbed predominantly by the lattice background rather than by internal phonon excitation.
-These estimates show that internal phonon excitation constitutes only a minor branching channel for momentum relaxation in copper, even at room temperature and for momentum transfers relevant to transport.
-The dominant mechanism for absorbing electron momentum is thus elastic or quasi-elastic coupling to the lattice background, consistent with the discussion above.
+These estimates show that, for momentum transfers relevant to transport, the overwhelming fraction of scattering events are internally elastic, in the sense that they leave the lattice in the same vibrational state.
+Phonon creation constitutes only a minor branching channel at room temperature, even for large-angle scattering.
+Momentum exchange with the lattice therefore occurs predominantly through elastic or quasi-elastic coupling to the background degrees of freedom, with internal phonon excitation representing a secondary correction.
 VI Experiments suggesting large elastic contributions
@@
 Quantum interference and dephasing
-Traditionally, the suppression of quantum-interference effects with increasing temperature has been attributed to inelastic scattering processes that transfer energy to the environment [22, 2].
-However, it is important to distinguish carefully between dephasing, as operationally defined in transport experiments, and irreversible decoherence in the strict quantum-mechanical sense.
-A wide range of experiments demonstrate that genuine energy relaxation remains slow over the same temperature range in which phase coherence is strongly suppressed.
-In particular, mesoscopic interference experiments show a rapid decrease—and in some cases saturation—of the phase coherence time $\tau_{\phi}$ at low temperatures, even as independent probes indicate that electron-phonon energy exchange rates continue to vanish [26, 16, 35].
-This separation of timescales shows that momentum and phase randomization need not be accompanied by significant energy transfer.
-As observed in weak-localization experiments on mesoscopic metals, the loss of quantum interference reflects the destruction of coherent backscattering between time-reversed trajectories.
-This coherence is far more sensitive than the transport relaxation time and can be suppressed by elastic, time-dependent scattering with negligible energy exchange and without irreversible decoherence [22, 26].
-From this perspective, dephasing reflects the extreme sensitivity of quantum interference to weak, time-dependent fluctuations of the environment.
-Even in the absence of thermal excitations, lattice coordinates remain quantum operators with nonvanishing variance due to zero-point motion.
-As a result, elastic scattering amplitudes are not unity even at $T=0$, as expressed by the Debye-Waller factor.
-When such quantum fluctuations are slow on electronic timescales—as occurs for long-wavelength modes in finite samples—they act as an effectively time-dependent scattering potential, suppressing interference under ensemble and time averaging.
-These processes are elastic on electronic energy scales and do not constitute irreversible decoherence.
+The suppression of quantum-interference effects with increasing temperature is commonly described in terms of inelastic scattering processes that exchange energy with environmental degrees of freedom [20, 1].
+It is useful, however, to distinguish between dephasing as operationally defined in transport experiments and irreversible decoherence in the strict quantum-mechanical sense.
+A broad class of experiments indicates that substantial suppression of interference can occur even when direct energy relaxation to the lattice remains weak.
+Mesoscopic interference measurements, for example, show a pronounced reduction— and in some cases an apparent saturation—of the phase-coherence time $\tau_{\phi}$ at low temperatures, while independent probes suggest that electron–phonon energy exchange rates continue to decrease in the same regime [23, 15, 31].
+This empirical separation of timescales illustrates that phase and momentum randomization need not scale directly with the rate of substantial energy transfer to internal lattice excitations.
+Weak-localization is robust in experiments, implying that the backscattering is coherent and therefore elastic.
+[20, 23].Elastic scattering amplitudes, including diffuse non-Bragg amplitudes, are reduced by the well known Debye–Waller factor.
 Hot-electron relaxation
-A particularly direct separation of momentum relaxation from energy relaxation is provided by hot-electron experiments in metals [36, 30, 11, 12].
-In these measurements, nonequilibrium electron distributions are injected and their subsequent momentum randomization and energy loss are probed independently.
-Over wide temperature ranges, the momentum-relaxation time inferred from transport or magnetotransport is found to be much shorter than the energy-relaxation time associated with thermalization.
-Electrons therefore undergo many momentum-deflecting collisions while remaining essentially isoenergetic.
-Such a pronounced separation of timescales cannot be explained by purely inelastic phonon scattering and instead points to an effectively elastic deflection channel intrinsic to the metallic state.
-Johnson-noise thermometry
-Johnson-noise thermometry provides a particularly stringent probe of energy relaxation in driven metallic microstructures [36, 30].
-In these experiments, a dc bias drives a substantial current while the local electronic temperature is inferred directly from the Johnson-Nyquist noise.
-Because the noise power is proportional to the electronic energy distribution, this technique provides a quantitative and noninvasive measure of energy flow, independent of momentum transport.
-A key observation is that under bias the electrical resistivity and momentum randomization rate increase strongly, while the electronic temperature remains comparatively low and spatially nonuniform.
-This demonstrates that electrons can undergo many momentum-deflecting scattering events without efficiently transferring energy to the lattice.
-Such behavior is difficult to reconcile with a picture in which inelastic phonon scattering controls momentum relaxation, but follows naturally if the dominant scattering mechanism is elastic deflection from a dynamically fluctuating lattice background.
+A particularly direct separation of momentum and energy relaxation is provided by hot-electron experiments in metals [32, 26, 10, 11].
+In these measurements, nonequilibrium electron distributions are injected, and the subsequent momentum randomization and energy loss are probed independently.
+Over broad temperature ranges, the momentum-relaxation time inferred from transport or magnetotransport is found to be substantially shorter than the energy-relaxation time associated with thermalization.
+Electrons therefore undergo many momentum-deflecting collisions while remaining close to isoenergetic on the same timescale.
+This pronounced separation of timescales indicates that efficient momentum randomization does not require comparably rapid energy transfer to lattice excitations.
+While inelastic electron–phonon scattering ultimately governs thermalization, the observed hierarchy suggests that additional scattering channels, with weak net energy exchange on electronic scales, can play an important role in momentum relaxation within the metallic state.
 VII Discussion
 The purpose of this section is twofold.
-First, we examine the consistency of the present framework with well-established experimental facts, including the Wiedemann-Franz law, weak localization, and quantum oscillations.
-These topics are not speculative; they address potential objections that might otherwise arise when elastic scattering plays a dominant microscopic role.
-Second, we outline broader interpretive implications for Planckian transport, linear-in-$T$ resistivity, and the Mott-Ioffe-Regel crossover.
-These latter points are necessarily more exploratory and are presented as a coherent physical interpretation rather than as definitive conclusions.
+First, we examine the consistency of the present framework with well-established experimental results, including the Wiedemann–Franz law, weak localization, and quantum oscillations.
+These considerations address potential concerns regarding the role of predominantly elastic momentum-relaxing channels.
+Second, we outline broader interpretive implications for Planckian transport, linear-in-$T$ resistivity, and the Mott–Ioffe–Regel crossover.
+The latter discussion is exploratory and is presented as a coherent physical interpretation rather than as a definitive claim.
 VII.1 Implications for the Lorenz ratio
-Because elastic scattering randomizes momentum just as effectively as inelastic scattering, no contradiction arises between predominantly elastic microscopic dynamics and the well-established success of conventional transport theory.
-In a degenerate metal, the Wiedemann-Franz (WF) law states that the ratio of the electronic thermal conductivity $\kappa$ to the electrical conductivity $\sigma$ times temperature approaches the universal Sommerfeld value
+Elastic scattering randomizes momentum as effectively as inelastic scattering within standard transport formalisms.
+Accordingly, a predominantly elastic microscopic channel does not conflict with the established success of conventional transport theory.
+In a degenerate metal, the Wiedemann–Franz (WF) law states that the ratio of the electronic thermal conductivity $\kappa$ to the electrical conductivity $\sigma$ times temperature approaches the Sommerfeld value
 $L\equiv\frac{\kappa}{\sigma T}\;\xrightarrow[T\to 0]{}\;L_{0}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}.$ (35)
-While some fraction of inelastic processes must exist to relax the electronic energy current and establish global thermal equilibrium, it has long been understood that these processes need not control the transport coefficients themselves [39, 21, 31].
-If the dominant momentum-relaxing channel is elastic and only weakly dependent on energy over the thermal window $\sim k_{B}T$ about the Fermi level, then both electrical and thermal currents are governed by the same relaxation time $\tau_{\mathrm{el}}$,
+Although some inelastic processes must ultimately relax the electronic energy current and establish global thermal equilibrium, it has long been understood that these processes need not determine the leading transport coefficients themselves [35, 19, 27].
+If the dominant momentum-relaxing channel is elastic and only weakly energy dependent over the thermal window $\sim k_{B}T$ about the Fermi level, then both electrical and thermal currents are governed by the same relaxation time $\tau_{\mathrm{el}}$,
 $\sigma\propto\tau_{\mathrm{el}},\qquad\kappa\propto\tau_{\mathrm{el}},$ (36)
-and the Sommerfeld value $L_{0}$ is recovered.
-Rare inelastic events, with rate $1/\tau_{\mathrm{in}}\ll 1/\tau_{\mathrm{el}}$, serve to transfer energy from the electronic system to internal lattice degrees of freedom, but do not enter the leading expressions for $\sigma$ or $\kappa$.
-A transport regime dominated by elastic momentum randomization is therefore fully compatible with the WF law, even when true energy relaxation is parametrically slow.
+and the Sommerfeld value $L_{0}$ follows in the usual way.
+Inelastic processes with rate $1/\tau_{\mathrm{in}}\ll 1/\tau_{\mathrm{el}}$ transfer energy between electrons and lattice degrees of freedom but do not enter the leading expressions for $\sigma$ or $\kappa$.
+A transport regime dominated by elastic momentum randomization is therefore fully compatible with the WF law, even when true energy relaxation is parametrically slower.
 VII.2 Weak localization
-Weak localization (WL) provides a particularly clear illustration of the dominance of elastic backscattering in metals at low temperature.
-The phenomenon arises from constructive interference between pairs of time-reversed electronic trajectories that return to the same point after multiple elastic scattering events [6, 22].
-The very existence of a WL correction therefore implies that the dominant backscattering processes are coherent and elastic.
-If backscattering were predominantly inelastic, interference between time-reversed paths would be washed out and no WL correction could develop.
-Instead, experiments show a well-defined suppression of the conductivity at low temperature together with a characteristic negative magnetoresistance when time-reversal symmetry is broken by an applied magnetic field.
-As temperature is increased, WL is gradually suppressed.
-Although this suppression is often attributed to “inelastic scattering,” experimentally it reflects the loss of phase coherence rather than rapid energy relaxation.
-Independent probes such as hot-electron relaxation and Johnson-noise thermometry show that electron-phonon energy exchange remains slow over the same temperature range.
-The suppression of WL therefore signals the onset of a time-dependent scattering environment that randomizes phase while leaving electronic energy essentially unchanged.
-Within the present framework, WL fits naturally into the hierarchy of timescales
+Weak localization (WL) provides a clear illustration of the central role of elastic backscattering in low-temperature metallic transport.
+The phenomenon arises from constructive interference between pairs of time-reversed electronic trajectories that revisit the same spatial region after multiple scattering events [6, 20].
+The WL correction therefore requires that a substantial fraction of backscattering processes preserve phase coherence over the relevant timescale.
+If scattering were strongly inelastic on the transport timescale, interference between time-reversed paths would be suppressed before a WL correction could develop.
+Experimentally, however, a well-defined suppression of conductivity and characteristic negative magnetoresistance are observed at low temperature.
+As temperature increases, the WL correction is gradually reduced.
+Although this reduction is commonly described in terms of increasing “inelastic scattering,” what is operationally measured is the loss of phase coherence rather than a rapid increase in electronic energy relaxation.
+Independent probes, including hot-electron relaxation and Johnson-noise thermometry, indicate that electron–phonon energy exchange can remain comparatively slow in the same temperature range.
+The suppression of WL is therefore consistent with the onset of a time-dependent scattering environment that randomizes phase while involving only modest energy transfer on electronic scales.
+Within the present framework this behavior fits naturally into the hierarchy of timescales
 $\tau_{p}\ll\tau_{\phi}\ll\tau_{E},$
-with elastic scattering producing diffusive transport and slower temporal evolution of the lattice background destroying phase coherence without necessitating phonon creation or significant energy dissipation.
-VII.3 Shubnikov-de Haas oscillations
-Quantum oscillation phenomena such as the Shubnikov-de Haas and de Haas-van Alphen effects provide an independent and equally stringent demonstration that electronic scattering in metals is predominantly elastic at low temperature [33, 3].
+where elastic momentum randomization produces diffusive transport, while slower temporal evolution of the lattice background reduces phase coherence without requiring substantial phonon excitation or rapid energy dissipation.
+VII.3 Shubnikov–de Haas oscillations
+Quantum oscillation phenomena such as the Shubnikov–de Haas and de Haas–van Alphen effects provide an independent probe of electronic scattering in metals at low temperature [29, 2].
 These oscillations arise from the quantization of electronic orbits in a magnetic field and reflect coherent motion around the Fermi surface over many cyclotron periods.
@@
 where $\tau_{q}$ is the quantum lifetime associated with Landau-level broadening.
-Importantly, $\tau_{q}$ is primarily sensitive to elastic momentum-randomizing processes, while genuinely inelastic energy exchange rapidly suppresses oscillations.
-Any scattering event that transfers an energy comparable to $\hbar\omega_{c}$ destroys the coherence required to complete a cyclotron orbit.
-The observation of quantum oscillations over broad temperature ranges and down to low magnetic fields therefore implies that electrons undergo many momentum-deflecting collisions without significant energy exchange.
-As in the case of weak localization, the temperature dependence of the oscillation amplitude reflects phase smearing due to a time-dependent scattering environment rather than rapid inelastic backscattering.
-Quantum oscillations thus provide direct experimental evidence that elastic scattering dominates electron dynamics in normal metals at low temperature.
-VII.4 Sound attenuation
-Ultrasonic attenuation in metals provides a complementary probe of electron-lattice dynamics from the lattice side.
-Experimentally, sound waves propagating through a metal are strongly attenuated even in very clean samples, and this attenuation is often interpreted as evidence for inelastic electron-phonon scattering [28, 25].
-However, attenuation of a coherent acoustic wave does not, by itself, imply irreversible energy dissipation.
-From a microscopic perspective, attenuation measures the loss of coherence of a propagating phonon mode.
-This loss can arise through elastic scattering processes that redistribute acoustic momentum among phonon modes or transfer momentum to the electronic subsystem, even when the total energy of the combined electron-lattice system is conserved.
-In this sense, sound attenuation is directly analogous to the suppression of quantum interference in electronic transport: coherent propagation is degraded without requiring strong energy relaxation [19, 1].
-Standard theories of ultrasonic attenuation distinguish regimes based on the relative timescales of the sound frequency and electronic or phononic relaxation rates.
-In the collisionless (Landau-Rumer) regime, attenuation arises from scattering of sound by thermal excitations and mode conversion rather than from viscous energy dissipation [20, 25].
-More generally, the acoustic wave loses forward coherence through elastic momentum exchange with a dynamically fluctuating background, while energy relaxation enters only at longer timescales.
-Within the present framework, elastic momentum exchange between electrons and the lattice background provides a natural mechanism for ultrasonic attenuation.
-Time-dependent lattice distortions scatter sound waves, redirecting acoustic momentum and reducing the amplitude of the forward-propagating mode, while leaving electronic and lattice energies largely unchanged.
-Energy dissipation occurs subsequently through rare inelastic processes that establish thermal equilibrium.
-This interpretation is consistent with experimental observations in normal metals, where substantial ultrasonic attenuation persists even in temperature regimes where independent measurements show that electron-phonon energy transfer is strongly suppressed [28, 25].
-As in weak localization and quantum oscillation phenomena, sound attenuation therefore supports a picture in which elastic scattering dominates momentum transfer, while true energy relaxation plays a secondary role.
-VII.5 Planckian diffusion and linear-in-$T$ resistivity
-Much of the recent literature introduces Planckian behavior as a bound on a microscopic relaxation time $\tau_{\mathrm{Pl}}\sim\hbar/k_{B}T$, often interpreted as the fastest possible rate of inelastic scattering or energy dissipation [37, 14, 13, 32].
-Within this narrative, transport coefficients are obtained by inserting $\tau_{\mathrm{Pl}}$ into kinetic or hydrodynamic expressions.
-The framework developed in Ref. [38] and here suggests a different interpretation. The primary microscopic result is the emergence of real-space diffusion with a diffusion constant
+This lifetime is sensitive to momentum-randomizing processes that broaden Landau levels, while sufficiently strong inelastic energy exchange suppresses the coherence required to complete cyclotron motion.
+Scattering events that transfer energies comparable to $\hbar\omega_{c}$ disrupt the phase coherence of the orbit.
+The observation of well-defined quantum oscillations over broad temperature ranges and down to relatively low magnetic fields therefore indicates that electrons can undergo many momentum-deflecting collisions while maintaining phase coherence over multiple cyclotron periods.
+As in weak localization, the temperature dependence of the oscillation amplitude reflects phase smearing due to temporal fluctuations of the scattering environment.
+These considerations are consistent with a regime in which momentum randomization can occur without rapid or large energy exchange on electronic scales.
+VII.4 Planckian diffusion and linear-in-$T$ resistivity
+Much recent literature discusses Planckian behavior in terms of a microscopic relaxation time
+$\tau_{\mathrm{Pl}}\sim\frac{\hbar}{k_{B}T},$
+often interpreted as setting a characteristic scale for scattering or energy dissipation [33, 13, 12, 28].
+Within this perspective, transport coefficients are commonly estimated by inserting $\tau_{\mathrm{Pl}}$ into kinetic or hydrodynamic expressions.
+The framework developed in Ref. [5, 34] and extended here suggests a complementary interpretation. The primary microscopic result is the emergence of real-space diffusion with diffusion constant
 $D\sim\frac{\hbar}{m^{\ast}},$ (38)
 arising from elastic momentum scrambling in a time-dependent lattice background.
-This form of Planckian diffusion does not rely on quasiparticles, strong inelastic scattering, or rapid energy dissipation; it follows directly from quantum dynamics in a fluctuating potential and has been demonstrated explicitly in wave-on-wave simulations of the full electron-lattice problem [5, 38].
+In this formulation, diffusion emerges from quantum dynamics in a fluctuating potential and does not rely explicitly on strong inelastic scattering or rapid energy dissipation.
+Wave-on-wave simulations of the coupled electron–lattice problem demonstrate this mechanism explicitly [5, 16, 34].
 Once diffusion is established, Einstein relations connect $D$ to transport coefficients,
 $\sigma=\chi D,\qquad D=\mu\frac{k_{B}T}{e},$ (39)
-so that recasting $D$ in terms of an effective relaxation time yields
+so that expressing $D$ in terms of an effective relaxation time yields [5]
 $\tau_{\mathrm{eff}}\sim\frac{\hbar}{k_{B}T},$ (40)
 up to factors of order unity.
-From this perspective, the Planckian timescale is not a fundamental microscopic bound, but an emergent parametrization of quantum-limited diffusion combined with equilibrium thermodynamics.
-It constrains how rapidly charge spreads in space, not how rapidly energy must be irreversibly transferred to internal degrees of freedom.
-This viewpoint also clarifies the emergence of linear-in-$T$ resistivity across and below the Debye scale.
-Within a conventional Bloch-Grüneisen picture, the Debye temperature sets the scale at which all phonon modes become available for inelastic scattering.
-In contrast, within the quantum acoustics framework there is no special anomaly associated with crossing the Debye temperature.
-Linear-in-$T$ resistivity arises smoothly as coherence is lost in a time-dependent elastic background, while energy relaxation remains comparatively slow.
-The Debye scale therefore marks neither the onset of momentum relaxation nor a transition to a new inelastic regime, but rather a coherence crossover.
-VII.6 Mott-Ioffe-Regel crossover as loss of quasiparticle coherence
-The Mott-Ioffe-Regel (MIR) condition $k_{F}\ell\sim 1$ is often interpreted as a boundary beyond which quasiparticles “cease to exist” and transport must become anomalous.
-From the present perspective, this interpretation is misleading.
-The MIR condition marks the point at which a semiclassical mean free path loses a literal geometric meaning, not a dynamical singularity in transport.
-Within the quantum acoustics framework, transport evolves smoothly through $k_{F}\ell\sim 1$.
-Diffusion persists, but it is no longer usefully described in terms of well-defined trajectories between discrete scattering events.
-Crucially, this breakdown of a quasiparticle picture does not coincide with a loss of coherence.
-Were strongly inelastic processes truly dominant at the MIR scale, interference phenomena such as weak localization and quantum oscillations would be rapidly suppressed, contrary to experimental observation.
-Instead, experiments show that momentum relaxation can be strong while inelastic energy relaxation remains weak.
-Wave-on-wave simulations sharpen this picture by explicitly demonstrating that momentum can be transferred elastically to a thermally distorted lattice background while preserving electronic energy.
-The MIR crossover therefore reflects a loss of quasiparticle coherence rather than the onset of strong dissipation, fully consistent with elastic-dominated momentum relaxation in a dynamical lattice environment.
+From this viewpoint, the Planckian timescale may be understood not as a fundamental bound on inelastic scattering, but as an emergent parametrization of quantum-limited diffusion combined with equilibrium thermodynamics.
+It characterizes the rate at which charge spreads spatially rather than the rate at which energy must be irreversibly transferred to internal degrees of freedom.
 VIII Summary
-Summary
-In the conventional clamped-lattice formulation of electron-phonon scattering, the factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$ associated with rigid translation of the lattice background is removed by assumption.
-This elimination suppresses an otherwise allowed elastic momentum-transfer channel, forcing momentum relaxation to be identified exclusively with inelastic phonon creation or annihilation.
-By restoring the lattice translational zero mode, we show that this identification is not required by symmetry, conservation laws, or kinematics.
-In a finite crystal, elastic pseudomomentum transfer between an electron and the lattice background is generically present and can dominate momentum relaxation at low and intermediate temperatures.
-Making this elastic channel explicit requires abandoning several silent but consequential restrictions embedded in the traditional formulation.
-First, the lattice must be treated as a finite, isolated object so that its center-of-mass degree of freedom remains dynamical.
-(If the crystal is attached to a laboratory or substrate, this simply enlarges the background without altering the underlying momentum bookkeeping.) Second, discrete translational symmetry must be implemented symmetrically, with foreground and background pseudomomenta treated as equally real and conserved quantities.
-Third, the electron-lattice interaction must be retained in its full density-density form, rather than truncated in a way that projects out elastic momentum-transfer processes by construction.
-When these elements are retained, elastic scattering of electrons from thermal lattice distortions is not merely allowed but unavoidable.
-In this framework, pseudomomentum is exchanged between the electron and the lattice background, while total mechanical momentum is conserved through recoil of the lattice zero mode, without the necessity of phonon creation.
-The long-standing identification of lattice motion with inelasticity is thus revealed to be an artifact of an incomplete theoretical description.
-Restoring elastic momentum-transfer processes to their proper role provides a more complete and internally consistent foundation for electronic transport in crystals.
-It offers a natural interpretation of experimental observations that indicate robust elastic scattering in clean materials and clarifies how momentum relaxation can be efficient even when energy relaxation is slow.
-More broadly, these results raise the possibility that in at least some strange-metal regimes, momentum relaxation is dominated by elastic scattering in a time-dependent background, with genuinely inelastic processes playing a secondary role in establishing energy equilibration rather than setting the transport rate.
-IX Acknowledgements
+In the conventional clamped-lattice formulation of electron–phonon scattering, the lattice is treated as a fixed reference frame and its translational zero mode does not appear explicitly among the dynamical variables.
+Within that representation, momentum transfer is naturally described in terms of internal phonon excitations.
+When the lattice center-of-mass degree of freedom is retained explicitly, an additional elastic momentum-transfer channel becomes manifest.
+In a finite crystal, pseudomomentum can be exchanged between an electron and the lattice background while total mechanical momentum is conserved through recoil of the lattice zero mode.
+Such processes do not require phonon creation and may contribute significantly to momentum relaxation at low and intermediate temperatures.
+Making this channel explicit requires adopting a formulation in which (i) the lattice is treated as a finite, mechanically isolated object so that its center-of-mass motion remains dynamical, (ii) discrete translational symmetry is implemented symmetrically, with foreground and background pseudomomenta treated on equal footing, and (iii) the electron–lattice interaction is retained in its full density–density form.
+These steps do not alter the microscopic Hamiltonian; rather, they retain degrees of freedom that are commonly suppressed by choice of reference frame.
+When treated in this manner, elastic momentum exchange between electron and lattice background appears naturally within the same theoretical structure.
+In this framework, pseudomomentum is redistributed between the electron and the lattice background, while total mechanical momentum is conserved through recoil of the translational zero mode.
+Momentum relaxation need not coincide with phonon creation, and energy dissipation need not set the transport rate.
+This separation of roles clarifies how efficient momentum randomization can coexist with comparatively slow energy relaxation.
+Restoring explicit center-of-mass dynamics provides a more complete description of momentum bookkeeping in finite crystals.
+It offers a unified interpretation of experimental observations that indicate robust elastic scattering in clean materials, including weak localization, quantum oscillations, and ultrasonic attenuation.
+More broadly, the results suggest that in certain regimes, including some materials exhibiting linear-in-$T$ resistivity, momentum relaxation may be governed primarily by elastic scattering in a time-dependent lattice background [5], while inelastic processes play a secondary role in establishing global thermal equilibrium.
+In a perfectly clean, isolated, translation-invariant crystal, total momentum is conserved exactly.
+Such a system may exhibit internal equilibration and decay of electronic current, but it cannot exhibit finite DC resistivity in the strict steady-state sense.
+True resistivity requires a mechanism for removing total momentum from the system.
+IX Acknowledgments
 E.
@@
 DE-SC0025489.
-Appendix A Elastic and inelastic scattering in a tight-binding, isolated one-dimensional chain
-We consider a finite one-dimensional tight-binding chain of $N$ identical atoms, isolated in space, with nearest-neighbor springs and an electron moving in a basis of localized $p$ orbitals.
-The chain possesses an exact translational zero mode, corresponding to its center-of-mass (CM) motion.
-This allows the electron to exchange momentum with the lattice as a whole without necessarily exciting internal phonon modes.
-As a result, at finite temperature there exists a strictly elastic backscattering channel for an electronic wavepacket.
+Figure 3: Isolated systems of mass $M+m$ containing a foreground particle of mass $m$ and a background of mass $M$.
+The particle is attracted to the background and moves frictionlessly, except in C and D it may excite vibrations of the chain.
+In A, the particle is free except for walls.
+In B, the particle is in a fixed periodic potential within its host, the box.
+In C we replace the fixed corrugation with a chain of masses connected by springs.
+In D, we dispense with the box, coming close to the situation that matters here.
+Of course the box could be sitting on a table, which could be fixed to a floating laboratory, which could be upon something so massive that it has significant gravitational attraction, yet that too is floating in space…
+Appendix A Particle in a Free Box
+We consider simple models that clarify the treatment of momentum.
+The usual particle-in-a-box problem fixes the walls and therefore violates momentum conservation.
+Here the box is given finite mass and allowed to move freely, so that total momentum is exactly conserved.
+Consider a one-dimensional rigid box of mass $M$ and length $L$ with center-of-mass coordinate $R$, containing a particle of mass $m$ with coordinate $r$ (Fig. 3).
+Introduce total and relative coordinates
+$\mathcal{R}=\frac{mr+MR}{m+M},\qquad\rho=r-R,$
+with total mass $M_{T}=m+M$ and reduced mass
+$\mu=\frac{mM}{m+M}.$
+The hard-wall constraint is $|\rho|<L/2$.
+The Hamiltonian separates exactly:
+$H=-\frac{\hbar^{2}}{2M_{T}}\frac{\partial^{2}}{\partial\mathcal{R}^{2}}+\left[-\frac{\hbar^{2}}{2\mu}\frac{\partial^{2}}{\partial\rho^{2}}+V_{\rm hw}(\rho)\right].$
+Total momentum is conserved:
+$\hat{P}_{\rm tot}=-i\hbar\frac{\partial}{\partial\mathcal{R}}.$
+Case A: Flat interior
+The relative Hamiltonian is
+$H_{\rm rel}=-\frac{\hbar^{2}}{2\mu}\frac{d^{2}}{d\rho^{2}}+V_{\rm hw}(\rho).$
+The normalized eigenfunctions are
+$\phi_{n}(\rho)=\sqrt{\frac{2}{L}}\sin\!\left[\frac{n\pi}{L}\left(\rho+\frac{L}{2}\right)\right],\qquad n=1,2,\dots,$
+with energies
+$E_{n}=\frac{\hbar^{2}\pi^{2}n^{2}}{2\mu L^{2}}.$
+Any internal state may be written
+$\psi(\rho,t)=\sum_{n=1}^{\infty}c_{n}\,\phi_{n}(\rho)\,e^{-iE_{n}t/\hbar}.$
+A Gaussian superposition centered at $\rho_{0}$ with mean wavenumber $k_{0}$,
+$c_{n}=\mathcal{N}\exp\!\left[-\frac{(n-n_{0})^{2}}{4\sigma_{n}^{2}}\right]\exp\!\left[-ik_{n}\rho_{0}\right],\qquad k_{n}=\frac{n\pi}{L},$
+produces a localized packet
+$\psi(\rho,0)\approx A(\rho-\rho_{0})e^{ik_{0}(\rho-\rho_{0})}.$
+Although each $\phi_{n}$ carries no relative momentum, the particle and box momenta may be nonzero individually, while $P_{\rm tot}$ remains fixed.
+At a wall collision both reverse sign, preserving total momentum.
+Case B: Periodic interior
+Let the interior potential be periodic relative to the box:
+$V_{\rm per}(\rho)=V_{\rm per}(\rho+a).$
+Then
+$H_{\rm rel}=-\frac{\hbar^{2}}{2\mu}\frac{d^{2}}{d\rho^{2}}+V_{\rm per}(\rho)+V_{\rm hw}(\rho).$
+Eigenfunctions take Bloch standing-wave form
+$\varphi_{bn}(\rho)=\sqrt{\frac{2}{L}}u_{bn}(\rho)\sin\!\left(\frac{n\pi}{L}(\rho+L/2)\right),$
+with
+$u_{bn}(\rho+a)=u_{bn}(\rho).$
+Wavepackets are superpositions
+$\psi(\rho,0)=\sum_{n=1}^{\infty}A_{n}\varphi_{bn}(\rho).$
+Shared pseudomomentum
+Because $\rho=r-R$,
+$e^{iq\rho}=e^{iqr}e^{-iqR}.$
+Internal pseudomomentum $\hbar q$ is therefore shared between particle and box. For fixed total momentum $P_{0}$,
+$p_{r}=\frac{m}{m+M}P_{0}+\hbar q,\qquad P_{R}=\frac{M}{m+M}P_{0}-\hbar q,$
+so $p_{r}+P_{R}=P_{0}$.
+Cases C and D: Dynamical periodic background
+Let the periodic background be a dynamical elastic chain with atomic coordinates $R_{j}$:
+$H=\frac{p_{r}^{2}}{2m}+\sum_{j}\frac{p_{j}^{2}}{2M}+\frac{K}{2}\sum_{j}(R_{j+1}-R_{j}-a)^{2}+H_{\rm int}.$
+With density–density coupling
+$H_{\rm int}=\int dr\,dr^{\prime}\,\rho_{p}(r)V(r-r^{\prime})\rho_{\rm lat}(r^{\prime}),$
+where
+$\rho_{p}(r)=\delta(r-\hat{r}),\qquad\rho_{\rm lat}(r)=\sum_{j}\delta(r-\hat{R}_{j}),$
+the Hamiltonian depends only on coordinate differences and is invariant under rigid translation
+$r\rightarrow r+a_{0},\qquad R_{j}\rightarrow R_{j}+a_{0}.$
+Thus total momentum
+$P_{\rm tot}=p_{r}+\sum_{j}p_{j}$
+is exactly conserved. Diagonalizing the lattice,
+$R_{j}=R_{\rm lat}+\sum_{q}u_{q}e^{iqja},$
+reveals the zero mode $R_{\rm lat}$. The density operator contains the factor
+$e^{-iqR_{\rm lat}},$
+showing explicitly that momentum $\hbar q$ transferred to the particle is balanced by recoil of the lattice center of mass.
+Internal phonons redistribute momentum but are not required for momentum conservation.
+Appendix B Elastic and inelastic scattering in a tight-binding, isolated one-dimensional chain
+We consider a finite one-dimensional tight-binding chain of $N$ identical atoms, isolated in space, with nearest-neighbor springs and an electron moving in a basis of localized orbitals.
+The chain possesses an exact translational zero mode (center-of-mass motion).
+An electron can therefore exchange momentum with the lattice as a whole without necessarily exciting internal phonon modes.
+This opens a phonon-diagonal (internally elastic) backscattering channel for an electronic wavepacket, with the required momentum reversal balanced by center-of-mass recoil.
 Hamiltonian and separation of the CM mode
-Let the atomic coordinates be written as
+Write the atomic coordinates as
 $x_{n}=R+na+u_{n},\qquad\sum_{n=1}^{N}u_{n}=0,$ (41)
-where $R$ is the center-of-mass coordinate, $a$ the equilibrium lattice spacing, and $\{u_{n}\}$ the internal displacements. The lattice Hamiltonian separates exactly into CM and internal parts,
+where $R$ is the center-of-mass coordinate, $a$ the equilibrium lattice spacing, and $\{u_{n}\}$ the internal displacements. The lattice Hamiltonian separates into CM and internal parts,
 $H_{\rm lat}=\frac{P_{R}^{2}}{2M_{\rm tot}}+\sum_{n=1}^{N}\frac{p_{n}^{2}}{2M}+\frac{K}{2}\sum_{n=1}^{N-1}(u_{n+1}-u_{n})^{2},$ (42)
-with total mass $M_{\rm tot}=NM$ and $\sum_{n}p_{n}=0$.
+with $M_{\rm tot}=NM$ and $\sum_{n}p_{n}=0$.
 The electron is described by a nearest-neighbor tight-binding Hamiltonian,
 $H_{e}=-\sum_{n=1}^{N-1}t_{n}\left(c_{n+1}^{\dagger}c_{n}+c_{n}^{\dagger}c_{n+1}\right),$ (43)
-where the hopping amplitudes depend linearly on the bond lengths,
+with bond-length-dependent hopping
 $t_{n}\equiv t(x_{n+1}-x_{n})\simeq t_{0}-\alpha(u_{n+1}-u_{n}).$ (44)
-The electron-lattice coupling depends only on internal coordinates; the CM coordinate $R$ drops out exactly.
-Strictly elastic backscattering via CM recoil
-Because the full electron-lattice system is translationally invariant, there exists an exactly conserved total momentum.
-The electronic band energy satisfies $\varepsilon(k)=\varepsilon(-k)$, so reversing the electronic momentum does not change the electronic energy.
-Consider an incoming electronic wavepacket narrowly peaked at momentum +k+k, and an internal lattice eigenstate $|\nu\rangle$. A scattering process of the form
-$|k\rangle\otimes|\nu\rangle\otimes|P_{R}\rangle\;\longrightarrow\;|-k\rangle\otimes|\nu\rangle\otimes|P_{R}+2\hbar k\rangle$ (45)
-leaves the internal phonon state unchanged and transfers the momentum $2\hbar k$ entirely to the CM of the chain.
+Thus the electron–lattice coupling depends only on internal coordinates $u_{n+1}-u_{n}$; the CM coordinate $R$ drops out exactly.
+Phonon-diagonal backscattering via CM recoil
+Because the full electron+lattice Hamiltonian is translationally invariant, the total momentum is exactly conserved.
+Since the band energy satisfies $\varepsilon(k)=\varepsilon(-k)$, electronic reversal $k\to-k$ does not change the electronic energy.
+Consider an incoming wavepacket narrowly peaked at +k+k, an internal lattice eigenstate $|\nu\rangle$, and a CM momentum eigenstate $|P_{R}\rangle$.
+A phonon-diagonal backscattering channel has the form
+$|k\rangle\otimes|\nu\rangle\otimes|P_{R}\rangle\longrightarrow|-k\rangle\otimes|\nu\rangle\otimes|P_{R}+2\hbar k\rangle,$ (45)
+i.e. the internal state is unchanged while the CM absorbs the recoil $2\hbar k$.
 The associated change in CM kinetic energy is
 $\Delta E_{\rm CM}=\frac{(P_{R}+2\hbar k)^{2}-P_{R}^{2}}{2M_{\rm tot}}=\frac{4\hbar k\,P_{R}+4\hbar^{2}k^{2}}{2M_{\rm tot}}.$ (46)
-Exact energy conservation requires $\Delta E_{\rm CM}=0$, which is satisfied when
+Exact energy conservation with strictly unchanged internal state requires $\Delta E_{\rm CM}=0$, i.e.
 $P_{R}=-\hbar k.$ (47)
-For this value of the CM momentum, the backscattering process $k\rightarrow-k$ is strictly elastic, involving no excitation or absorption of internal phonon modes.
-Alternately, if $N$ is very large, the requirement $P_{R}=-\hbar k$ can be relaxed.
-For a finite, free chain of total mass $M$, electron backscattering $k\rightarrow-k$ changes the electron momentum by $\Delta p_{e}=-2\hbar k$.
-Momentum conservation then requires the chain to absorb a recoil $\Delta P_{\rm chain}=+2\hbar k$.
-If the chain initially has CM momentum $P$, energy conservation implies an associated recoil energy
-$\Delta E_{\rm recoil}=\frac{(P+2\hbar k)^{2}-P^{2}}{2M}=\frac{2\hbar k\,(P+\hbar k)}{M}.$ (48)
-Strictly elastic backscattering without internal excitation is therefore possible only for the special value $P=-\hbar k$; for generic initial CM momentum, an energy mismatch must be accommodated.
-For a macroscopic chain, however, $M\propto N$ and the recoil energy scales as $\Delta E_{\rm recoil}\sim 1/N$, becoming parametrically small in the thermodynamic limit.
-In this regime the energy mismatch can be absorbed without resolving a discrete internal excitation.
-In particular, an electron wavepacket of spatial extent $L$ has an intrinsic energy spread
-$\Delta E_{\rm pkt}\sim\hbar v_{g}\,\Delta k\sim\hbar v_{g}/L,$ (49)
-which is independent of $N$ and typically exceeds $\Delta E_{\rm recoil}$ for sufficiently large chains.
-Equivalently, if the scattering occurs over a finite time $\tau_{\rm scatt}$, the associated time-energy uncertainty $\hbar/\tau_{\rm scatt}$ exceeds the recoil energy, rendering the process effectively elastic from the electron’s perspective.
-In a phonon number-state description, the accommodation of this small energy mismatch corresponds to the creation or annihilation of one or more extremely low-frequency acoustic phonons.
-Importantly, however, such number-state changes do not account for the electron’s momentum reversal $k\rightarrow-k$: a single low-$q$ phonon carries momentum $\hbar q\ll 2\hbar k$.
-The dominant momentum transfer is instead absorbed by the collective motion of the chain (center-of-mass recoil), while the phonon sector merely provides an energetically soft channel required by exact energy conservation.
-Thus, for large $N$, electron backscattering involves an exact momentum transfer of $2\hbar k$ to the background with a vanishing energy cost, and any phonon number-state changes reflect quasi-elastic energy bookkeeping rather than the mechanism of momentum reversal itself.
+For generic $P_{R}$ there is a recoil-energy mismatch
+$\Delta E_{\rm recoil}=\frac{(P_{R}+2\hbar k)^{2}-P_{R}^{2}}{2M_{\rm tot}}=\frac{2\hbar k\,(P_{R}+\hbar k)}{M_{\rm tot}}.$ (48)
+However, for a macroscopic chain $M_{\rm tot}\propto N$ so $|\Delta E_{\rm recoil}|\sim 1/N$ becomes parametrically small.
+For a wavepacket of spatial extent $L$ with intrinsic energy width $\Delta E_{\rm pkt}\sim\hbar v_{g}/L$, or for scattering occurring over a finite time $\tau_{\rm scatt}$ with uncertainty $\hbar/\tau_{\rm scatt}$, the process is effectively elastic whenever $\Delta E_{\rm recoil}$ is below the available energy resolution.
+In a number-state phonon description, any residual mismatch may be accommodated by very soft internal modes, but the momentum reversal $k\to-k$ is carried by CM recoil; internal excitations, if present, serve primarily as an energetically low-cost bookkeeping channel.
 Role of finite temperature
@@
 with width $\sigma_{P}=\sqrt{M_{\rm tot}k_{B}T}$.
-The resonant value $P_{R}=-\hbar k$ therefore occurs with nonzero probability density for any $T>0$.
-Consequently, strictly elastic electronic backscattering events exist at finite temperature, mediated entirely by recoil of the lattice center of mass.
-Bulk and boundary backscattering in a isolated chain
-The elastic CM-recoil backscattering channel does not require the electron to encounter the ends of the chain and may occur locally in the bulk.
-The electron-lattice interaction in Eq.
-(44) depends only on nearest-neighbor bond distortions and is independent of absolute position.
-For a wavepacket localized well away from the ends, with spatial extent $\ell\ll Na$, the matrix element connecting $|k\rangle$ and $|-k\rangle$ takes the form
-$\langle-k,P_{R}+2\hbar k|H_{\rm int}|k,P_{R}\rangle\propto\sum_{n}e^{2ikna}\,\langle P_{R}+2\hbar k|P_{R}\rangle,$ (51)
-where the sum runs over bulk sites only. The momentum transfer is absorbed entirely by the CM degree of freedom, while the internal lattice state remains unchanged.
-Free boundaries of the finite chain can also reflect an electronic wavepacket once it reaches an end.
-Such boundary reflections are elastic and likewise involve a momentum reversal $k\rightarrow-k$ accompanied by CM recoil.
-Free boundaries therefore provide a spatial location for reflection but do not act as a momentum sink.
-Both bulk and boundary backscattering processes are governed by the same momentum conservation principle enforced by the CM zero mode.
+The resonant value $P_{R}=-\hbar k$ therefore occurs with nonzero probability density for any $T>0$, so strictly phonon-diagonal backscattering events exist at finite temperature in an isolated chain.
+Bulk and boundary backscattering
+The CM-recoil channel does not require the electron to reach the ends: it can occur in the bulk from the static component of thermal distortions (phonon-diagonal scattering). The SSH coupling
+$H_{\rm int}=\alpha\sum_{n=1}^{N-1}(u_{n+1}-u_{n})\left(c_{n+1}^{\dagger}c_{n}+c_{n}^{\dagger}c_{n+1}\right)$
+connects $|k\rangle$ and $|-k\rangle$ through the $2k$ Fourier component of the bond-distortion field. CM recoil is enforced by the CM translation factor. In particular, the CM shift operator produces
+$\langle P_{R}^{\prime}|e^{-i(2k)R}|P_{R}\rangle\propto\delta(P_{R}^{\prime}-P_{R}-2\hbar k),$ (50)
+so momentum conservation is implemented by a delta function (or Kronecker delta in a finite quantization volume), not by an inner product between distinct CM momentum eigenstates.
+Free boundaries can also reflect a wavepacket once it reaches an end.
+Such reflections reverse $k$ and are likewise accompanied by CM recoil.
+Boundaries provide a location for reflection but do not constitute a momentum sink; both bulk (distortion-mediated) and boundary reflections obey the same conservation principle enforced by the CM zero mode.
 Other scattering channels
-In addition to the strictly elastic CM recoil channel, a variety of inelastic and hybrid scattering processes are allowed within a isolated chain.
+In addition to the phonon-diagonal CM-recoil channel, inelastic and hybrid processes are allowed.
 Single-phonon emission and absorption
-If momentum is transferred to internal phonon modes, the internal lattice state changes. A generic single-phonon process takes the form
-$|k\rangle\otimes|\{n_{q}\}\rangle\longrightarrow|k^{\prime}\rangle\otimes|\{n_{q}^{\prime}\}\rangle,$
-with $n_{q}^{\prime}=n_{q}\pm 1$. Energy conservation requires
+If momentum is transferred to internal modes, the internal lattice state changes:
+$|k\rangle\otimes|\{n_{q}\}\rangle\to|k^{\prime}\rangle\otimes|\{n_{q}^{\prime}\}\rangle,\qquad n_{q}^{\prime}=n_{q}\pm 1.$
+Energy conservation requires
 $\varepsilon(k^{\prime})=\varepsilon(k)\pm\hbar\omega_{q}+\Delta E_{\rm CM}.$
-The corresponding Golden-Rule rate is
+A Golden-Rule rate has the form
 $W_{k\to k^{\prime};q}=\frac{2\pi}{\hbar}|M_{k,k^{\prime};q}|^{2}\,\delta\!\left(\varepsilon(k^{\prime})-\varepsilon(k)\mp\hbar\omega_{q}-\Delta E_{\rm CM}\right),$
@@
 $M_{k,k^{\prime};q}\sim\alpha\sqrt{\frac{\hbar}{2M\omega_{q}}}\,F_{k,k^{\prime};q}.$
-Multi-phonon and hybrid CM-phonon processes
-When single-phonon processes are kinematically suppressed, higher-order multi-phonon processes and hybrid processes involving both CM recoil and internal phonon exchange are allowed.
-These channels relax the strict resonance condition Eq.
-(47) at the cost of reduced matrix elements.
-Kinematic summary and temperature dependence
-  • Elastic channels: exact backscattering $k\rightarrow-k$ without phonon excitation occurs via CM recoil and is enabled by translational invariance.
-  • Inelastic channels: internal phonon emission and absorption become increasingly important as temperature increases and as phonon spectra permit energy matching.
-  • Finite size: the discreteness of the phonon spectrum in a finite chain strongly constrains available inelastic channels and can enhance the relative importance of CM-mediated elastic processes.
+Multi-phonon and hybrid CM–phonon processes
+Higher-order multi-phonon processes and hybrid processes involving both CM recoil and internal phonon exchange relax the strict resonance condition (47) at the cost of reduced matrix elements.
+Kinematic summary
+  • Phonon-diagonal channel: $k\to-k$ with unchanged internal state is allowed by translational invariance and proceeds via CM recoil; it is strictly elastic only on resonance (47) and effectively elastic for large $N$.
+  • Inelastic channels: internal phonon emission/absorption become important when the phonon spectrum permits energy matching.
+  • Finite size: discrete internal spectra constrain inelastic channels and can enhance the relative importance of CM-mediated phonon-diagonal scattering.
 Practical consequences
-  • In wavepacket simulations, elastic backscattering events are identified by momentum reversal without change in internal phonon occupation, accompanied by a compensating shift in CM momentum.
-  • Both bulk and boundary reflections obey the same momentum conservation law and are distinguished only by their spatial location.
-  • The CM zero mode plays a central dynamical role and cannot be neglected without qualitatively altering the scattering physics.
+  • In wavepacket simulations, phonon-diagonal backscattering is identified by momentum reversal with unchanged internal state and a compensating shift of $P_{R}$ by $2\hbar k$.
+  • Bulk and boundary reflections differ only in spatial location; both satisfy the same momentum conservation enforced by the CM zero mode.
+  • Neglecting the CM zero mode (implicit clamping) qualitatively changes the available scattering channels.
+Appendix C Whither Umklapp?
+The traditional role of Umklapp for finite resistivity in an infinite, perfect crystal arises from freezing the background and projecting out its pseudomomentum.
+Once the background is treated as dynamical, with an explicit translational degree of freedom, momentum relaxation can occur through elastic pseudomomentum exchange between electron and lattice.
+Phonons then represent one possible basis for describing background dynamics rather than the exclusive channel for momentum transfer.
+This perspective clarifies long-standing ambiguities surrounding Umklapp and explains why momentum relaxation is observed experimentally even in regimes where Umklapp processes are expected to be suppressed.
+It also underscores that the association of Umklapp with dissipation and entropy production is not intrinsic, but follows from additional assumptions about the background.
+The expectation that Umklapp processes are necessary for finite resistivity in a perfect crystal has deep roots in transport theory.
+Peierls first distinguished between normal and Umklapp processes in his analysis of phonon thermal conductivity, showing that only the latter can relax total crystal momentum in a closed system.
+This reasoning was later extended to electrons by Bloch, Peierls, Ziman, and others, and appears explicitly in standard texts (e.g.
+Peierls, Quantum Theory of Solids; Ziman, Electrons and Phonons).
+If the Fermi surface lies well inside the Brillouin zone, or at low temperature where large-wavevector phonons are thermally suppressed, Umklapp processes are expected to be exponentially rare.
+Under these conditions conventional theory predicts extremely long momentum lifetimes in a clean crystal.
+The present formulation revisits this conclusion by restoring the dynamical translational degree of freedom of the background.
+Momentum exchange between subsystems does not require phonon creation, and the strict identification of Umklapp with momentum relaxation relies on additional assumptions about background clamping or environmental coupling.
+Umklapp processes are best understood as a classification of crystal-momentum conservation rather than as the fundamental cause of momentum relaxation.
+Their elevated role in conventional transport theory stems from treating the lattice background as inert.
+Recognizing the background as a dynamical object carrying its own pseudomomentum restores a more symmetric and physically complete description in which elastic momentum exchange is allowed and Umklapp is no longer indispensable.

### v3

@@
-Elastic, Quasielastic, and Superelastic Electron Scattering from Thermal Lattice Distortions in Perfect Crystals
+Elastic and Quasielastic Electron Scattering in Perfect Crystals
 Abstract
-In standard treatments of electron transport, momentum relaxation in a perfect, defect-free crystal is linked with phonon creation or annihilation.
-In this work, we reconsider this problem for a finite, isolated crystal, retaining the lattice center-of-mass (recoil) degree of freedom and enforcing conservation of total mechanical momentum together with discrete crystal pseudomomentum.
-Starting from the density–density form of the electron–lattice interaction, we show that an electron in the interior of a perfect crystal admits elastic momentum-transfer channels in which total momentum is conserved by recoil of the lattice background without phonon excitation.
-These elastic channels can provide the leading contribution to momentum relaxation.
-We further identify mixed quasielastic and superelastic processes in which phonon occupations change but do not account entirely for the electron’s momentum transfer.
-The elastic channels arise within the standard microscopic Hamiltonian and do not require additional disorder or defects.
-The resulting framework provides a complementary microscopic perspective on momentum relaxation in clean crystals, and is consistent with experimental phenomena such as weak localization, quantum oscillations, ultrasonic attenuation, and the observed separation of momentum and energy relaxation times.
+Momentum relaxation of electrons in perfect crystals is conventionally associated with phonon emission or absorption and is therefore regarded as intrinsically inelastic.
+We show that this is not a fundamental requirement.
+Starting from the Fröhlich Hamiltonian, but making fewer assumptions than usual—retaining total mechanical momentum and the full density-density electron-lattice interaction—we find a substantially larger class of allowed processes.
+When total momentum conservation is enforced together with all pseudomomenta, the lattice center-of-mass provides a continuous momentum reservoir, and phonons are no longer required to account for the electron’s momentum change.
+The conventional phonon-mediated processes emerge as a restricted subset of a higher-dimensional manifold of transitions.
+Within this expanded space, elastic and quasielastic momentum-transfer channels appear already at first order, even in a perfect crystal.
+Momentum relaxation can therefore proceed rapidly while energy relaxation remains slow, consistent with a wide range of experimental observations including hot-electron injection, quantum oscillations, weak localization, and Coulomb drag in low-dimensional conductors.
+In particular, scattering in a perfect one-dimensional wire is not kinematically constrained once the lattice recoil degree of freedom is retained.
 preprint: APS/123-QED
@@
 They govern both momentum relaxation and energy exchange in crystalline solids, and therefore underlie electrical resistivity and the loss of electronic coherence.
-Experimentally, however, momentum and energy relaxation often occur on markedly different timescales.
+In the modern transport theory of metals (Bloch–Peierls–Ziman–Mott–Ashcroft–Mermin), the electron–lattice problem is formulated by imposing periodic boundary conditions and expanding lattice displacements using internal normal modes.
+In this construction, going back to Born and von K’arm’an, the lattice center-of-mass (zero) mode is removed at the outset, leaving only internal vibrational degrees of freedom.
+When an electron applies force to a lattice atom, it can only respond with finite energy phonons.
+The zero energy recoil, without which Mössbauer scattering would falter, has been removed.
+In first order, momentum transfer is forced to reside in a single phonon, leading to the familiar selection rules $k’=k\pm q$ and $\varepsilon_{k’}=\varepsilon_{k}\pm\hbar\omega_{q}$.
+In a perfect crystal, a change in electron momentum thus requires phonon emission or absorption and is therefore inelastic.
+Although this is true within the formulation, it is an unnatural result, not followed in the laboratory.
+When the lattice center-of-mass degree of freedom is retained, the total pseudomomentum is conserved and can be exchanged between the electron and the background.
+If the zero mode is removed, this exchange channel is absent.
+The phonon sector alone does not form a closed system for pseudomomentum, and the familiar selection rule
+$\Delta K_{e}+\Delta K_{\rm ph}=Gm$
+need not hold for individual processes.
+Indeed, it is violated once the density-density interaction is treated beyond the linear Fröhlich approximation.
+Pseudomomentum is therefore not a conserved quantity within the electron-phonon subspace alone.
+The traditional clean pseudomomentum accounting falters if even the next terms are kept in the the density-density expansion, e.g.
+Purely elastic processes are a rarity in the classical world, but are commonplace in the quantum world.
+Consider a linear atom-diatom collision, i.e.
+three translational degrees of freedom, separating to a center mass of the diatom, the atom, and the internal virbatipnal coordinate: if not too violent, upon measurement the diatom is likely to be in the state it started in, even though it has definitely been disturbed.
+Again, Mössbauer emission comes to mind.
+In fact for pass-through experiments, like fast electrons, both Bragg and a good portion of diffuse scattering is known to be elastic, with the crystal recoiling as a whole.
+Somehow an electron residing inside a crystal has found itself treated a different way, but it should not be.
+Experimentally, momentum and energy relaxation often occur on markedly different timescales.
 In clean metals over broad temperature ranges, electrical resistivity and magnetotransport measurements indicate rapid momentum randomization, while hot-electron experiments, Johnson-noise thermometry, and related nonequilibrium probes show that energy relaxation to the lattice can remain comparatively slow.
@@
 where $\tau_{p}$ is the momentum-relaxation time and $\tau_{E}$ the energy-relaxation time, indicates that electronic momentum can be efficiently scrambled without substantial net energy transfer to internal lattice excitations.
-In conventional transport theory, momentum relaxation in a perfect, defect-free crystal is typically described in terms of phonon-mediated scattering processes.
-Within the Bloch–Grüneisen and semiclassical Boltzmann frameworks, transport is treated as a sequence of transitions between electronic states, with rates evaluated using the electron–phonon interaction [35, 3].
-In the standard Fröhlich Hamiltonian [9], the interaction is expressed explicitly in terms of phonon creation and annihilation operators, and momentum transfer is therefore associated, at lowest order, with changes in phonon occupation.
-In this commonly employed description, elastic momentum relaxation is primarily attributed to impurities or other forms of explicit disorder.
-The observed separation between $\tau_{p}$ and $\tau_{E}$ motivates a careful examination of how momentum conservation is implemented microscopically in a perfectly periodic crystal.
-The conceptual distinction between momentum relaxation and energy dissipation has long been recognized.
-Peierls emphasized that electrical resistance requires degradation of electronic (crystal) momentum rather than dissipation of electronic energy, and noted that the lattice, as a massive collective object, can absorb momentum with minimal energy cost [24].
-Pippard similarly stressed that while defects often play an important role, momentum relaxation need not be uniquely tied to strongly inelastic processes [25].
-In this work we revisit the problem starting from the same microscopic electron–lattice Hamiltonian, but retaining explicitly the lattice center-of-mass degree of freedom.
-In conventional treatments the crystal is effectively considered as a fixed reference frame, and the translational zero mode does not appear among the dynamical variables.
-Here we adopt an expanded viewpoint in which the crystal total mechanical momentum is retained.
-This does not modify the underlying Hamiltonian; rather, it makes explicit a degree of freedom that is commonly suppressed.
-When total momentum conservation is enforced in this manner, a transfer of momentum $\mathbf{q}$ to an electron is accompanied by an equal and opposite momentum $\mathbf{q}$ carried by the remaining degrees of freedom, which we identify as the lattice background.
-This exchange can occur without excitation of internal phonon modes, leading to an elastic recoil channel even in a perfectly periodic crystal.
-We also identify mixed quasielastic and superelastic processes in which phonon occupations change but do not fully account for the electron’s momentum transfer.
-These channels arise within the standard density–density form of the electron–lattice interaction [22, 9] and require no additional disorder.
-The physical content of the elastic recoil channel is elementary.
-In any untethered composite object, an impulse imparted to one constituent immediately changes the total momentum of the whole.
-While redistribution of that momentum among internal degrees of freedom proceeds causally and at finite velocity, the collective momentum of the backgroud crystal appears in the wavefunction and is not dependent of the details of redistribution of the momentum among the constitutuents of the whole.
-.
-An instructive analogue is provided by Mössbauer scattering, in which the dominant elastic line is recoil of the entire lattice without internal phonon excitation.
-A conceptual roadblock has been that, in contrast to processes such as Mössbauer and neutron scattering, where elastic, non-Bragg diffuse deflection is widely acknowledged, conduction electrons do not enter and exit; they start and stay in the crystal.
-The roadblock is addressed by assigning pseudomomentum to both a “foreground” electron density and a “background” consisting of everything else.
-We hasten to point out that the question here involves real physics, and not mere representation: Must phonons be created or destroyed when electrons deflect in a pure crystal, or not?
-On the meaning of elastic scattering
-Elastic scattering of interior electrons is most familiar in the context of impurity or defect scattering, where translational symmetry is explicitly broken.
-Here, we restore pseudomomentum to the status of a conserved momentum, and as in free space what then is conserved in such an elastic deflection?
-The answer will not be controversial: the total momentum of the whole system is conserved.
-What we believe is new is that two pseudmomomenta have changed equally and oppositely in such a collision: one belonging to the electron, another to the rest of the crystal.
-The total momentum, the sum of the two, is unchanged.
-In a perfectly periodic crystal, the relationship between changes in electronic pseudomomentum and the mechanical momentum of the lattice background is less often made explicit.
-In the present formulation this bookkeeping becomes transparent once (i) the lattice is treated as fully dynamical, (ii) pseudomomentum is assigned both to the electron and to the background in the absence of the electron density, and (iii) the interaction is expressed in density–density form, making the exchange of pseudomomentum explicit.
-Finally, we clarify our terminology.
-In a finite-temperature many-body system, strictly elastic processes do not exist in an absolute sense.
-Even the dominant line in Mössbauer scattering exhibits slight thermal broadening.
-We therefore use the term elastic to denote processes in which phonon occupation does not change, and quasielastic to denote processes involving energy exchange that is small compared with the momentum transfer involved.
-The landscape seen by an electron
-At temperatures of order $T\sim 100\,$K, thermally excited acoustic distortions in a typical metal generate substantial deformation-potential landscapes on nanometer length scales.
-Using standard values for the deformation potential ($D\sim 5$$15\,\mathrm{eV}$), sound velocity ($v_{s}\sim 3$$5\times 10^{3}\,\mathrm{m/s}$), and the characteristic thermal phonon wavevector $q_{T}\sim k_{B}T/\hbar v_{s}$, one finds order-of-magnitude estimates for the resulting deformation-potential field gradients of
-$|\nabla V_{\rm def}|\sim 10^{4}\text{–}10^{6}\ \mathrm{V/cm}\qquad(T\sim 100\,\mathrm{K}).$ (2)
-Over a representative lateral scale of $L\sim 10\,$nm, this corresponds to peak-to-peak variations of the screened scalar potential of order
-$\Delta V_{\rm pp}\sim|\nabla V_{\rm def}|\,L\sim 10\text{–}300\ \mathrm{mV}.$ (3)
-[…]
-[…]
-II Foreground, Background, and Pseudomomentum
-The conceptual distinction between momentum, crystal momentum, and electrical current has been recognized since the early development of band theory, but its implications for microscopic transport have remained subtle.
-In a particularly prescient analysis, Kohn emphasized that crystal momentum is a conserved quantity associated with translational symmetry, while the electrical current need not be proportional to it, and that momentum may be exchanged with the lattice without the necessity of phonon excitation or energy dissipation [17].
-This observation already points to the possibility that momentum conservation can be satisfied through recoil of the lattice background rather than through inelastic scattering processes.
-What has remained largely implicit, however, is how this distinction should be implemented at the level of a local scattering event occurring deep in the interior of a crystal.
-Here, we explicitly assign pseudomomentum to both a foreground electron and the background lattice, with both pseudomomenta appearing as dynamical variables in the many-body wavefunction.
-Making this partition explicit is a realization of Kohn’s insight, within a framework in which elastic deflection, coherent lattice recoil, and diffusive transport can coexist naturally in a perfect crystal.
-In what follows, we formalize a symmetric foreground-background structure, derive the elastic scattering amplitude within first-order perturbation theory, and compute the elastic fraction explicitly.
-We show that this fraction is close to unity for copper at room temperature.
-We also clarify how the present formulation extends the Fröhlich framework without contradicting its successful predictions, and discuss the implications for the interpretation of transport phenomena in crystalline solids.
-We distinguish explicitly between foreground and background degrees of freedom.
-The foreground is the object of interest, here a single electron or a collection of them.
+The experimentally observed separation between $\tau_{p}$ and $\tau_{E}$ does not seem to comport with the “always inelastic” narrative, and this reality check motivates a careful examination of how momentum conservation is implemented in a perfectly periodic crystal.
+At finite temperature, lattice vibrations introduce perturbations that allow electrons to scatter between Bloch states.
+This mechanism was first analyzed by Bloch, who obtained the characteristic low-temperature $T^{5}$ resistivity associated with electron-phonon scattering in metals [8].
+These ideas were incorporated into the Bloch-Grüneisen-Boltzmann transport framework [8, 13], in which the electronic distribution function evolves according to a Boltzmann equation for Bloch electrons.
+Within this formulation, scattering is described as transitions between Bloch states accompanied by phonon absorption or emission, enforced by energy-conserving delta functions.
+In the standard Fröhlich Hamiltonian [10], the interaction is expressed in terms of phonon creation and annihilation operators, so that momentum transfer is associated, at lowest order, with changes in phonon occupation and is therefore intrinsically inelastic, subject to pseudomomentum and energy conservation [39, 4].
+Elastic momentum relaxation is accordingly attributed to impurities or other forms of explicit disorder.
+The subsequent development of Migdal-Eliashberg theory further reinforced this picture, embedding the identification of momentum relaxation with inelastic electron-phonon scattering into the standard theoretical framework.
+The situation has been different for external, “pass through” probes such as X-rays or neutrons or even fast electrons.
+In such circumstances, elastic and even diffuse elastic recoil of the entire crystal is routinely included through Debye-Waller theory.
+Here, we will see that there is no solid basis for the asymmetry in the treatment of an interior electron.
+As we will see, bringing back the center of mass dramatically changes the narrative of internal scattering processes.
+Somehow that seems implausible, which is one reason perhaps that the inelastic narrative has persisted for so long.
+We revisit this issue determined to be rigorous about momentum conservation in a fully isolated electron-crystal system with all applicable momenta retained.
+(The isolation can later be broken once a full understanding is reached of the isolated system.) When the electron-lattice interaction is written in its full density-density form, elastic and quasi-elastic momentum exchange channels appear already at first order, without any defects present other than the transient ones implied by the thermal deformation potential.
+A few preliminary, qualitative remarks are in order.
+First, removing the center of mass makes electrons in crystals is akin to the venerable particle in a box model.
+There too there is no center of mass of the box, and no box recoil.
+Momentum is not conserved.
+In the Appendix we show what happens when the box is isolated, given mass, and allowed to recoil.
+Second, if the center of mass is absent, the only degree of freedom that could carry the recoil resulting from an elastic interior electron deflection is also absent.
+Phonons will be forced to take on that role and indeed by construction there can be no elastic events of this sort.
+If the crystal total mechanical momentum is retained, and total momentum conservation is again enforceable, a transfer of momentum $\mathbf{q}$ to an electron is accompanied by an equal and opposite momentum $\mathbf{q}$ carried by the remaining degrees of freedom, which we identify as the lattice background.
+Those degrees of freedom include the several lattice zero frequency recoil modes, and internal vibrational modes.
+An instructive analog is provided by Mössbauer scattering, in which the dominant elastic line, perhaps 97% of the total probability, implies recoil of the entire lattice without internal phonon excitation.
+We hasten to point out that the question here involves real physics, and not mere representation: are phonons created or destroyed when electrons deflect in a pure crystal, or not?
+That is an essential, physical difference, with profound experimental consequences.
+In a finite-temperature many-body system, strictly elastic processes do not exist in a provable, absolute sense.
+Even the dominant “elastic” line in Mössbauer scattering exhibits slight thermal broadening.
+We therefore use the term elastic to denote processes in which phonon occupation does not change, and quasielastic to denote processes involving small energy exchanges that do not balance electron momentum transfer; that task will be left to the crystal center of mass recoil.
+In fact there are so many ways that momentum conservation can be accommodated that the textbook phonon carrying away exactly the right momentum is a measure zero event, in spite of its utility as a model.
+This will become clearer below.
+II A minimal example: the electron on a free atomic ring
+Figure 1: Left: Ring lattice with a localized phonon excitation.
+All angular coordinates are retained, including the 0 mode (angular momentum).
+Right: The tight-binding electronic wavefunction is added.
+The color wheel represents the amplitude and phase of the electronic amplitude $e^{ikx}$, winding here by $4\pi$ around the ring.
+Only total angular momentum $L_{\rm tot}=L+R\hbar k$.
+and energy are conserved quantities.
+Before turning to the general many-body problem, it is useful to display the dramatic effect of including the zero mode in the simplest finite crystal that has exact momentum bookkeeping: a ring of $N$ identical atoms of mass $m$, constrained to move on a circle of radius $R$ and coupled by nearest-neighbor springs (figure 1, left).
+Completing the picture, an electron moves on the same ring by dynamic tight binding nearest-neighbor hopping (figure 1, right).
+Dynamic means the atoms move along the ring, unconstrained except for their neighbors, and carry their orbitals with them.
+An early example of this approach applied to graphene is found in ref.
+[28].
+The importance of this model is its simplicity, yet it retains essential electron-phonon interactions.
+The lattice zero mode is present explicitly, total angular momentum is conserved, and the distinction between internal phonons and rigid-body recoil is completely transparent.
+All lattice degrees of freedom are retained, including the global rotational (zero) mode, which carries the total angular momentum of the lattice.
+The standard conclusion for such a 1D model, like a perfect wire, is that only elastic backscattering is allowed, and even then under restricted circumstances, due to the simultaneous constraints of energy and momentum conservation when only the electron and a phonon are present.
+Efficient relaxation requires an additional degree of freedom.
+In conventional treatments[26, 18] this is introduced in the form of a third electronic excitation, such as a deep hole drawn from the Fermi sea to satisfy the combined energy and momentum constraints that cannot be met within a strictly two-body electron-phonon process.
+Multiphonon processes may also be invoked, but at higher order.
+An especially clean case is provided by Coulomb drag[24] in wire pairs, where exponential suppression of drag with temperature is again predicted, but instead power law dependence is seen in experiments.
+We will find something completely different, with manifolds of transitions possible obeying all conserved quantities.
+The third, “new” degree of freedom turns out to be the long lost center of mass.
+In drag, current in one wire creates fluctuating electric fields that transfer momentum to the other wire through the Coulomb interaction.
+For that transfer to occur, both wires must be able to absorb the same small energy $\hbar\omega$ and momentum $q$.
+The drag rate is therefore governed by the overlap of the two wires’ low-energy dynamical response functions.
+If each wire has gapless excitations near the Fermi points, then as $T\to 0$ there remains a shrinking but nonzero window of allowed $\omega$ and $q$, typically of size set by $k_{B}T$.
+Integrating over that thermal window gives a power of $T$.
+II.1 Hamiltonian and collective coordinates
+Let $\phi$ denote the angular coordinate of the electron, and let $\theta_{\ell}$ denote the instantaneous angular position of atom $\ell=1,\dots,N$ on the ring.
+The atomic coordinates are dynamical variables and include thermal displacements.
+For each instantaneous lattice configuration $\{\theta_{\ell}\}$, define a set of localized electronic orbitals
+$\chi_{\ell}(\phi;\theta_{\ell})=\chi(\phi-\theta_{\ell}),$ (2)
+where $\chi(\phi-\theta_{\ell})$ is an electron wavefunction centered on the nucleus at $\theta_{\ell}$. A general single-electron state may then be written as
+$\Psi(\phi;\{\theta_{\ell}\})=\sum_{\ell=1}^{N}\psi_{\ell}\,\chi(\phi-\theta_{\ell}),$ (3)
+where the coefficients $\psi_{\ell}$ give the electronic amplitudes on the lattice-following orbitals.
+The full Hamiltonian acts both on the lattice coordinates $\{\theta_{\ell}\}$ and on the electronic amplitudes $\{\psi_{\ell}\}$:
+$H=H_{\rm lat}(\{\theta_{\ell}\})+H_{\rm el}(\{\theta_{\ell}\}),$ (4)
+Here $H_{\rm el}(\{\theta_{\ell}\})$ denotes the electronic Hamiltonian in the lattice-following orbital basis.
+It encodes the electron-lattice interaction.
+Its matrix elements depend on the instantaneous atomic positions and therefore include both the electronic energies and the electron-lattice interaction.
+$H_{\rm lat}$ $=-\frac{\hbar^{2}}{2I_{\rm ion}}\sum_{\ell=1}^{N}\frac{\partial^{2}}{\partial\theta_{\ell}^{2}}+\frac{K}{2}\sum_{\ell=1}^{N}\left(\theta_{\ell+1}-\theta_{\ell}-\tfrac{2\pi}{N}\right)^{2},$ (5)
+$\qquad\theta_{N+1}\equiv\theta_{1}+2\pi.$
+In the nearest-neighbor tight-binding approximation, the electronic Hamiltonian acts on the coefficients $\psi_{\ell}$ as
+$(H_{e}\psi)_{\ell}=-\,t_{\ell-1}\,\psi_{\ell-1}-\,t_{\ell}\,\psi_{\ell+1},\qquad\psi_{\ell+N}\equiv\psi_{\ell},$ (6)
+where the hopping amplitudes $t_{\ell}$ are determined by the overlap of neighboring orbitals and therefore depend on the instantaneous bond lengths, i.e.
+on the differences $\theta_{\ell+1}-\theta_{\ell}$.
+II.2 Lattice: zero mode and internal phonons
+Let $\theta_{\ell}$ be the angular position of atom $\ell=1,\dots,N$, with periodic indexing $\theta_{N+1}\equiv\theta_{1}$. Introduce the center angle
+$\Theta=\frac{1}{N}\sum_{\ell=1}^{N}\theta_{\ell}$ (7)
+and internal coordinates
+$\varphi_{\ell}=\theta_{\ell}-\Theta,\qquad\sum_{\ell=1}^{N}\varphi_{\ell}=0.$ (8)
+The angular coordinates $\theta_{\ell}$ denote the instantaneous positions of the atoms and include thermal displacements. We therefore write
+$\theta_{\ell}=\Theta+\frac{2\pi\ell}{N}+\varphi_{\ell},$ (9)
+where $\Theta$ describes the rigid rotation of the lattice and $\varphi_{\ell}$ the internal (thermal) displacements, with $\sum_{\ell}\varphi_{\ell}=0$.
+The lattice Hamiltonian is
+$H_{\rm lat}=\sum_{\ell=1}^{N}\frac{L_{\ell}^{2}}{2mR^{2}}+\frac{\kappa R^{2}}{2}\sum_{\ell=1}^{N}(\varphi_{\ell+1}-\varphi_{\ell})^{2}.$ (10)
+In these variables the Hamiltonian separates exactly as
+$H_{\rm lat}=\frac{L^{2}}{2I_{\rm tot}}+H_{\rm ph},\qquad I_{\rm tot}=NmR^{2},$ (11)
+where
+$L=\sum_{\ell=1}^{N}L_{\ell}=-i\hbar\frac{\partial}{\partial\Theta}$ (12)
+is the angular momentum of rigid rotation of the entire ring. The term $H_{\rm ph}$ contains the N1N-1 internal normal modes built from the coordinates $\varphi_{\ell}$.
+Thus the $q=0$ mode is not discarded: it appears explicitly as the rotor term $L^{2}/2I_{\rm tot}$, while the remaining modes are the usual phonons.
+In the lattice-following description, the electron is represented in a basis of localized orbitals $\chi(\phi-\theta_{\ell})$ centered on the instantaneous atomic positions.
+A general state may be written as
+$\Psi(\phi;\{\theta_{\ell}\})=\sum_{\ell}\psi_{\ell}\,\chi(\phi-\theta_{\ell}),$ (13)
+where $\psi_{\ell}$ are the amplitudes on the orbitals.
+In this basis the electronic Hamiltonian acts on the amplitudes through nearest-neighbor matrix elements,
+$(H_{\rm el}\psi)_{\ell}=\varepsilon_{\ell}(\{\theta\})\,\psi_{\ell}-t_{\ell-1}(\{\theta\})\,\psi_{\ell-1}-t_{\ell}(\{\theta\})\,\psi_{\ell+1},$ (14)
+with periodic boundary conditions $\psi_{\ell+N}\equiv\psi_{\ell}$.
+The matrix elements depend on the instantaneous lattice configuration and therefore incorporate both the electronic structure and the electron-lattice coupling.
+For small distortions, the bond length depends only on internal differences,
+$d_{\ell}\simeq a+R(\varphi_{\ell+1}-\varphi_{\ell}),\qquad a=\frac{2\pi R}{N},$ (15)
+so that, to linear order,
+$t_{\ell}\simeq t_{0}-\alpha R(\varphi_{\ell+1}-\varphi_{\ell}).$ (16)
+Thus the electronic Hamiltonian depends only on the internal coordinates $\{\varphi_{\ell}\}$: the rigid rotation $\Theta$ drops out.
+II.3 Relative coordinate and Bloch-like structure
+Because the Hamiltonian depends on $\phi$ and $\Theta$ only through the relative coordinate $\chi=\phi-\Theta$, the eigenstates may be chosen in Bloch-like form,
+$\Psi_{k}(\phi,\Theta,\{\varphi_{\ell}\})=e^{ik(\phi-\Theta)}\sum_{\ell}u_{\ell k}(\{\varphi_{\ell}\})\,\chi(\phi-\theta_{\ell}),$ (17)
+where the coefficients $u_{\ell k}$ describe the internal structure of the state in the lattice-following basis.
+Here the instantaneous atomic positions are
+$\theta_{\ell}=\Theta+\frac{2\pi\ell}{N}+\varphi_{\ell},$ (18)
+so that the orbitals $\chi(\phi-\theta_{\ell})$ move with the lattice atoms11.
+In this form the exponential factor $e^{ik(\phi-\Theta)}$ carries the crystal pseudomomentum, while the coefficients $u_{\ell k}$ encode the internal, lattice-following structure of the electronic state.
+In the absence of distortions, $u_{\ell k}\propto e^{ik\frac{2\pi\ell}{N}}$, and Eq.
+(17) reduces to the standard tight-binding Bloch wave.
+The factor $e^{ik(\phi-\Theta)}$ in equation 17 is revealing and promising.
+The overall zero mode rotation angle $\Theta$ now takes its place among the dynamical variables.
+It is not benign, because it is not quite the angle belonging to the the total angular momentum, since that includes the electron.
+The key point is that the evolution of $\Theta$ is nontrivial, depending on recoil from changes in the angular momentum of the electron.
+It is clear from this form that if the electron $\phi$ has pseudomomentum, so must $\Theta$, and both can change, always equally and oppositely.
+This will be developed in the next section.
+The sum of the two pseudomomenta, call them $k_{\phi}=k$ and $k_{\Theta}=-k$, vanishes, and remains zero under changes of $k$, assuring the the conservation of total pseudomomentum.
+Because zero mode rotation it is now included, a world of elastic and quasielastic processess will open up.
+II.4 Total angular momentum balance
+Rotational invariance of the full Hamiltonian implies conservation of the total angular momentum operator
+$L_{\rm tot}=-i\hbar\frac{\partial}{\partial\phi}-i\hbar\frac{\partial}{\partial\Theta}.$ (19)
+Acting on the Bloch-like state gives the one-line conservation rule
+$L_{\rm tot}=L+R\hbar k,$ (20)
+where $L$ is the angular momentum conjugate to $\Theta$.
+Equation (20) implies that any change in the electronic pseudomomentum $k$ must be accompanied by a compensating change in $L$, corresponding to recoil of the ring as a whole.
+This is of course necessary to keep total angular momentum constant (lattice plus electron).
+Interestingly, there is no role yet for phonon momentum, and we can leave it this way if we want.
+As Peierls pointed out, phonon momentum can be assigned to the lattice as a whole[29], Chpts 1-2.
+Here the momentum appeared as belonging to the lattice, and indeed in must, since $L=\sum_{\ell}L_{\ell},$ which is a sum over all the motion of every atom.
+Any supposed phonon momentum would be double counting.
+We have already remarked that the form $\exp[ik(\psi-\Theta)]$ already implies the $L$ (not $L_{tot}$) must be regarded as a pseudomomentum.
+This will br defended on symmetry grounds in the next section.
+Both the electron and its host have pseudomomentum, which is actually a very comfortable circumstance.
+We should indeed be uncomfortable if we suppose that a pseudomomentum and a mechanical momentum could combine.
+II.5 Phonon-diagonal (elastic) backscattering
+The simplest illustration of the dramatic effect of keeping the zero recoil mode is provided by phonon-diagonal backscattering.
+In the absence of lattice distortions, the electronic eigenstates are plane waves
+$\left|k\right\rangle=\frac{1}{\sqrt{N}}\sum_{\ell}e^{ika\ell}c_{\ell}^{\dagger}\left|0\right\rangle,$ (21)
+with even dispersion $\varepsilon(k)=\varepsilon(-k).$ For an initial state
+$\left|\Psi_{i}\right\rangle=\left|k\right\rangle\left|\nu\right\rangle\left|L\right\rangle,$ (22)
+one allowed elastic transition is
+$\left|k\right\rangle\left|\nu\right\rangle\left|L\right\rangle\;\rightarrow\;\left|-k\right\rangle\left|\nu\right\rangle\left|L+2\hbar kR\right\rangle,$ (23)
+in which the electron reverses its direction while the internal phonon state remains unchanged. The compensating angular momentum is absorbed by rigid-body recoil of the lattice.
+The elastic backscattering channel
+$\left|k\right\rangle\left|\nu\right\rangle\to\left|-k\right\rangle\left|\nu\right\rangle$
+is kinematically forbidden in the conventional phonon-only representation unless Umklapp is invoked.
+In this process $\Delta K_{e}=-2k$ while $\Delta K_{\rm ph}=0$, so the standard selection rule cannot be satisfied by phonons alone.
+More fundamentally, the process has no carrier of the corresponding mechanical momentum once the lattice center-of-mass degree of freedom is removed.
+The absence of this channel is therefore not a dynamical result, but a consequence of having eliminated the recoil degree of freedom that would otherwise carry the compensating momentum.With the pseudoangularmomentum intact, it is kinematically robust, but to verify that this channel is dynamically allowed, we evaluate the corresponding matrix element of the electron-lattice interaction.
+In the density-density (bond-length dependent) form, the hopping amplitude depends on the relative displacements of neighboring sites, $t_{\ell}=t_{0}+\alpha(u_{\ell+1}-u_{\ell}).$ Because the interaction depends only on internal coordinate differences, it is independent of the global rotational coordinate, and therefore does not couple directly to $L$.
+Nonvanishing elastic recoil matrix element (moving basis).
+We evaluate the matrix element in an instantaneous lattice basis, in which localized electronic orbitals follow the atomic positions. Let
+$R_{\ell}(\Theta,\{\xi\})=R\Theta+\ell a+\xi_{\ell},$ (24)
+where $\Theta$ is the global rotational coordinate and $\{\xi_{\ell}\}$ are internal displacements. The electronic states are
+$\left|k;\Theta,\{\xi\}\right\rangle=\frac{1}{\sqrt{N}}\sum_{\ell}e^{ik\ell a}\,c_{\ell}^{\dagger}(\Theta,\{\xi\})\left|0\right\rangle.$ (25)
+For fixed configuration, the Hamiltonian is
+$H_{e}=-\sum_{\ell}t_{\ell}(\{\xi\})\left(c_{\ell+1}^{\dagger}c_{\ell}+\mathrm{h.c.}\right),$ (26)
+with
+$t_{\ell}(\{\xi\})=t\!\left(a+\xi_{\ell+1}-\xi_{\ell}\right).$ (27)
+The global coordinate $\Theta$ drops out, as $t_{\ell}$ depends only on internal differences.
+For the elastic recoil channel
+$\left|k,\nu,L\right\rangle\rightarrow\left|-k,\nu,L+2\hbar kR\right\rangle,$ (28)
+the electronic matrix element at fixed configuration is
+$\langle-k|H_{e}|k\rangle=-\frac{2\cos ka}{N}\sum_{\ell}t_{\ell}(\{\xi\})\,e^{2ika\ell}.$ (29)
+For a uniform lattice this vanishes, but for a generic instantaneous configuration $t_{\ell}$ is nonuniform and the $2k$ component is nonzero. To leading order,
+$t_{\ell}\simeq t_{0}+\alpha(\xi_{\ell+1}-\xi_{\ell}),$ (30)
+so that
+$\langle-k|H_{e}|k\rangle\propto\sum_{\ell}e^{2ika\ell}(\xi_{\ell+1}-\xi_{\ell}).$ (31)
+This is a satisfying outcome: the backscattering amplitude is controlled by the $2k$ Fourier component of the instantaneous deformation field and is generically nonvanishing.
+The operator acts only on internal coordinates, while the required compensating change is carried by the global mode, $L\to L+2\hbar kR$.
+The process therefore remains elastic in the phonon sector..
+The associated rotational energy shift
+$\Delta E_{\rm rot}=\frac{(L+2\hbar kR)^{2}-L^{2}}{2I_{\rm tot}}$ (32)
+is negligible for macroscopic $I_{\rm tot}=NmR^{2}$, so the process is effectively elastic.
+This example shows explicitly that momentum transfer need not be tied to a change in the internal phonon state: it may instead be carried by the lattice zero mode.
+It is worth emphasizing that the phonon-diagonal backscattering process in Eq.
+(23) is absent in standard treatments in which the lattice center-of-mass (zero mode) is eliminated.
+In that representation, momentum conservation is enforced solely through the internal phonon modes, implying $k^{\prime}=k\pm q$ and excluding the channel $\nu^{\prime}=\nu$ for $k\to-k$.
+The absence of this process therefore reflects the removal of the recoil degree of freedom rather than a dynamical constraint.
+II.6 Elastic and quasi-elastic scattering
+The general kinematics of electron-lattice scattering in the ring geometry follow directly from the conservation law Eq. (20). If the electron changes crystal momentum
+$k\rightarrow k^{\prime},$ (33)
+the lattice angular momentum must adjust according to
+$L^{\prime}=L+\hbar R(k-k^{\prime}).$ (34)
+Thus the exact momentum-conserving transitions take the form
+$\left|k,\nu,L\right\rangle\;\rightarrow\;\left|k^{\prime},\nu^{\prime},L^{\prime}\right\rangle.$ (35)
+The corresponding energy balance is
+$\varepsilon(k)-\varepsilon(k^{\prime})=\sum_{q}\hbar\omega_{q}(n_{q}^{\prime}-n_{q})+\frac{(L^{\prime})^{2}-L^{2}}{2I_{\rm tot}},$ (36)
+where the final term represents the recoil energy of the lattice. For a macroscopic ring this contribution is typically negligible, so that
+$\varepsilon(k)-\varepsilon(k^{\prime})\simeq\sum_{q}\hbar\omega_{q}(n_{q}^{\prime}-n_{q}).$ (37)
+Because the lattice angular momentum $L$ is retained as a dynamical degree of freedom, each electronic transition $k\to k^{\prime}$ is compatible with a broad family of vibrational responses:
+  1. 1.
+    Elastic recoil
+    $\left|k,\nu,L\right\rangle\rightarrow\left|k^{\prime},\nu,L+\hbar R(k-k^{\prime})\right\rangle,$ (38)
+    in which the phonon configuration is unchanged and the rigid rotation of the lattice carries the momentum transfer.
+  2. 2.
+    Quasi-elastic single-phonon processes
+    $\left|k,\nu,L\right\rangle\rightarrow\left|k^{\prime},\nu\pm 1_{q},L+\hbar R(k-k^{\prime})\right\rangle,$ (39)
+    in which the zero mode absorbs the angular momentum while a phonon adjusts the energy balance.
+  3. 3.
+    Multiphonon processes
+    $\left|k,\nu,L\right\rangle\rightarrow\left|k^{\prime},\nu+\{\Delta n_{q}\},L+\hbar R(k-k^{\prime})\right\rangle.$ (40)
+  4. 4.
+    Hybrid recoil-phonon processes in which the momentum transfer is shared between the rigid lattice rotation and the internal phonon manifold.
+Restoring the zero mode therefore enlarges the accessible final states from $\left|k^{\prime},\nu^{\prime}\right\rangle$ to $\left|k^{\prime},\nu^{\prime},L^{\prime}\right\rangle$.
+The additional angular momentum degree of freedom opens a wide manifold of momentum-conserving elastic and quasi-elastic scattering channels.
+The interaction matrix elements
+$\langle k^{\prime},\nu^{\prime},L^{\prime}|H|k,\nu,L\rangle$ (41)
+are therefore generically nonzero across this enlarged space of states.
+II.7 Consequences for one-dimensional transport and ring geometries
+In strictly one-dimensional electronic systems, two-particle collisions are strongly constrained by the simultaneous conservation of energy and momentum,
+$k_{1}+k_{2}=k_{1}^{\prime}+k_{2}^{\prime},\qquad\varepsilon(k_{1})+\varepsilon(k_{2})=\varepsilon(k_{1}^{\prime})+\varepsilon(k_{2}^{\prime}).$ (42)
+For a monotone dispersion these relations admit only trivial rearrangements, so that two-body processes cannot efficiently relax a near-Fermi distribution.
+In particular, within conventional electron-phonon treatments, even backscattering is highly constrained, since momentum transfer must be accompanied by phonon creation or annihilation, and the combined conservation laws severely restrict the available phase space.
+Standard approaches therefore invoke an additional dynamical participant. This is typically taken to be a third electronic excitation,
+$k_{1}+k_{2}+k_{3}=k_{1}^{\prime}+k_{2}^{\prime}+k_{3}^{\prime},$ (43)
+with $k_{3}$ a deep hole drawn from the Fermi sea [26, 18, 24].
+Because such excitations are thermally rare, the resulting relaxation rates are predicted to be exponentially suppressed at low temperature.
+Multiphonon processes can also relax the constraints, but only at higher order.
+A particularly clear manifestation of this reasoning appears in Coulomb drag between parallel quantum wires, where theory predicts exponential suppression of drag at low temperature, in contrast with the power-law behavior observed experimentally.
+Once the lattice center-of-mass mode is retained, however, an additional dynamical participant is always present. The conserved momentum becomes
+$P_{\rm tot}=P_{\rm CM}+\sum_{i}\hbar k_{i},$ (44)
+or in the ring geometry,
+$L_{\rm tot}=L+\sum_{i}\hbar k_{i}R.$ (45)
+An electronic momentum change $k\rightarrow k^{\prime}$ can then be balanced by recoil of the lattice,
+$\Delta P_{\rm CM}=-\hbar(k^{\prime}-k),$ (46)
+with associated energy
+$\Delta E_{\rm CM}=\frac{(\Delta P_{\rm CM})^{2}}{2M_{\rm tot}},$ (47)
+which is negligible for macroscopic $M_{\rm tot}$.
+The additional degree of freedom required by one-dimensional kinematics is therefore always available and need not be supplied by a thermally activated deep electronic excitation.
+The “third particle” is the collective zero mode of the lattice.
+As a result, large manifolds of elastic and quasi-elastic transitions exist that satisfy all conservation laws.
+Energy redistribution occurs within the phonon manifold rather than through rare electronic excitations, leading naturally to power-law, rather than exponentially suppressed, low-temperature behavior.
+Such power-law relaxation is widely observed in clean quantum wires and related one-dimensional conductors [6, 1].
+In this picture the lattice zero mode provides the natural third participant in the scattering process, enabling momentum relaxation through a broad class of elastic and quasi-elastic channels.
+The same kinematics appears in the ring geometry familiar from the theory of persistent currents.
+Once the lattice center-of-mass motion is retained, the electronic phase winding can transfer angular momentum to the lattice through recoil, with the zero mode acting as the collective degree of freedom that accommodates the momentum balance.
+Finally, we emphasize that the appearance of this additional degree of freedom is not an assumption but a consequence of enforcing the exact symmetries of the underlying Hamiltonian.
+No new interaction has been introduced, and no approximation has been invoked beyond those already implicit in the microscopic model.
+Retaining the lattice center-of-mass simply restores the full momentum conservation law of the isolated system.
+Any objection that the zero mode should not act as the required third participant must therefore confront the fact that, once this conservation law is imposed, a large manifold of elastic and quasi-elastic scattering channels follows directly.
+These channels satisfy all conservation laws and arise without fine tuning or higher-order processes.
+The resulting kinematics is thus not a hypothesis but an unavoidable consequence of the complete Hamiltonian description.
+III Foreground, Background, and Pseudomomentum
+III.1 Defining Pseudomomentum
+It is important that we clearly distinguish component parts We distinguish foreground and background degrees of freedom.
+The foreground is the object of interest, here a single electron, or it could be a collection of them, or a defect, or a lattice phonoon.
 The background consists of all remaining degrees of freedom, including the lattice atoms and any other electrons not treated explicitly.
@@
 The associated conserved quantity is pseudomomentum (crystal momentum); for an electron in a periodic potential this is the familiar Bloch momentum $\hbar\mathbf{k}$.
-Crucially, the same discrete translational symmetry applies to the background (see Fig.
-1).
+The same discrete translational symmetry applies to the background.
+Translating the entire lattice while holding the electron fixed leaves the Hamiltonian invariant, and therefore assigns a pseudomomentum to the background as well.
+(see Fig.
+2).
 If the electron (foreground A) coordinate is held fixed and the entire background B is translated by a lattice vector, the Hamiltonian is again unchanged.
@@
 Figure 2: Sliding Blocks.
-Floating in space, two blocks attract one another and slide frictionlessly along their common long axis.
+Untethered, two blocks attract one another and slide frictionlessly along their common long axis.
 At the top, both blocks carry conserved mechanical momenta, and clearly $P_{\rm tot}=p_{A}+p_{B}$.
-At the bottom, the same relation holds, but $p_{A}$ and $p_{B}$ are now pseudomomenta; their sum nevertheless equals the conserved mechanical momentum.
+At the bottom, the same relation holds, but $p_{A}$ and $p_{B}$ are now not independently conserved; if the boxes are sliding over one another,their sum nevertheless equals the conserved mechanical momentum.
 The two blocks serve as surrogates for a foreground electron (A) in a perfect background crystal (B), illustrating that both the foreground and the background acquire pseudomomentum in each other’s presence.
@@
 […]
-so the corresponding Noether generator is the conserved total mechanical momentum, denoted by $\hat{\mathbf{Q}}$.
+so the corresponding symmetry is the conserved total mechanical momentum, denoted by $\hat{\mathbf{Q}}$.
 In the foreground/background formulation $\hat{\mathbf{Q}}$ is obtained as the sum of two subsystem pseudomomenta.
@@
 Here $\hat{\mathbf{p}}$ is the pseudomomentum carried by the foreground electron (the generator of its translation), and $\hat{\mathbf{P}}$ is the pseudomomentum carried by the background.
+The total background pseudomomentum decomposes into
+$\hat{\mathbf{P}}=\hat{\mathbf{P}}_{\bf R}+\hat{\mathbf{P}}_{\rm ph}$ (51)
+the center-of-mass (zero-mode) contribution and the internal phonon contribution.
 (the generator of translating the background as a whole).
@@
 For convenience, we can set it to 0; it never changes and we can now ignore it.
-Even under the usual rule stating that defects, such as a crystal edge or impurity, scatter electrons elastically, there is a subtlety: the electron changes its pseudomomentum, but the lattice cannot the respond or recoil in kind.
+Even under the usual rule stating that defects, such as a crystal edge or impurity, scatter electrons elastically, there is a subtlety within the prior framework: the electron changes its pseudomomentum, but as is conventionally understood, the lattice cannot respond or recoil in kind.
 A pseudomomentum cannot be combined with a mechanical momentum.
-We have now seen, however, that both foreground and background have pseudomomentum, and can exchage it freely.
+We have now seen, however, that both foreground and background have pseudomomentum, and can exchange it freely.
 Total momentum is conserved for example in the very real deflection of an electron at an edge by equal and opposite pseudomomentum changes.
-These rules are analogous to those governing the collision of two objects in free space, except that here the interaction is periodic in their relative coordinate, apart from defects and lattice vibrations.
-The total momentum $\hat{\mathbf{Q}}$ is conserved, but the foreground and background momenta, $\hat{\mathbf{p}}$ and $\hat{\mathbf{P}}$, undergo equal and opposite changes, in exact analogy with collisions in free space.
-This exchange does not require the creation or annihilation of internal lattice excitations and occurs immediately at the collision event (see Fig.
-2).
+Historically, there has been a trend to treat electron scattering in a crystal as if it were Compton scattering in free space.
+There is no ether, no center of mass of a substrate.
+This is elegant, but also it misses something important.
+Here, we restore for a periodic lattice a full set of rules that apply in free space, as one would wish.
+There is a total psuedomomentum conserved even if an electron encounters a defect.
+However, unlike in free space, in crystals there is an aether, so to speak, a silent partner: the lattice.
+It is more than its phonons, it has a center of mass.
+Figure 3 reflects the revised understanding and diagram including the participation of the lattice. This collision takes place in an “ether.”
+Figure 3: (top) Traditional diagram of electron deflection by phonon emission or absorption.
+(bottom) Momentum transfer in a perfect crystal when the lattice zero mode is retained.
+The collision occurs on a dynamical lattice: quasielastic phonons may be produced, while the change in electron momentum is balanced jointly by the phonon momentum and the center-of-mass recoil of the lattice.
+Note: When momentum $\Delta p=\hbar q$ is transferred to the lattice center of mass, the associated de Broglie wavelength is $\lambda=2\pi/q$, identical to that of the scattering process.
+The large lattice mass suppresses the recoil energy but not the wavelength or the momentum transfer.
+The total momentum $\hat{\mathbf{Q}}$ is conserved, but the foreground and background momenta, $\hat{\mathbf{p}}$ and $\hat{\mathbf{P}}=\hat{\mathbf{P}}_{R}+\hat{\mathbf{P}}_{\rm int}$, undergo equal and opposite changes, in exact analogy with collisions in free space.
+In particular, any change in $\hat{\mathbf{p}}$ necessarily induces a compensating change in the global (zero-mode) contribution $\hat{\mathbf{P}}_{R}$, while changes in the internal degrees of freedom are optional This exchange does not require the creation or annihilation of internal lattice excitations and occurs immediately at the collision event (see Fig.
+4).
 The very real recoil of the atom labeled in red in Fig.
@@
 This is just the elementary fact that a slightly displaced ground state of a harmonic oscillator is still mostly in the ground state, if measured.
-Two measurable subsystem momenta and the total momentum
-Figure 1 shows two variations on a mechanical system. At the top, two smooth blocks interact frictionlessly. At the bottom, both blocks are corrugated, leading to periodic interaction.
+In what follows, we formalize a symmetric foreground-background structure, derive the elastic scattering amplitude within first-order perturbation theory, and compute the elastic fraction explicitly.
+We show that this fraction is close to unity for copper at room temperature.
+We also clarify how the present formulation extends the Fröhlich framework without contradicting its successful predictions, and discuss the implications for the interpretation of transport phenomena in crystalline solids.
 Figure 4: The total momentum (black) remains constant as the foreground (electron, blue) and background (lattice, red) momenta jump abruptly at the collision time $t^{*}$.
@@
 The total appears in the wavefunctions.
-Finite-window symmetry
+As with Mössbauer scattering, the red atom has gained momentum and with it the whole lattice, but when measured, the lattice may still be in the same vibrational state as before the collision (elastic scattering).
+III.2 Finite-window symmetry
 In any physically realizable crystal the interaction is periodic only over a finite spatial extent $L$.
@@
 Any “phonon carrying momentum $\hbar q$” is then not an eigenstate; it’s a wavepacket (a superposition of standing modes) that only behaves like a traveling object transiently until it reflects, mode-mixes, etc.
-III Lattice degrees of freedom and zero modes
-We consider a finite crystal consisting of atoms with coordinates $\hat{\mathbf{R}}_{j}$ and masses $m_{j}$, with total lattice mass $M=\sum_{j}m_{j}$, minus the foreground electron.
+IV Lattice degrees of freedom and zero modes
+We have already encountered a zero mode in the minimal model of an electron and an atom chain. section II. The mode there is the angular momentum.
+Now we venture out more generally, and find the same large phase space becomes available for elastic and quasielastic scattering processes, a phase space that is simply absent if the center of mass is absent in the Hamiltonian.
+We consider a finite crystal consisting of atoms with coordinates $\hat{\mathbf{R}}_{j}$ and masses $m_{j}$, with a total lattice mass $M=\sum_{j}m_{j}$, minus the foreground electron.
 We define the center-of-mass coordinate of the lattice background as
@@
 All mechanical momentum of the crystal resides in the translational zero mode.
-IV Electron-Lattice interaction
-IV.1 Foreground and background Hamiltonians
+V Electron-Lattice interaction
+V.1 Foreground and background Hamiltonians
 . We write the total Hamiltonian as
@@
 The use of a bare kinetic-energy operator reflects the fact that the electron has not been pre-diagonalized into Bloch states.
-All effects of the periodic crystal potential, including band structure, pseudomomentum exchange, and scattering, are instead generated dynamically through the interaction Hamiltonian $H_{\rm int}$.
+All effects of the periodic crystal potential, including band structure, pseudomomentum exchange, and scattering, are generated through the interaction Hamiltonian $H_{\rm int}$.
 This is the same starting point adopted by Fröhlich [10].
@@
 Their degrees of freedom are therefore implicit in $H_{\rm bg}$.
-Density operator and exact factorization
+V.2 Density operator and exact factorization
 The lattice density operator admits an exact factorization into a center-of-mass translation and an internal part:
@@
 This identity expresses a purely kinematic fact: a rigid translation of the lattice multiplies every density Fourier component by the phase $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$, while all internal lattice dynamics are contained entirely in $\hat{\rho}^{\rm int}_{\mathbf{q}}$.
-Density–density interaction and pseudomomentum conservation
+V.3 Density–density interaction and pseudomomentum conservation
 The electron–lattice interaction may be written in density–density form as
-$H_{\rm int}=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}\,\hat{\rho}_{-\mathbf{q}},$ (12)
-following the formalism of Van Hove [30]. No assumption about plane-wave eigenstates is required; $\mathbf{q}$ labels Fourier components of the interaction kernel.
-Using Eq. (11), this becomes
-$H_{\rm int}$ $=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i{\mathbf{q}}\cdot\hat{\bm{\rho}}}\,\,\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (13)
-$=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}\,\hat{\rho}^{\rm int}_{-\mathbf{q}}.$
-[…]
-Crucially, this exchange of pseudomomentum applies whether or not there is a change a change in phonon occupation numbers.
-Elastic matrix elements of $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ are generically nonzero.
-The background zero mode is therefore capable of absorbing finite pseudomomentum with negligible energy cost, providing the kinematic basis for elastic but momentum-relaxing scattering.
-Any combination of phonon involvement and lattice background recoil is allowed.
-It seems this should matter for measurable properties.
-V Elastic internal scattering and the Debye-Waller fraction
-In this section we show that, even within a strictly perturbative treatment based on Bloch electrons and phonon number states, robust elastic scattering channels exist already at first order in the electron–lattice coupling.
-These channels become explicit once the full lattice density operator is retained and the translational zero mode of the lattice is treated as a dynamical degree of freedom.
+$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,\rho_{\rm fg}(\mathbf{q})\,\rho_{\rm bg}(-\mathbf{q}),$ (57)
+following the formalism of Van Hove [34].
+No assumption about Bloch-wave eigenstates is required; $\mathbf{q}$ labels Fourier components of the interaction kernel.
+This equation is the essence of the term: density-density interaction.
+For a tagged foreground electron, we write the aperiodic part of the Bloch waves as
+$\rho_{\rm fg}(\mathbf{q})=e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}$. (58)
+Substituting into Eq. (57) gives
+$\boxed{H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}\,\hat{\rho}^{\rm int}_{-\mathbf{q}}}$ (59)
+Equation (59) shows that the interaction depends only on the relative coordinate $\hat{\bm{\rho}}=\hat{\mathbf{r}}-\hat{\mathbf{R}}$.
+The symmetry of the phase factor $e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}$ is the central structural feature: pseudomomentum $\mathbf{q}$ is exchanged between foreground and background with change‘s always equal and opposite.
+When the electron momentum changes from $k$ to $k^{\prime}$, the lattice background absorbs the compensating momentum
+$\Delta P_{\rm bg}=-\hbar(k^{\prime}-k).$ (60)
+This momentum is carried entirely by the center-of-mass (zero mode) of the background. Internal lattice excitations describe only relative motion and carry no net mechanical momentum.
+The associated energy change separates into internal and recoil contributions,
+$\Delta E_{\rm lat}=\Delta E_{\rm int}+\frac{\Delta P_{R}^{\,2}}{2M_{\rm tot}},$ (61)
+where $\Delta P_{R}=-\hbar(k^{\prime}-k)$ is the center-of-mass recoil momentum.
+For macroscopic $M_{\rm tot}$, the recoil energy is negligible, so large momentum transfer can occur with little or no change in internal lattice energy.
+Translational invariance implies that the background degrees of freedom contain a rigid (center-of-mass) coordinate $\mathbf{R}$ with conjugate momentum
+$\hat{\mathbf{P}}_{\bf R}=-\,i\hbar\frac{\partial}{\partial\mathbf{R}}.$
+The remaining background variables describe internal excitations (phonons or distortions) defined relative to the center of mass. Accordingly, the background Hilbert space factorizes as
+$\mathcal{H}_{\rm bg}=\mathcal{H}_{R}\otimes\mathcal{H}_{\rm int}.$
+All mechanical momentum of the background resides in the zero mode,
+$\hat{\mathbf{P}}=\hat{\mathbf{P}}_{\bf R},$ (62)
+while the internal degrees of freedom carry no net mechanical momentum.
+The interaction operator
+$e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}$
+transfers momentum $\hbar\mathbf{q}$ from the electron to the background as a whole.
+This momentum is carried entirely by the center-of-mass degree of freedom.
+Lattice recoil is therefore not an additional dynamical channel but a direct consequence of translational symmetry.
+A nonvanishing matrix element of $H_{\rm int}$ requires
+$\Delta\mathbf{p}=+\hbar\mathbf{q},\qquad\Delta\mathbf{P}=-\hbar\mathbf{q},$ (63)
+so that the total mechanical momentum $\hat{\mathbf{Q}}=\hat{\mathbf{p}}+\hat{\mathbf{P}}$ is conserved exactly.
+Crucially, this momentum exchange is independent of phonon occupation.
+Elastic matrix elements of $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ are generically nonzero, so the lattice can absorb finite momentum through center-of-mass recoil without any change in its internal state.
+VI Elastic internal scattering and the Debye-Waller fraction
+We now show that, even within a strictly perturbative treatment based on Bloch electrons and phonon number states, robust elastic scattering channels exist already at first order in the electron–lattice coupling.
+These channels become explicit once the full lattice density operator is retained and the translational zero mode is treated as a dynamical degree of freedom.
 The resulting processes are elastic with respect to internal lattice excitations, despite involving momentum exchange with the lattice as a whole.
-Density–density coupling and emergence of the Bloch potential
-We take the foreground–background interaction in density–density form,
-$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,\rho_{\rm fg}(\mathbf{q})\,\rho_{\rm bg}(-\mathbf{q}),$ (14)
-with the tagged-electron density operator
-$\rho_{\rm fg}(\mathbf{q})=e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}$. (15)
-For the background we use the exact center-of-mass factorization
-$\rho_{\rm bg}(-\mathbf{q})=\sum_{j}e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}_{j}}=e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (16)
-where $\hat{\mathbf{R}}_{j}=\hat{\mathbf{R}}+\bm{\xi}_{j}^{(0)}+\hat{\mathbf{u}}_{j}$.
-Separating the background density operator into its equilibrium expectation value (taken in the background center-of-mass frame) and fluctuations,
+VI.1 Emergence of the Bloch potential
+Separating the background density operator into its equilibrium expectation value (taken in the lattice center-of-mass frame) and fluctuations,
 $\rho_{\rm bg}(-\mathbf{q})=\langle\rho_{\rm bg}(-\mathbf{q})\rangle+\delta\rho_{\rm bg}(-\mathbf{q}),$ (64)
-the static component produces an effective periodic potential acting on the foreground electron,
+the static component produces an effective periodic potential
 $U_{\rm Bloch}(\hat{\mathbf{r}})=\sum_{\mathbf{G}}V(\mathbf{G})\,e^{i\mathbf{G}\cdot\hat{\mathbf{r}}}\,\langle\rho_{\rm bg}(-\mathbf{G})\rangle,$ (65)
 where $\mathbf{G}$ are reciprocal lattice vectors.
-The corresponding single-particle foreground Hamiltonian becomes
+The corresponding foreground Hamiltonian becomes
 […]
-which is the usual Bloch Hamiltonian. This periodic potential is not assumed a priori, but emerges as the static mean-field component of the underlying density–density coupling.
-The remaining part of $H_{\rm int}$ couples the foreground electron to background density fluctuations $\delta\rho_{\rm bg}$, but with the crucial difference from the clamped-lattice formulation: the full operator $\rho_{\rm bg}(-\mathbf{q})$ retains the explicit center-of-mass translation factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$, which enforces exact momentum bookkeeping in the finite, untethered crystal.
-Initial and final states
-For a perturbative treatment we consider product states of foreground and background,
-$\Psi_{i}=\Psi_{n\mathbf{k}}(\mathbf{r})\,\Psi^{\rm bg}_{i}(\{\mathbf{u}_{j}\}),\qquad\Psi_{f}=\Psi_{n^{\prime}\mathbf{k}^{\prime}}(\mathbf{r})\,\Psi^{\rm bg}_{f}(\{\mathbf{u}_{j}\}),$ (20)
-where $\Psi^{\rm bg}_{i,f}$ are phonon number states or superpositions thereof. The electronic states are Bloch eigenstates of the static lattice potential.
-Using the factorized form of the interaction, any nonvanishing matrix element $\langle\Psi_{f}|H_{\rm int}|\Psi_{i}\rangle$ implies the pseudomomentum transfers
-$\Delta\mathbf{p}=+\hbar\mathbf{q},\qquad\Delta\mathbf{P}=-\hbar\mathbf{q},$ (21)
-so that the total mechanical momentum of the combined system is conserved. The recoil pseudomomentum is carried by the lattice center-of-mass degree of freedom, regardless of phonon number changes.
-Elastic internal channel and the Debye-Waller factor
-For a harmonic lattice, the internal displacements admit the normal-mode expansion
-$\hat{\mathbf{u}}_{j}=\frac{1}{\sqrt{N}}\sum_{\mathbf{k},\lambda}\sqrt{\frac{\hbar}{2M\omega_{\mathbf{k}\lambda}}}\,\mathbf{e}_{\mathbf{k}\lambda}\left(a_{\mathbf{k}\lambda}e^{i\mathbf{k}\cdot\bm{\xi}^{(0)}_{j}}+a^{\dagger}_{\mathbf{k}\lambda}e^{-i\mathbf{k}\cdot\bm{\xi}^{(0)}_{j}}\right)$.
-(22)
-The operator $e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}$ therefore contains all orders in the phonon operators and automatically generates elastic, quasielastic, and inelastic processes.
-In particular, in a phonon number state basis the diagonal matrix elements are generically nonzero,
-$\left\langle\{n\}\left|e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\right|\{n\}\right\rangle\neq 0,$ (23)
-so that a scattering channel strictly elastic with respect to internal lattice excitations is present already at first order in $H_{\rm int}$.
-Upon thermal averaging (or in the ground state), this diagonal factor reduces to the Debye-Waller form
-$\left\langle e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\right\rangle_{T}=\exp[-W(\mathbf{q},T)],$ (24)
-which is the direct analog of the recoil-free (zero-phonon-line) fraction in Mössbauer spectroscopy and in neutron or x-ray scattering.
-While much stronger coupling regimes can be treated nonperturbatively—for example within wave-on-wave (WoW) simulations [4] the present analysis shows that even within standard perturbation theory there exist elastic scattering processes that are absent when the center of mass is hidden or eliminated.
-Relation to the Fröhlich truncation
-In the original Fröhlich formulation, the exact factorization Eq.
-(11) is not employed.
-The lattice center-of-mass coordinate $\hat{\mathbf{R}}$ is not retained explicitly, which is equivalent to setting the recoil factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$ to unity and treating the lattice as a clamped reference frame.
-At the same time, the internal displacement exponential is truncated to linear order,
-$e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\rightarrow 1-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}.$ (25)
-The zeroth-order term reproduces the static periodic potential responsible for Bloch band formation, while the linear term generates single-phonon processes.
-Although this procedure is internally consistent and highly successful in the weak-coupling regime, it removes the background recoil zero mode by construction.
-As a result, even if the full exponential in $\hat{\mathbf{u}}_{j}$ were retained, the truncated theory would still lack an explicit degree of freedom capable of absorbing the total recoil momentum of the lattice.
-The standard Fröhlich Hamiltonian therefore describes momentum exchange with internal lattice modes, but the 0 mode, the lattice recoil as a whole, is absent.
-Thermal averaging and the Debye-Waller fraction
+which is the usual Bloch Hamiltonian.
+In this way, the periodic potential is not imposed a priori, but emerges as the mean-field component of the translationally invariant density-density interaction in the background center-of-mass frame.
+The remaining part of $H_{\rm int}$ describes coupling to fluctuations of the background density.
+Crucially, it retains the explicit center-of-mass translation factor, ensuring that momentum transfer is accounted for by recoil of the background as a whole, rather than being assigned solely to internal phonon modes.
+VI.2 Initial and final states
+For a perturbative treatment, we consider product states
+$\Psi_{i}=\Psi_{n\mathbf{k}}(\mathbf{r})\,\Psi^{\rm bg}_{i},\qquad\Psi_{f}=\Psi_{n^{\prime}\mathbf{k}^{\prime}}(\mathbf{r})\,\Psi^{\rm bg}_{f},$ (67)
+where $\Psi_{n\mathbf{k}}$ are extended-zone Bloch eigenstates of the static lattice Hamiltonian and $\Psi^{\rm bg}_{i,f}$ are arbitrary background states in the full lattice Hilbert space (not restricted to phonon number eigenstates).
+VI.3 Interaction and matrix element
+From Eq. (59), the interaction is
+$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (68)
+where $\hat{\mathbf{R}}$ is the lattice center-of-mass operator and $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ acts on internal background degrees of freedom.
+Between product states the interaction matrix element is
+$\mathcal{M}_{fi}$ $=\big\langle n^{\prime}\mathbf{k}^{\prime},\Psi_{f}^{\rm bg}\big|H_{\rm int}\big|n\mathbf{k},\Psi_{i}^{\rm bg}\big\rangle$
+$=V(\mathbf{q})\,\langle n^{\prime}\mathbf{k}^{\prime}|e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}|n\mathbf{k}\rangle\,\langle\Psi_{f}^{\rm bg}|e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\hat{\rho}^{\rm int}_{-\mathbf{q}}|\Psi_{i}^{\rm bg}\rangle.$ (69)
+VI.4 Total pseudomomentum conservation
+Translational invariance of the full electron-lattice system implies exact conservation of total pseudomomentum,
+$\mathbf{k}^{\prime}-\mathbf{k}=-\big(\mathbf{K}^{\rm bg}_{f}-\mathbf{K}^{\rm bg}_{i}\big),$ (70)
+where $\mathbf{K}^{\rm bg}_{i,f}$ are eigenvalues of the total background pseudomomentum operator, including the zero (center-of-mass) mode.
+Equation (70) is the only kinematic constraint imposed by translational symmetry. No additional reciprocal lattice vector is required when electronic states are treated in the extended-zone scheme.
+In mechanical form this reads
+$\Delta\mathbf{p}=-\Delta\mathbf{P},$ (71)
+so that the total mechanical momentum $\hat{\mathbf{Q}}=\hat{\mathbf{p}}+\hat{\mathbf{P}}$ is conserved exactly.
+VII Partition of pseudomomentum transfer
+We now examine how the transferred pseudomomentum $\hbar(\mathbf{k}^{\prime}-\mathbf{k})$ is stored within the background.
+The background matrix element contains two distinct structures,
+$e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\qquad\text{(center-of-mass recoil)},$ (72)
+and
+$\hat{\rho}^{\rm int}_{-\mathbf{q}}\qquad\text{(internal lattice operator)}.$ (73)
+The conservation law (70) constrains only the total background pseudomomentum change $\Delta\mathbf{K}^{\rm bg}.$ It places no requirement on how that momentum is distributed between internal modes and the center-of-mass degree of freedom.
+VII.1 Conventional phonon bookkeeping
+In the clamped-lattice approximation one neglects the recoil operator and expands background states in phonon number eigenstates. In that restricted description one identifies
+$\Delta\mathbf{K}^{\rm bg}=\sum_{\mathbf{q}}\Delta n_{\mathbf{q}}\,\mathbf{q},$ (74)
+so that electronic momentum change is compensated entirely by phonon creation or annihilation.
+Energy conservation then requires
+$\epsilon_{n^{\prime}\mathbf{k}^{\prime}}=\epsilon_{n\mathbf{k}}\pm\hbar\omega_{\mathbf{q}\lambda}.$ (75)
+This phonon-counting condition is therefore a consequence of the harmonic number-state basis together with suppression of the center-of-mass degree of freedom.
+It restricts the dynamics to a low-codimension subset of the full many-body phase space, in which momentum and energy must be matched by discrete phonon quanta.
+VII.2 Full Hilbert-space description
+In the complete electron-lattice Hilbert space, Eq.
+(70) remains the only exact constraint.
+The compensating background momentum is carried entirely by the center-of-mass (zero mode) of the lattice.
+Internal degrees of freedom describe only relative motion and carry no net mechanical momentum.
+Thus, the association of background momentum transfer with specific phonon occupation changes is not fundamental, but a consequence of discarding the zero mode.
+Phonons may accompany a scattering event, but they do not provide the required momentum balance.
+The zero mode is the unique carrier of mechanical momentum.
+In the clamped-lattice formulation, an electronic transition $k\to k^{\prime}$ is restricted to those phonon modes whose wavevector and frequency simultaneously satisfy both momentum and energy matching conditions.
+In low dimensions or at low temperature, such solutions are sparse and may be absent altogether except through Umklapp.
+When it is said that Umklapp momentum is “absorbed by the lattice,” the role of the zero mode is implicitly reintroduced for the special case of a reciprocal lattice vector $G$, even though the lattice can in fact absorb arbitrary momentum.
+By contrast, when the translational zero mode is retained, momentum conservation is enforced by recoil of the background as a whole, while energy conservation is governed independently by the internal spectrum.
+The internal lattice states form an exponentially dense many-body manifold, and all configurations within an energy window $|E_{f}-E_{i}|\lesssim\Gamma$ contribute.
+The number of quasi-elastic channels therefore scales with the many-body density of states and becomes macroscopic in the thermodynamic limit.
+This vast phase space is absent when the zero mode is omitted.
+VII.3 Recoil-dominated channel
+When the zero recoil mode is retained, the required pseudomomentum transfer is taken up entirely by the background center-of-mass degree of freedom:
+$\langle\mathbf{K}^{\prime}|e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}|\mathbf{K}\rangle=\delta_{\mathbf{K}^{\prime},\,\mathbf{K}-\mathbf{q}},$ (76)
+so that total pseudomomentum conservation $\mathbf{k}^{\prime}-\mathbf{k}=-(\mathbf{K}^{\rm bg}_{f}-\mathbf{K}^{\rm bg}_{i})$ is satisfied without reference to phonon occupation changes.
+The associated recoil energy is
+$\Delta E_{\rm recoil}=\frac{\hbar^{2}}{2M_{\rm tot}}\left[(\mathbf{K}-\mathbf{q})^{2}-\mathbf{K}^{2}\right]=\frac{(\hbar q)^{2}}{2M_{\rm tot}}-\frac{\hbar^{2}}{M_{\rm tot}}\mathbf{K}\!\cdot\!\mathbf{q},$ (77)
+which scales as $1/M_{\rm tot}$ and therefore vanishes in the thermodynamic limit. To leading order in $1/N$, the scattering is thus elastic with respect to internal lattice excitations.
+The internal factor
+$\langle\Psi_{f}^{\rm bg}|\hat{\rho}^{\rm int}_{-\mathbf{q}}|\Psi_{i}^{\rm bg}\rangle$ (78)
+remains finite even when no phonon quanta are created or annihilated.
+For a thermally fluctuating background its diagonal contribution reduces, upon averaging, to the Debye-Waller factor $e^{-W(\mathbf{q},T)}.$
+VII.4 Quasielastic and superelastic processes
+More generally, $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ contains all orders in internal lattice operators.
+Transitions therefore exist in which the center-of-mass absorbs the entire pseudomomentum transfer, while the internal modes undergo small changes that adjust the energy balance.
+Such processes are quasielastic:
+$\epsilon_{n^{\prime}\mathbf{k}^{\prime}}\approx\epsilon_{n\mathbf{k}},\qquad|\Delta E_{\rm int}|\ll\epsilon_{n\mathbf{k}},$ (79)
+with the exact kinematic constraint always enforced by center-of-mass recoil.
+At finite temperature, superelastic channels occur when pre-existing internal excitations are partially annihilated, transferring energy to the electron while the net pseudomomentum remains balanced by the center-of-mass degree of freedom.
+VII.5 Elastic and quasi-elastic processes beyond the Fröhlich term
+Consider the electron-lattice interaction in its full density-density form
+$H_{\rm int}=\sum_{q}V(q)\,\rho_{e}(-q)\,\rho_{\rm bg}(q),\qquad\rho_{e}(-q)=\sum_{k}c^{\dagger}_{k+q}c_{k}.$ (80)
+The background density may be written as
+$\rho_{\rm bg}(q)$ $=\sum_{j}e^{-iq(R_{j}+u_{j})}$ (81)
+$=\sum_{j}e^{-iqR_{j}}\left(1-iq\,u_{j}-\frac{q^{2}}{2}u_{j}^{2}+\cdots\right).$
+where $R_{j}$ are equilibrium positions and $u_{j}$ displacements.
+The zeroth-order term gives
+$H^{(0)}=\sum_{k,q}V(q)\,c^{\dagger}_{k+q}c_{k}\sum_{j}e^{-iqR_{j}},$ (82)
+which contains no phonon operators and is therefore purely elastic.
+In a perfect periodic crystal this reduces to reciprocal lattice scattering ($q=G$), i.e.
+Bragg processes.
+It is responsible for the static band structure.
+As Kohn noted[20], there is plenty of elastic deflection going into the construction of bands, which gets masked in making Bloch waves.
+Figure 5 makes this clear.
+Figure 5: Interrupted Bloch wave Both panels feature a circular absorbing zone seen clearly, as the real part of a plain wave (left) and of a Bloch wave (right) enter from above.
+A shadow persists and slowly fills by diffraction on the left, plane wave case, but strong scattering by the triangular lattice present on the right quickly fills in the void.
+This scattering, which is obscured by periodic coherence in the unobstructed Bloch waves, is plainly revealed as the void fills.
+The point is that the strong unprotected scattering in the shadow region cannot be strongly inelastic, or else passive barriers in periodic lattices would cause heating downstream.
+The linear term
+$H^{(1)}=-i\sum_{k,q}q\,V(q)\,c^{\dagger}_{k+q}c_{k}\sum_{j}e^{-iqR_{j}}u_{j},$ (83)
+yields, upon expanding $u_{j}$ in phonon operators, the familiar Fröhlich interaction in which each term carries a definite phonon pseudomomentum.
+At quadratic order one finds
+$H^{(2)}=-\frac{1}{2}\sum_{k,q}q^{2}V(q)\,c^{\dagger}_{k+q}c_{k}\sum_{j}e^{-iqR_{j}}u_{j}^{2},$ (84)
+and writing
+$u_{j}=\frac{1}{\sqrt{N}}\sum_{Q}U_{Q}\,e^{iQR_{j}}(a_{Q}+a^{\dagger}_{-Q}),$ (85)
+one obtains terms of the form
+$\begin{split}H^{(2)}\sim\sum_{k,q}\sum_{Q,Q^{\prime}}c^{\dagger}_{k+q}c_{k}\,(a_{Q}+a^{\dagger}_{-Q})(a_{Q^{\prime}}+a^{\dagger}_{-Q^{\prime}})\\ \times\sum_{j}e^{-i(q-Q-Q^{\prime})R_{j}}.\end{split}$ (86)
+Among these are phonon-diagonal contributions with $Q^{\prime}=-Q$,
+$H^{(2)}_{\rm el}\sim\sum_{k,q,Q}c^{\dagger}_{k+q}c_{k}\,\bigl(2a_{Q}^{\dagger}a_{Q}+1\bigr)\sum_{j}e^{-iqR_{j}},$ (87)
+which have nonzero matrix elements
+$\langle\nu|H^{(2)}_{\rm el}|\nu\rangle\neq 0.$ (88)
+These describe elastic electron scattering with $\nu\to\nu$, and more generally quasi-elastic processes when $\nu\to\nu^{\prime}$ with small energy change.
+Thus, once the interaction is treated beyond the linear Fröhlich approximation, elastic and quasi-elastic channels appear already at first order in the density-density interaction.
+In such processes $\Delta K_{\rm ph}=0$, while in general $\Delta K_{e}\neq Gm$, so the standard selection rule
+$\Delta K_{e}+\Delta K_{\rm ph}=Gm$ (89)
+is not satisfied within the zero-mode-removed theory.
+This is consistent with the observation, emphasized by Kohn, that the electronic band structure itself arises from coherent elastic scattering from the periodic lattice potential.
+Elastic scattering is therefore not an exceptional process but the dominant one that defines the electronic states.
+The restriction to inelastic, one-phonon processes is a feature of the linearized phonon representation, not of the underlying dynamics.
+The appearance of elastic and quasi-elastic channels thus exposes the limitations of the reduced description.
+Once the lattice center-of-mass degree of freedom is omitted, neither mechanical momentum nor pseudomomentum is properly accounted for within the phonon subspace alone.
+The full problem requires a dynamical background capable of carrying the compensating momentum.
+VII.6 Thermal averaging and the Debye-Waller fraction
 In thermal equilibrium the internal lattice degrees of freedom are described by the density matrix
@@
 $f_{\mathrm{el}}(\mathbf{q},T)\simeq e^{-2W(\mathbf{q},T)},\qquad W(\mathbf{q},T)=\frac{1}{2}\left\langle(\mathbf{q}\cdot\hat{\mathbf{u}})^{2}\right\rangle_{T},$ (95)
-the familiar Debye–Waller result. Even at finite temperature a substantial fraction of momentum-transfer events leave the lattice in the same internal state.
-Estimate of the elastic internal fraction at representative momentum transfers
+the familiar Debye-Waller result. Even at finite temperature a substantial fraction of momentum-transfer events leave the lattice in the same internal state.
+VII.7 Estimate of the elastic internal fraction at representative momentum transfers
 For momentum relaxation in a metal it is natural to consider momentum transfers of order the Fermi momentum, since large-angle scattering dominates the transport relaxation rate through the usual $(1-\cos\theta)$ weighting.
@@
 Momentum exchange with the lattice therefore occurs predominantly through elastic or quasi-elastic coupling to the background degrees of freedom, with internal phonon excitation representing a secondary correction.
-VI Experiments suggesting large elastic contributions
-Experimental evidence for predominantly elastic momentum relaxation in metals at low temperature has accumulated for decades, particularly in the mesoscopic-transport and quantum-interference literature.
-In this section we show that the framework developed here resolves several longstanding disparities between theory and experiment regarding electronic transport.
+We can motivate this result as follows: For a Cu atom modeled as an Einstein oscillator, a sudden momentum transfer $\Delta p=\hbar q$ leaves the oscillator in its ground state with probability
+$P_{0\to 0}=\left|\langle 0|e^{iqx}|0\rangle\right|^{2}=\exp\!\left(-\frac{E_{R}}{\hbar\omega_{E}}\right),\qquad E_{R}=\frac{(\hbar q)^{2}}{2M}.$ (98)
+Taking $q=2k_{F}$ for copper, with $k_{F}\approx 1.36\,\mathrm{\AA }^{-1}$ and $M\approx 63.5\,u$, gives $E_{R}\approx 0.24\,\mathrm{meV}$.
+For a typical local vibrational energy $\hbar\omega_{E}\sim 20$-$30\,\mathrm{meV}$, one finds $P_{0\to 0}\approx 0.99$.
+VIII Experiments suggesting large elastic contributions
+We have already touched on pure wire resistivity and persistent currents above, as allowing a more satisfactory explanation of the experiments through inclusion of the zero modes and the expansion to elastic and quasielastic channels that they enable.
+Other experimental evidence suggesting predominantly elastic momentum relaxation in metals at low temperature has accumulated for decades, particularly in the mesoscopic-transport and quantum-interference literature.
+In this section we show that the framework developed here can resolve several longstanding disparities between theory and experiment regarding electronic transport.
 We review a range of experiments that demonstrate a clear separation between momentum relaxation and energy relaxation, consistent with elastic or quasi-elastic scattering dominating momentum randomization.
-Hierarchy of momentum and energy relaxation times
+VIII.1 Hierarchy of momentum and energy relaxation times
 Electrical resistivity and optical conductivity probe momentum relaxation, quantum-interference phenomena such as weak localization probe phase coherence, and nonequilibrium measurements including shot noise, Johnson-noise thermometry, and hot-electron relaxation directly access energy exchange.
@@
 The consistent appearance of this hierarchy demonstrates that momentum relaxation and energy relaxation are distinct physical processes.
-While conventional language often associates momentum randomization with inelastic scattering, these experiments show that electronic momentum can be efficiently scrambled by a time-dependent environment without significant energy exchange.
-The lattice need not act as an energy sink in order to act as a momentum sink.
-This separation is precisely what is expected when electrons scatter elastically from a dynamically fluctuating lattice background.
-Quantum interference and dephasing
+These experiments show that electronic momentum can be efficiently scrambled by a time-dependent environment without significant energy exchange.
+In other words, the lattice need not act as an energy sink in order to act as a momentum sink.
+This separation is precisely what is expected when electrons scatter quasi-elastically from a dynamically fluctuating lattice background.
+VIII.2 Quantum interference and dephasing
 The suppression of quantum-interference effects with increasing temperature is commonly described in terms of inelastic scattering processes that exchange energy with environmental degrees of freedom [23, 2].
@@
 [23, 27].Elastic scattering amplitudes, including diffuse non-Bragg amplitudes, are reduced by the well known Debye–Waller factor.
-Hot-electron relaxation
+VIII.3 Hot-electron relaxation
 A particularly direct separation of momentum and energy relaxation is provided by hot-electron experiments in metals [36, 30, 11, 12].
@@
 While inelastic electron–phonon scattering ultimately governs thermalization, the observed hierarchy suggests that additional scattering channels, with weak net energy exchange on electronic scales, can play an important role in momentum relaxation within the metallic state.
-VII Discussion
-The purpose of this section is twofold.
-First, we examine the consistency of the present framework with well-established experimental results, including the Wiedemann–Franz law, weak localization, and quantum oscillations.
-These considerations address potential concerns regarding the role of predominantly elastic momentum-relaxing channels.
+Johnson-noise thermometry provides a direct probe of electron–lattice energy exchange.
+Experiments consistently show that the electron temperature relaxes on timescales much longer than those associated with momentum relaxation inferred from transport[11].
+This separation, $\tau_{p}\ll\tau_{E}$, is difficult to reconcile with a picture in which momentum relaxation is intrinsically inelastic.
+It is, however, naturally explained if momentum is transferred predominantly through elastic or quasielastic processes, while energy relaxation proceeds through slower coupling to internal lattice degrees of freedom.
+VIII.4 Shubnikov-de Haas oscillations
+Quantum oscillation phenomena such as the Shubnikov-de Haas and de Haas-van Alphen effects provide a precise probe of electronic coherence and scattering in metals at low temperature [33, 3].
+These oscillations arise from Landau quantization and require that electrons execute many cyclotron orbits without loss of phase coherence.
+Within the conventional Lifshitz-Kosevich (LK) framework, the oscillation amplitude may be written as
+$A(T,B)\propto\frac{X}{\sinh X}\exp\!\left(-\frac{\pi}{\omega_{c}\tau_{q}}\right),\qquad X=\frac{2\pi^{2}k_{B}T}{\hbar\omega_{c}},$ (100)
+where the factor $X/\sinh X$ describes thermal smearing of the Fermi surface, while the Dingle factor
+$\exp\!\left(-\frac{\pi}{\omega_{c}\tau_{q}}\right)$ (101)
+encodes Landau-level broadening through the quantum lifetime $\tau_{q}$.
+In standard treatments, $\tau_{q}$ is associated with inelastic or quasi-inelastic processes, and is often assumed to become weakly temperature dependent at low $T$.
+From the present viewpoint, this interpretation is unnecessarily restrictive.
+When the lattice center-of-mass degree of freedom is retained, momentum relaxation need not proceed through phonon creation or annihilation.
+Instead, the electron can exchange momentum elastically with the lattice background via recoil of the zero mode, while total mechanical momentum is conserved.
+These processes are elastic in energy, but nonetheless randomize the electronic phase.
+As a result, the Dingle factor should be understood more generally as encoding phase randomization rather than strictly inelastic scattering.
+Elastic recoil processes provide a continuous channel for Landau-level broadening, even in the absence of phonon excitation.
+The temperature dependence then arises not from phonon occupation factors, but from the growth of available recoil phase space in the thermally fluctuating background.
+This leads naturally to a temperature-dependent quantum lifetime of the form
+$\frac{1}{\tau_{q}(T)}\sim T^{\alpha},$ (102)
+with $\alpha\sim 1$ in a Planckian regime, implying
+$\exp\!\left(-\frac{\pi}{\omega_{c}\tau_{q}(T)}\right)\;\sim\;\exp\!\left(-\mathrm{const}\,\frac{T}{B}\right).$ (103)
+Thus, the damping of quantum oscillations can remain strongly temperature dependent down to low $T$ without invoking inelastic scattering.
+The persistence of Shubnikov-de Haas oscillations over wide temperature ranges therefore indicates that substantial momentum randomization can occur through predominantly elastic processes, while phase coherence is degraded only gradually.
+This resolves the apparent tension between strong scattering and long-lived cyclotron coherence, and places quantum oscillation damping on the same footing as weak localization: both probe phase randomization in a dynamically fluctuating, but not necessarily energy-relaxing, environment.
+This interpretation is consistent with experimental observations of a marked temperature dependence of the Dingle factor, often extending to low $T$ with approximately power-law or linear-in-$T$ behavior, rather than saturation.
+In contrast to conventional theory, which attributes Landau-level broadening primarily to inelastic processes and therefore expects weak temperature dependence at low $T$, the present framework predicts a continued $T$-dependent damping arising from elastic recoil-driven phase randomization.
+IX Discussion
+Here we examine the consistency of the present framework with well-established experimental results, including the Wiedemann–Franz law, weak localization, and quantum oscillations.
 Second, we outline broader interpretive implications for Planckian transport, linear-in-$T$ resistivity, and the Mott–Ioffe–Regel crossover.
 The latter discussion is exploratory and is presented as a coherent physical interpretation rather than as a definitive claim.
-VII.1 Implications for the Lorenz ratio
+IX.1 Implications for the Lorenz ratio
 Elastic scattering randomizes momentum as effectively as inelastic scattering within standard transport formalisms.
@@
 A transport regime dominated by elastic momentum randomization is therefore fully compatible with the WF law, even when true energy relaxation is parametrically slower.
-VII.2 Weak localization
+IX.2 Weak localization
 Weak localization (WL) provides a clear illustration of the central role of elastic backscattering in low-temperature metallic transport.
 The phenomenon arises from constructive interference between pairs of time-reversed electronic trajectories that revisit the same spatial region after multiple scattering events [7, 23].
-The WL correction therefore requires that a substantial fraction of backscattering processes preserve phase coherence over the relevant timescale.
+It is often the case that the WL correction is important only with defects causing elastic backscattering.
+However the electrons still have to arrive back intact, so to speak (i.e.
+in phase), so other inelastic events cannot have decohered the electron on the way out and back.
+This is a tall order if any “natural” deflections are inelastic.
+elcrtherefore requires that a substantial fraction of backscattering processes preserve phase coherence over the relevant timescale.
 If scattering were strongly inelastic on the transport timescale, interference between time-reversed paths would be suppressed before a WL correction could develop.
@@
 where elastic momentum randomization produces diffusive transport, while slower temporal evolution of the lattice background reduces phase coherence without requiring substantial phonon excitation or rapid energy dissipation.
-VII.3 Shubnikov–de Haas oscillations
-Quantum oscillation phenomena such as the Shubnikov–de Haas and de Haas–van Alphen effects provide an independent probe of electronic scattering in metals at low temperature [29, 2].
-These oscillations arise from the quantization of electronic orbits in a magnetic field and reflect coherent motion around the Fermi surface over many cyclotron periods.
-The oscillation amplitude is controlled by the Dingle factor [8],
-$\exp\!\left(-\frac{\pi}{\omega_{c}\tau_{q}}\right),$ (37)
-where $\tau_{q}$ is the quantum lifetime associated with Landau-level broadening.
-This lifetime is sensitive to momentum-randomizing processes that broaden Landau levels, while sufficiently strong inelastic energy exchange suppresses the coherence required to complete cyclotron motion.
-Scattering events that transfer energies comparable to $\hbar\omega_{c}$ disrupt the phase coherence of the orbit.
-The observation of well-defined quantum oscillations over broad temperature ranges and down to relatively low magnetic fields therefore indicates that electrons can undergo many momentum-deflecting collisions while maintaining phase coherence over multiple cyclotron periods.
-As in weak localization, the temperature dependence of the oscillation amplitude reflects phase smearing due to temporal fluctuations of the scattering environment.
-These considerations are consistent with a regime in which momentum randomization can occur without rapid or large energy exchange on electronic scales.
-VII.4 Planckian diffusion and linear-in-$T$ resistivity
+IX.3 Planckian diffusion and linear-in-$T$ resistivity
 Much recent literature discusses Planckian behavior in terms of a microscopic relaxation time
@@
 From this viewpoint, the Planckian timescale may be understood not as a fundamental bound on inelastic scattering, but as an emergent parametrization of quantum-limited diffusion combined with equilibrium thermodynamics.
-It characterizes the rate at which charge spreads spatially rather than the rate at which energy must be irreversibly transferred to internal degrees of freedom.
-VIII Summary
-In the conventional clamped-lattice formulation of electron–phonon scattering, the lattice is treated as a fixed reference frame and its translational zero mode does not appear explicitly among the dynamical variables.
-Within that representation, momentum transfer is naturally described in terms of internal phonon excitations.
-When the lattice center-of-mass degree of freedom is retained explicitly, an additional elastic momentum-transfer channel becomes manifest.
-In a finite crystal, pseudomomentum can be exchanged between an electron and the lattice background while total mechanical momentum is conserved through recoil of the lattice zero mode.
-Such processes do not require phonon creation and may contribute significantly to momentum relaxation at low and intermediate temperatures.
-Making this channel explicit requires adopting a formulation in which (i) the lattice is treated as a finite, mechanically isolated object so that its center-of-mass motion remains dynamical, (ii) discrete translational symmetry is implemented symmetrically, with foreground and background pseudomomenta treated on equal footing, and (iii) the electron–lattice interaction is retained in its full density–density form.
-These steps do not alter the microscopic Hamiltonian; rather, they retain degrees of freedom that are commonly suppressed by choice of reference frame.
-When treated in this manner, elastic momentum exchange between electron and lattice background appears naturally within the same theoretical structure.
-In this framework, pseudomomentum is redistributed between the electron and the lattice background, while total mechanical momentum is conserved through recoil of the translational zero mode.
-Momentum relaxation need not coincide with phonon creation, and energy dissipation need not set the transport rate.
-This separation of roles clarifies how efficient momentum randomization can coexist with comparatively slow energy relaxation.
-Restoring explicit center-of-mass dynamics provides a more complete description of momentum bookkeeping in finite crystals.
-It offers a unified interpretation of experimental observations that indicate robust elastic scattering in clean materials, including weak localization, quantum oscillations, and ultrasonic attenuation.
-More broadly, the results suggest that in certain regimes, including some materials exhibiting linear-in-$T$ resistivity, momentum relaxation may be governed primarily by elastic scattering in a time-dependent lattice background [5], while inelastic processes play a secondary role in establishing global thermal equilibrium.
-In a perfectly clean, isolated, translation-invariant crystal, total momentum is conserved exactly.
-Such a system may exhibit internal equilibration and decay of electronic current, but it cannot exhibit finite DC resistivity in the strict steady-state sense.
-True resistivity requires a mechanism for removing total momentum from the system.
-IX Acknowledgments
+It characterizes the rate at which charge spreads spatially rather than the rate at which energy must be irreversibly transferred to internal degrees of freedom, and thus fits nicely into the main elastic and quasielastic theme of this paper.
+IX.4 Coulomb drag
+We began discussing Coulomb drag in section II.
+It is an ideal case in which the zero recoil mode, missing in the literature on coulomb drag, is likely the agent of the power law seen in experiments.
+Theoretical analyses of one-dimensional transport consistently emphasize the strong kinematic constraints imposed by simultaneous energy and momentum conservation, leading to the conclusion that relaxation processes should be strongly suppressed at low temperature [26, 18] .
+Experimentally, however, relaxation and drag in quantum wires are often observed to exhibit power-law temperature dependence rather than strong suppression.
+This discrepancy is typically attributed to additional mechanisms such as disorder, band curvature, or coupling to external degrees of freedom that relax the kinematic constraints.
+This will be the topic of a future publication.
+IX.5 Phonon drag
+In conventional transport theory, the phonon-drag contribution to the thermopower is attributed to momentum transfer from a nonequilibrium phonon population to the electrons, and is therefore tied to phonon creation and annihilation.
+This reflects a formulation in which the lattice center-of-mass (zero mode) is removed and momentum exchange is forced into the phonon sector.
+When the zero mode is retained, momentum transfer is no longer tied to phonon number.
+The electron couples to the lattice deformation field, and momentum can be exchanged with the lattice as a whole through recoil, without requiring phonon excitation.
+The conventional phonon-drag contribution therefore represents only a subset of a more general lattice-drag response.
+Accordingly, the thermopower should be written as
+$S=S_{\rm diff}+S_{\rm lat},\qquad S_{\rm lat}=S_{\rm el}+S_{\rm inel},$ (109)
+where $S_{\rm inel}$ is the usual phonon-drag term and $S_{\rm el}$ arises from elastic or quasielastic momentum transfer via lattice recoil.
+In regimes where scattering is predominantly elastic, $S_{\rm el}$ need not be small and may dominate the drag response.
+X Summary
+We have revisited electron-lattice scattering starting from the standard microscopic Hamiltonian, but retaining a degree of freedom that is nearly always removed: the translational center-of-mass (zero mode) of the lattice.
+In conventional formulations, this mode is eliminated (e.g., through Born-von Kármán boundary conditions), leaving momentum transfer to be carried entirely by internal phonon excitations.
+Together with the form of the first order Fröhlich expansion, changes in electronic momentum are tied to phonon creation or annihilation, and scattering is treated as intrinsically inelastic.
+When the lattice center of mass is retained explicitly, and the density-density form of the interaction is retained and not expanded, this picture changes qualitatively.
+Total mechanical momentum is conserved by recoil of the lattice as a whole, while pseudomomentum is redistributed between electron and background.
+A broad class of elastic and quasielastic momentum-transfer processes then becomes available, in which electronic momentum changes are balanced by the zero mode without requiring phonon excitation.
+This enlargement of the allowed phase space follows directly from the exact symmetries of the Hamiltonian.
+No new interaction has been introduced.
+Retaining the full density-density interaction reveals that elastic matrix elements arise already at the level of the state-to-state transition amplitude $\langle f|H_{\rm int}|i\rangle$, without invoking higher-order or multiphonon processes.
+The familiar phonon-mediated processes remain, but represent a restricted subset of a much larger set of allowed transitions.
+The resulting kinematics has broad implications.
+In one-dimensional systems, the need for thermally activated three-particle processes is removed, and power-law relaxation follows naturally.
+More generally, momentum relaxation need not be tied to energy dissipation: momentum can randomize rapidly while energy relaxes on much longer timescales.
+This resolves a long-standing tension between theory and experiment in clean metals, including observations from weak localization, quantum oscillations, hot-electron transport, and Coulomb drag.
+We emphasize that these conclusions are not model-dependent.
+Once the full momentum conservation law of the isolated electron-lattice system is enforced, the existence of elastic and quasielastic scattering channels is unavoidable.
+The restriction to phonon-mediated inelastic scattering is therefore not a fundamental feature of the physics, but a consequence of having discarded the lattice center-of-mass degree of freedom.
+XI Acknowledgments
 E.
@@
 Of course the box could be sitting on a table, which could be fixed to a floating laboratory, which could be upon something so massive that it has significant gravitational attraction, yet that too is floating in space…
-Appendix A Particle in a Free Box
-We consider simple models that clarify the treatment of momentum.
-The usual particle-in-a-box problem fixes the walls and therefore violates momentum conservation.
-Here the box is given finite mass and allowed to move freely, so that total momentum is exactly conserved.
-Consider a one-dimensional rigid box of mass $M$ and length $L$ with center-of-mass coordinate $R$, containing a particle of mass $m$ with coordinate $r$ (Fig. 3).
-Introduce total and relative coordinates
-$\mathcal{R}=\frac{mr+MR}{m+M},\qquad\rho=r-R,$
-with total mass $M_{T}=m+M$ and reduced mass
-$\mu=\frac{mM}{m+M}.$
-The hard-wall constraint is $|\rho|<L/2$.
-The Hamiltonian separates exactly:
-$H=-\frac{\hbar^{2}}{2M_{T}}\frac{\partial^{2}}{\partial\mathcal{R}^{2}}+\left[-\frac{\hbar^{2}}{2\mu}\frac{\partial^{2}}{\partial\rho^{2}}+V_{\rm hw}(\rho)\right].$
-Total momentum is conserved:
-$\hat{P}_{\rm tot}=-i\hbar\frac{\partial}{\partial\mathcal{R}}.$
-Case A: Flat interior
-The relative Hamiltonian is
-$H_{\rm rel}=-\frac{\hbar^{2}}{2\mu}\frac{d^{2}}{d\rho^{2}}+V_{\rm hw}(\rho).$
-The normalized eigenfunctions are
-$\phi_{n}(\rho)=\sqrt{\frac{2}{L}}\sin\!\left[\frac{n\pi}{L}\left(\rho+\frac{L}{2}\right)\right],\qquad n=1,2,\dots,$
-with energies
-$E_{n}=\frac{\hbar^{2}\pi^{2}n^{2}}{2\mu L^{2}}.$
-Any internal state may be written
-$\psi(\rho,t)=\sum_{n=1}^{\infty}c_{n}\,\phi_{n}(\rho)\,e^{-iE_{n}t/\hbar}.$
-A Gaussian superposition centered at $\rho_{0}$ with mean wavenumber $k_{0}$,
-$c_{n}=\mathcal{N}\exp\!\left[-\frac{(n-n_{0})^{2}}{4\sigma_{n}^{2}}\right]\exp\!\left[-ik_{n}\rho_{0}\right],\qquad k_{n}=\frac{n\pi}{L},$
-produces a localized packet
-$\psi(\rho,0)\approx A(\rho-\rho_{0})e^{ik_{0}(\rho-\rho_{0})}.$
-Although each $\phi_{n}$ carries no relative momentum, the particle and box momenta may be nonzero individually, while $P_{\rm tot}$ remains fixed.
-At a wall collision both reverse sign, preserving total momentum.
-Case B: Periodic interior
-Let the interior potential be periodic relative to the box:
-$V_{\rm per}(\rho)=V_{\rm per}(\rho+a).$
-Then
-$H_{\rm rel}=-\frac{\hbar^{2}}{2\mu}\frac{d^{2}}{d\rho^{2}}+V_{\rm per}(\rho)+V_{\rm hw}(\rho).$
-Eigenfunctions take Bloch standing-wave form
-$\varphi_{bn}(\rho)=\sqrt{\frac{2}{L}}u_{bn}(\rho)\sin\!\left(\frac{n\pi}{L}(\rho+L/2)\right),$
-with
-$u_{bn}(\rho+a)=u_{bn}(\rho).$
-Wavepackets are superpositions
-$\psi(\rho,0)=\sum_{n=1}^{\infty}A_{n}\varphi_{bn}(\rho).$
-Shared pseudomomentum
-Because $\rho=r-R$,
-$e^{iq\rho}=e^{iqr}e^{-iqR}.$
-Internal pseudomomentum $\hbar q$ is therefore shared between particle and box. For fixed total momentum $P_{0}$,
-$p_{r}=\frac{m}{m+M}P_{0}+\hbar q,\qquad P_{R}=\frac{M}{m+M}P_{0}-\hbar q,$
-so $p_{r}+P_{R}=P_{0}$.
-Cases C and D: Dynamical periodic background
-Let the periodic background be a dynamical elastic chain with atomic coordinates $R_{j}$:
+Appendix A Bloch waves form from elastic scattering of electrons in a lattice
+By way or reinforcing the conclusion that elastic internal electron scattering must be very common in pure metals, we raise a point about the rather violent elastic scattering that is taking place, cl oacked by the symmetry of Bloch waves.
+The formation of Bloch waves in a periodic crystal may be viewed as the result of coherent multiple elastic scattering from the atoms of the lattice[20].
+Band formation itself is therefore an elastic scattering phenomenon.
+One might then suppose that Bloch waves are somehow protected from further elastic deflection by phase coherence.
+However, this inference is incorrect, as illustrated in Fig.
+5.
+We extend this insight slighty and suppose some atom were temporarily slightly out of place due to thermal acoustic waves.
+The atoms would individually scatter as before, but now the phase locking is imperfect, and the electron can deflect.
+There is no new “phonon needed” physics that sets in due to truly minuscule thermal displacements.
+If the electron thus deflects, momentum is exchanged with the lattice without phonon excitation or energy dissipation.
+Appendix B Tracking Momentum, and the Effects of Zero Mode Removal
+Figure 6 helps illuminate the choices that can be made in treating electron-phonon interaction.
+It helps to underline that there is a great difference between treating the background as a dynamical body (with a translational zero mode) or as frozen, without recoil.
+Frozen is an impossibility in the real world, total mometum is always conserved, although this a self-consistent model.
+The question is, does that model lead to good real-world predictions, or does it mislead?
+Retaining the zero mode total momentum is not a matter of interpretation; it fixes the form of the wavefunction and alters the physical predictions in important ways.
+Translation invariance of the whole requires that the Hamiltonian depends only on the center of mass and coordinate differences, so eigenstates separate into a plane wave for the center of mass coordinate and an internal state that depends only on relative coordinates.
+B.1 Uplifting the particle in a box to the real world.
+We therefore free the textbook particle in a box from its traditional and physically impossible bondage in case A.: considering a“free floating box” of mass $M$ containing a particle of mass $m$, with particle coordinate $r$ and box center-of-mass coordinate $R$, introduce
+$\mathcal{R}=\frac{mr+MR}{m+M},\qquad$
+Then a complete set of eigenstates (delta-normalized in $P$) may be written as
+$\Psi_{P,n}(r,R)=\frac{1}{\sqrt{2\pi\hbar}}\,e^{iP\mathcal{R}/\hbar}\,\phi_{n}(r-R).$ (110)
+Equation (110) is the operative content of keeping the zero mode: the internal structure is literally a function of rRr-R. The conserved total mechanical momentum is
+$\hat{P}_{\rm tot}=-i\hbar\frac{\partial}{\partial\mathcal{R}},\qquad[H,\hat{P}_{\rm tot}]=0.$
+Any internal state may be expanded in Fourier components on |rR|<L/2|r-R|<L/2,
+$\phi_{n}(r-R)=\sum_{q}a_{nq}\,e^{iq(r-R)},$
+and each component carries the shared phase
+$e^{iq(r-R)}=e^{iqr}e^{-iqR}.$ (111)
+Thus, in the plane-wave decomposition of the relative coordinate, momentum exchange is enforced at the level of the wavefunction: each Fourier component correlates $\hbar q$ on the particle with $\hbar q$ on the background.
+Phonons and other internal excitations may redistribute momentum within the background.
+B.2 Box is corrugated
+Suppose the box is as above but now contains a rigid, corrugated periodic potential (Fig.
+6, B).
+Before the corrugation, the particle and the box each carry mechanical momentum, and only the total $P_{\rm tot}=p+P$ is conserved.
+In collisions, momentum is exchanged between the two in equal and opposite amounts, exactly as in free space.
+The introduction of a periodic potential does not eliminate this structure, but reorganizes it.
+Continuous translational symmetry is reduced to discrete translational symmetry, and the corresponding conserved quantities become pseudomomenta defined modulo reciprocal lattice vectors.
+Importantly, the two-body structure of momentum exchange is preserved.
+The particle now carries a pseudomomentum $\hbar k$, while the background box carries a corresponding pseudomomentum.
+These two pseudomomenta are exchanged in equal and opposite amounts during scattering, just as the mechanical momenta were exchanged before the corrugation was introduced.
+Their sum remains fixed and corresponds to the conserved total momentum of the isolated system.
+Thus the corrugation does not replace two mechanical momenta with a single pseudomomentum; rather, it produces two pseudomomenta, one for the particle and one for the background.
+The center-of-mass degree of freedom of the box continues to carry the total mechanical momentum, while the periodic structure introduces a new, discrete labeling of the same underlying exchange process.
+The situation depicted in Fig.
+6, B therefore retains the essential physics of momentum exchange: the lattice remains a dynamical participant, even though its role is partially obscured by the use of pseudomomentum.
+B.3 Elastic Chain in a Box, Free CM (phonons plus the zero mode)
+Now take the background to be a dynamical chain (pink atoms) of total mass $M_{\rm lat}$ inside the box, with atomic coordinates $R_{j}$ and a center-of-mass coordinate $R_{\rm lat}$ (Fig.
+6, C).
+A minimal model is
 $H=\frac{p_{r}^{2}}{2m}+\sum_{j}\frac{p_{j}^{2}}{2M}+\frac{K}{2}\sum_{j}(R_{j+1}-R_{j}-a)^{2}+H_{\rm int}.$
-With density–density coupling
+With density-density coupling
 $H_{\rm int}=\int dr\,dr^{\prime}\,\rho_{p}(r)V(r-r^{\prime})\rho_{\rm lat}(r^{\prime}),$
-where
-$\rho_{p}(r)=\delta(r-\hat{r}),\qquad\rho_{\rm lat}(r)=\sum_{j}\delta(r-\hat{R}_{j}),$
 the Hamiltonian depends only on coordinate differences and is invariant under rigid translation
 $r\rightarrow r+a_{0},\qquad R_{j}\rightarrow R_{j}+a_{0}.$
-Thus total momentum
-$P_{\rm tot}=p_{r}+\sum_{j}p_{j}$
-is exactly conserved. Diagonalizing the lattice,
-$R_{j}=R_{\rm lat}+\sum_{q}u_{q}e^{iqja},$
-reveals the zero mode $R_{\rm lat}$. The density operator contains the factor
-$e^{-iqR_{\rm lat}},$
-showing explicitly that momentum $\hbar q$ transferred to the particle is balanced by recoil of the lattice center of mass.
-Internal phonons redistribute momentum but are not required for momentum conservation.
-Appendix B Elastic and inelastic scattering in a tight-binding, isolated one-dimensional chain
-We consider a finite one-dimensional tight-binding chain of $N$ identical atoms, isolated in space, with nearest-neighbor springs and an electron moving in a basis of localized orbitals.
-The chain possesses an exact translational zero mode (center-of-mass motion).
-An electron can therefore exchange momentum with the lattice as a whole without necessarily exciting internal phonon modes.
-This opens a phonon-diagonal (internally elastic) backscattering channel for an electronic wavepacket, with the required momentum reversal balanced by center-of-mass recoil.
-Hamiltonian and separation of the CM mode
-Write the atomic coordinates as
-$x_{n}=R+na+u_{n},\qquad\sum_{n=1}^{N}u_{n}=0,$ (41)
-where $R$ is the center-of-mass coordinate, $a$ the equilibrium lattice spacing, and $\{u_{n}\}$ the internal displacements. The lattice Hamiltonian separates into CM and internal parts,
-$H_{\rm lat}=\frac{P_{R}^{2}}{2M_{\rm tot}}+\sum_{n=1}^{N}\frac{p_{n}^{2}}{2M}+\frac{K}{2}\sum_{n=1}^{N-1}(u_{n+1}-u_{n})^{2},$ (42)
-with $M_{\rm tot}=NM$ and $\sum_{n}p_{n}=0$.
-The electron is described by a nearest-neighbor tight-binding Hamiltonian,
-$H_{e}=-\sum_{n=1}^{N-1}t_{n}\left(c_{n+1}^{\dagger}c_{n}+c_{n}^{\dagger}c_{n+1}\right),$ (43)
-with bond-length-dependent hopping
-$t_{n}\equiv t(x_{n+1}-x_{n})\simeq t_{0}-\alpha(u_{n+1}-u_{n}).$ (44)
-Thus the electron–lattice coupling depends only on internal coordinates $u_{n+1}-u_{n}$; the CM coordinate $R$ drops out exactly.
-Phonon-diagonal backscattering via CM recoil
-Because the full electron+lattice Hamiltonian is translationally invariant, the total momentum is exactly conserved.
-Since the band energy satisfies $\varepsilon(k)=\varepsilon(-k)$, electronic reversal $k\to-k$ does not change the electronic energy.
-Consider an incoming wavepacket narrowly peaked at +k+k, an internal lattice eigenstate $|\nu\rangle$, and a CM momentum eigenstate $|P_{R}\rangle$.
-A phonon-diagonal backscattering channel has the form
-$|k\rangle\otimes|\nu\rangle\otimes|P_{R}\rangle\longrightarrow|-k\rangle\otimes|\nu\rangle\otimes|P_{R}+2\hbar k\rangle,$ (45)
-i.e. the internal state is unchanged while the CM absorbs the recoil $2\hbar k$.
-The associated change in CM kinetic energy is
-$\Delta E_{\rm CM}=\frac{(P_{R}+2\hbar k)^{2}-P_{R}^{2}}{2M_{\rm tot}}=\frac{4\hbar k\,P_{R}+4\hbar^{2}k^{2}}{2M_{\rm tot}}.$ (46)
-Exact energy conservation with strictly unchanged internal state requires $\Delta E_{\rm CM}=0$, i.e.
-$P_{R}=-\hbar k.$ (47)
-For generic $P_{R}$ there is a recoil-energy mismatch
-$\Delta E_{\rm recoil}=\frac{(P_{R}+2\hbar k)^{2}-P_{R}^{2}}{2M_{\rm tot}}=\frac{2\hbar k\,(P_{R}+\hbar k)}{M_{\rm tot}}.$ (48)
-However, for a macroscopic chain $M_{\rm tot}\propto N$ so $|\Delta E_{\rm recoil}|\sim 1/N$ becomes parametrically small.
-For a wavepacket of spatial extent $L$ with intrinsic energy width $\Delta E_{\rm pkt}\sim\hbar v_{g}/L$, or for scattering occurring over a finite time $\tau_{\rm scatt}$ with uncertainty $\hbar/\tau_{\rm scatt}$, the process is effectively elastic whenever $\Delta E_{\rm recoil}$ is below the available energy resolution.
-In a number-state phonon description, any residual mismatch may be accommodated by very soft internal modes, but the momentum reversal $k\to-k$ is carried by CM recoil; internal excitations, if present, serve primarily as an energetically low-cost bookkeeping channel.
-Role of finite temperature
-At finite temperature $T$, the CM momentum is thermally distributed,
-$f(P_{R})\propto\exp\!\left(-\frac{P_{R}^{2}}{2M_{\rm tot}k_{B}T}\right),$ (49)
-with width $\sigma_{P}=\sqrt{M_{\rm tot}k_{B}T}$.
-The resonant value $P_{R}=-\hbar k$ therefore occurs with nonzero probability density for any $T>0$, so strictly phonon-diagonal backscattering events exist at finite temperature in an isolated chain.
-Bulk and boundary backscattering
-The CM-recoil channel does not require the electron to reach the ends: it can occur in the bulk from the static component of thermal distortions (phonon-diagonal scattering). The SSH coupling
-$H_{\rm int}=\alpha\sum_{n=1}^{N-1}(u_{n+1}-u_{n})\left(c_{n+1}^{\dagger}c_{n}+c_{n}^{\dagger}c_{n+1}\right)$
-connects $|k\rangle$ and $|-k\rangle$ through the $2k$ Fourier component of the bond-distortion field. CM recoil is enforced by the CM translation factor. In particular, the CM shift operator produces
-$\langle P_{R}^{\prime}|e^{-i(2k)R}|P_{R}\rangle\propto\delta(P_{R}^{\prime}-P_{R}-2\hbar k),$ (50)
-so momentum conservation is implemented by a delta function (or Kronecker delta in a finite quantization volume), not by an inner product between distinct CM momentum eigenstates.
-Free boundaries can also reflect a wavepacket once it reaches an end.
-Such reflections reverse $k$ and are likewise accompanied by CM recoil.
-Boundaries provide a location for reflection but do not constitute a momentum sink; both bulk (distortion-mediated) and boundary reflections obey the same conservation principle enforced by the CM zero mode.
-Other scattering channels
-In addition to the phonon-diagonal CM-recoil channel, inelastic and hybrid processes are allowed.
-Single-phonon emission and absorption
-If momentum is transferred to internal modes, the internal lattice state changes:
-$|k\rangle\otimes|\{n_{q}\}\rangle\to|k^{\prime}\rangle\otimes|\{n_{q}^{\prime}\}\rangle,\qquad n_{q}^{\prime}=n_{q}\pm 1.$
-Energy conservation requires
-$\varepsilon(k^{\prime})=\varepsilon(k)\pm\hbar\omega_{q}+\Delta E_{\rm CM}.$
-A Golden-Rule rate has the form
-$W_{k\to k^{\prime};q}=\frac{2\pi}{\hbar}|M_{k,k^{\prime};q}|^{2}\,\delta\!\left(\varepsilon(k^{\prime})-\varepsilon(k)\mp\hbar\omega_{q}-\Delta E_{\rm CM}\right),$
-with matrix elements scaling as
-$M_{k,k^{\prime};q}\sim\alpha\sqrt{\frac{\hbar}{2M\omega_{q}}}\,F_{k,k^{\prime};q}.$
-Multi-phonon and hybrid CM–phonon processes
-Higher-order multi-phonon processes and hybrid processes involving both CM recoil and internal phonon exchange relax the strict resonance condition (47) at the cost of reduced matrix elements.
-Kinematic summary
-  • Phonon-diagonal channel: $k\to-k$ with unchanged internal state is allowed by translational invariance and proceeds via CM recoil; it is strictly elastic only on resonance (47) and effectively elastic for large $N$.
-  • Inelastic channels: internal phonon emission/absorption become important when the phonon spectrum permits energy matching.
-  • Finite size: discrete internal spectra constrain inelastic channels and can enhance the relative importance of CM-mediated phonon-diagonal scattering.
-Practical consequences
-  • In wavepacket simulations, phonon-diagonal backscattering is identified by momentum reversal with unchanged internal state and a compensating shift of $P_{R}$ by $2\hbar k$.
-  • Bulk and boundary reflections differ only in spatial location; both satisfy the same momentum conservation enforced by the CM zero mode.
-  • Neglecting the CM zero mode (implicit clamping) qualitatively changes the available scattering channels.
+Consequently the conserved quantity is the true total mechanical momentum
+$P_{\rm tot}=P_{e}+P_{\rm lat},\qquad[H,P_{\rm tot}]=0.$
+Momentum exchange is enforced directly through the relative coordinate. Fourier transforming the lattice density yields
+$\rho_{\rm lat}(q)=e^{-iqR_{\rm lat}}\rho^{\rm int}_{q},\qquad H_{\rm int}\propto e^{iq(r-R_{\rm lat})}.$
+The phase factor $e^{iq(r-R_{\rm lat})}$ encodes the fundamental constraint: momentum $\hbar q$ transferred to the electron is accompanied by $\hbar q$ carried by the lattice center of mass.
+Internal degrees of freedom describe only relative motion and carry no net mechanical momentum.
+The translational zero mode $R_{\rm lat}$ is therefore the unique carrier of the compensating momentum.
+Internal modes (phonons) may change the energy of the system, but they are not required to enforce momentum conservation.
+Deleting the zero mode is not an innocuous simplification.
+It removes the only degree of freedom capable of carrying recoil and eliminates the shared-coordinate structure $\phi(r-R)$ from the Hilbert space.
+The resulting theory no longer represents true mechanical momentum conservation; instead it replaces it with pseudomomentum bookkeeping within a reduced space.
+Momentum exchange is then no longer enforced through the relative coordinate rRr-R, but appears only as a selection rule requiring phonon pseudomomentum to balance electron pseudomomentum (up to Umklapp).
+Thus the apparent kinematic necessity of phonon production does not arise from fundamental principles.
+It is imposed by amputating the translational degree of freedom that would otherwise carry the compensating momentum.
+B.4 Free Chain (no box), Free CM (translation invariance made manifest)
+Removing the box entirely leaves a floating chain.
+This is simply Fig.
+6, C with boundary constraints eliminated.
+The lesson is unchanged but becomes explicit: the conserved momentum is the generator of rigid translation of all coordinates.
+Any description that retains translational invariance necessarily contains the shared phase $e^{iq(r-R_{\rm lat})}$, since it is the Fourier representation of the relative coordinate.
+This phase structure is the direct expression of momentum conservation in the full system.
+B.5 Removing the Zero Mode
+The conventional electron-phonon kinematics is obtained by a construction that is the Hilbert-space analog of “nailing the box down,” but without introducing an external dynamical reservoir.
+The translational zero mode is removed at the outset.
+One expands the lattice displacement field in normal modes and omits the $q=0$ component.
+The resulting description retains only internal degrees of freedom.
+With the zero mode removed, continuous translation symmetry is no longer present, and the conserved quantity is not the true mechanical momentum.
+What remains is discrete lattice translation symmetry, with conserved crystal pseudomomentum
+$K_{\rm tot}=K_{e}+K_{\rm ph}\quad(\mathrm{mod}\;G),\qquad K_{\rm ph}=\sum_{q}\hbar q\,a_{q}^{\dagger}a_{q},$
+and the corresponding selection rule
+$\Delta K_{e}+\Delta K_{\rm ph}=Gm.$
+In this representation, momentum transfer is confined to phonons: changes in the electron pseudomomentum must be balanced entirely by phonon pseudomomentum, up to Umklapp.
+The lattice center-of-mass degree of freedom, which would otherwise provide a direct channel for momentum exchange through rigid recoil, is absent.
+The apparent kinematic necessity of phonon emission or absorption is therefore not a physical result.
+It is a consequence of having eliminated the translational degree of freedom that would otherwise carry the compensating momentum.
+It is not a choice one normally would wish to make, because it changes the predicted physical processes.
+It is surprising perhaps, but it does.
+The “thermodynamic limit” does not protect us here, and if it seems that it does, the limit $N\to\infty$ has not been taken properly.
+Jumping to $N=\infty$ is not taking the limit of course.
 Appendix C Whither Umklapp?
-The traditional role of Umklapp for finite resistivity in an infinite, perfect crystal arises from freezing the background and projecting out its pseudomomentum.
-Once the background is treated as dynamical, with an explicit translational degree of freedom, momentum relaxation can occur through elastic pseudomomentum exchange between electron and lattice.
-Phonons then represent one possible basis for describing background dynamics rather than the exclusive channel for momentum transfer.
-This perspective clarifies long-standing ambiguities surrounding Umklapp and explains why momentum relaxation is observed experimentally even in regimes where Umklapp processes are expected to be suppressed.
-It also underscores that the association of Umklapp with dissipation and entropy production is not intrinsic, but follows from additional assumptions about the background.
-The expectation that Umklapp processes are necessary for finite resistivity in a perfect crystal has deep roots in transport theory.
-Peierls first distinguished between normal and Umklapp processes in his analysis of phonon thermal conductivity, showing that only the latter can relax total crystal momentum in a closed system.
-This reasoning was later extended to electrons by Bloch, Peierls, Ziman, and others, and appears explicitly in standard texts (e.g.
-Peierls, Quantum Theory of Solids; Ziman, Electrons and Phonons).
-If the Fermi surface lies well inside the Brillouin zone, or at low temperature where large-wavevector phonons are thermally suppressed, Umklapp processes are expected to be exponentially rare.
-Under these conditions conventional theory predicts extremely long momentum lifetimes in a clean crystal.
-The present formulation revisits this conclusion by restoring the dynamical translational degree of freedom of the background.
-Momentum exchange between subsystems does not require phonon creation, and the strict identification of Umklapp with momentum relaxation relies on additional assumptions about background clamping or environmental coupling.
-Umklapp processes are best understood as a classification of crystal-momentum conservation rather than as the fundamental cause of momentum relaxation.
-Their elevated role in conventional transport theory stems from treating the lattice background as inert.
-Recognizing the background as a dynamical object carrying its own pseudomomentum restores a more symmetric and physically complete description in which elastic momentum exchange is allowed and Umklapp is no longer indispensable.
+The thermodynamic limit and Umklapp processes are often invoked to account for momentum relaxation in perfect crystals.
+These mechanisms are valid within the conventional formulation, but their apparent necessity reflects the prior removal of the lattice center-of-mass (zero-mode) degree of freedom.
+In the standard phonon description, momentum conservation is imposed only at the level of crystal pseudomomentum,
+$K_{\rm tot}=K_{e}+K_{\rm ph}\quad(\mathrm{mod}\;G),$
+and electronic momentum can relax only through phonon processes, supplemented by Umklapp when reciprocal lattice vectors are invoked.
+This framework is kinematically restrictive: the phonon system tracks momentum only modulo $G$, and the available channels are limited to discrete reciprocal lattice transfers.
+When the zero mode is retained, this restriction is lifted.
+The lattice background can recoil and absorb arbitrary mechanical momentum, and the scattering problem is no longer confined to Umklapp processes.
+Instead, momentum transfer occurs within a continuous manifold of elastic and quasielastic processes involving the full dynamical background.
+To make this explicit, we separate a single electron (the foreground) from the rest of the system (the background), which includes all ions and all other electrons. The total mechanical momentum is
+$\hat{P}_{\rm tot}=\hat{p}+\hat{P}_{\rm bg},\qquad\hat{P}_{\rm bg}=\sum_{\alpha\in\mathrm{bg}}\hat{P}_{\alpha}.$ (112)
+The operator $\hat{P}_{\rm bg}$ generates rigid translation of the background and constitutes its center-of-mass (zero-mode) momentum.
+Internal phonon modes describe only relative motion and do not furnish an independent contribution to $\hat{P}_{\rm bg}$.
+Thus any change in the electron’s mechanical momentum must be balanced by the background zero mode.
+This makes clear what is implicit in Umklapp.
+When a reciprocal lattice vector $G$ appears in a scattering process, the associated momentum $\hbar G$ cannot be attributed to the electron alone, nor is it carried by internal phonon modes.
+It must be carried by the lattice as a whole.
+In the conventional formulation, however, the zero mode that would carry this momentum has been removed, and the role of recoil is replaced by bookkeeping modulo $G$.
+The traditional role assigned to Umklapp in producing finite resistivity therefore arises within a restricted description in which the background is effectively clamped.
+Once the background is treated dynamically, electronic momentum need not decay through Umklapp.
+Instead, it can be transferred elastically to the lattice as a whole.
+Umklapp remains a useful classification of pseudomomentum conservation modulo reciprocal lattice vectors.
+It is not, however, the unique or even generic mechanism for momentum relaxation in a real crystal, where the background is free to recoil.
+Appendix D The landscape seen by an electron
+At temperatures of order $T\sim 100\,$K, thermally excited acoustic distortions in a typical metal generate substantial deformation-potential landscapes on nanometer length scales.
+Using standard values for the deformation potential ($g\sim 5$$15\,\mathrm{eV}$), sound velocity ($v_{s}\sim 3$$5\times 10^{3}\,\mathrm{m/s}$), and the characteristic thermal phonon wavevector $q_{T}\sim k_{B}T/\hbar v_{s}$, one finds order-of-magnitude estimates for the resulting deformation-potential field gradients of
+$|\nabla V_{\rm def}|\sim 10^{4}\text{–}10^{6}\ \mathrm{V/cm}\qquad(T\sim 100\,\mathrm{K}).$ (113)
+Over a representative lateral scale of $L\sim 10\,$nm, this corresponds to peak-to-peak variations of the screened scalar potential of order
+$\Delta V_{\rm pp}\sim|\nabla V_{\rm def}|\,L\sim 10\text{–}300\ \mathrm{mV}.$ (114)
+[…]
+[…]